GNU tar

GNU tar: an archiver tool

FTP release, version 1.16.1, 7 December 2006

John Gilmore, Jay Fenlason et al.

This manual is for GNU tar (version 1.16.1, 7 December 2006), which creates and extracts files from archives. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, 1994, 1995, 1996, 1997, 1999, 2000, 2001, 2003, 2004, 2005, 2006 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover Texts being “A GNU Manual,” and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled "GNU Free Documentation License".

(a) The FSF’s Back-Cover Text is: “You are free to copy and modify this GNU Manual. Buying copies from GNU Press supports the FSF in developing GNU and promoting software freedom.”

1 Introduction

GNU tar creates and manipulates archives which are actually collections of many other files; the program provides users with an organized and systematic method for controlling a large amount of data. The name “tar” originally came from the phrase “Tape ARchive”, but archives need not (and these days, typically do not) reside on tapes.

1.1 What this Book Contains

The first part of this chapter introduces you to various terms that will recur throughout the book. It also tells you who has worked on GNU tar and its documentation, and where you should send bug reports or comments.

The second chapter is a tutorial (see Chapter 2 [Tutorial], page 5) which provides a gentle introduction for people who are new to using tar. It is meant to be self contained, not requiring any reading from subsequent chapters to make sense. It moves from topic to topic in a logical, progressive order, building on information already explained.

Although the tutorial is paced and structured to allow beginners to learn how to use tar, it is not intended solely for beginners. The tutorial explains how to use the three most frequently used operations (‘create’, ‘list’, and ‘extract’) as well as two frequently used options (‘file’ and ‘verbose’). The other chapters do not refer to the tutorial frequently; however, if a section discusses something which is a complex variant of a basic concept, there may be a cross reference to that basic concept. (The entire book, including the tutorial, assumes that the reader understands some basic concepts of using a Unix-type operating system; see Chapter 2 [Tutorial], page 5.)

The third chapter presents the remaining five operations, and information about using tar options and option syntax.

The other chapters are meant to be used as a reference. Each chapter presents everything that needs to be said about a specific topic.

One of the chapters (see Chapter 7 [Date input formats], page 109) exists in its entirety in other GNU manuals, and is mostly self-contained. In addition, one section of this manual (see [Standard], page 165) contains a big quote which is taken directly from tar sources.

In general, we give both long and short (abbreviated) option names at least once in each section where the relevant option is covered, so that novice readers will become familiar with both styles. (A few options have no short versions, and the relevant sections will indicate this.)

1.2 Some Definitions

The tar program is used to create and manipulate tar archives. An archive is a single file which contains the contents of many files, while still identifying the names of the files, their owner(s), and so forth. (In addition, archives record access permissions, user and group, size in bytes, and data modification time. Some archives also record the file names in each archived directory, as well as other file and directory information.) You can use tar to create a new archive in a specified directory.

The files inside an archive are called members. Within this manual, we use the term file to refer only to files accessible in the normal ways (by ls, cat, and so forth), and the term member to refer only to the members of an archive. Similarly, a file name is the name of a file, as it resides in the file system, and a member name is the name of an archive member within the archive.

The term extraction refers to the process of copying an archive member (or multiple members) into a file in the file system. Extracting all the members of an archive is often called extracting the archive. The term unpack can also be used to refer to the extraction of many or all the members of an archive. Extracting an archive does not destroy the archive’s structure, just as creating an archive does not destroy the copies of the files that exist outside of the archive. You may also list the members in a given archive (this is often thought of as “printing” them to the standard output, or the command line), or append members to a pre-existing archive. All of these operations can be performed using tar.

1.3 What tar Does

The tar program provides the ability to create tar archives, as well as various other kinds of manipulation. For example, you can use tar on previously created archives to extract files, to store additional files, or to update or list files which were already stored.

Initially, tar archives were used to store files conveniently on magnetic tape. The name tar comes from this use; it stands for tape archiver. Despite the utility’s name, tar can direct its output to available devices, files, or other programs (using pipes). tar may even access remote devices or files (as archives).

You can use tar archives in many ways. We want to stress a few of them: storage, backup, and transportation.

Storage Often, tar archives are used to store related files for convenient file transfer over a network. For example, the GNU Project distributes its software bundled into tar archives, so that all the files relating to a particular program (or set of related programs) can be transferred as a single unit. A magnetic tape can store several files in sequence. However, the tape has no names for these files; it only knows their relative position on the tape. One way to store several files on one tape and retain their names is by creating a tar archive. Even when the basic transfer mechanism can keep track of names, as FTP can, the nuisance of handling multiple files, directories, and multiple links makes tar archives useful. Archive files are also used for long-term storage. You can think of this as transportation from the present into the future. (It is a science-fiction idiom that you can move through time as well as in space; the idea here is that tar can be used to move archives in all dimensions, even time!)

Backup Because the archive created by tar is capable of preserving file information and directory structure, tar is commonly used for performing full and incremental backups of disks. A backup puts a collection of files (possibly pertaining to many users and projects) together on a disk or a tape. This guards against accidental destruction of the information in those files. GNU tar has special features that allow it to be used to make incremental and full dumps of all the files in a file system.

Transportation You can create an archive on one system, transfer it to another system, and extract the contents there. This allows you to transport a group of files from one system to another.

1.4 How tar Archives are Named

Conventionally, tar archives are given names ending with ‘.tar’. This is not necessary for tar to operate properly, but this manual follows that convention in order to accustom readers to it and to make examples more clear.

Often, people refer to tar archives as “tar files,” and archive members as “files” or “entries”. For people familiar with the operation of tar, this causes no difficulty. However, in this manual, we consistently refer to “archives” and “archive members” to make learning to use tar easier for novice users.

1.5 GNU tar Authors

GNU tar was originally written by John Gilmore, and modified by many people. The GNU enhancements were written by Jay Fenlason, then Joy Kendall, and the whole package has been further maintained by Thomas Bushnell, n/BSG, François Pinard, Paul Eggert, and finally Sergey Poznyakoff with the help of numerous and kind users.

We wish to stress that tar is a collective work, and owes much to all those people who reported problems, offered solutions and other insights, or shared their thoughts and suggestions. An impressive, yet partial list of those contributors can be found in the ‘THANKS’ file from the GNU tar distribution.

Jay Fenlason put together a draft of a GNU tar manual, borrowing notes from the original man page from John Gilmore. This was withdrawn in version 1.11. Thomas Bushnell, n/BSG and Amy Gorin worked on a tutorial and manual for GNU tar. François Pinard put version 1.11.8 of the manual together by taking information from all these sources and merging them. Melissa Weisshaus finally edited and redesigned the book to create version 1.12. The book for versions from 1.14 up to 1.16.1 were edited by the current maintainer, Sergey Poznyakoff.

For version 1.12, Daniel Hagerty contributed a great deal of technical consulting. In particular, he is the primary author of Chapter 5 [Backups], page 73.

In July, 2003 GNU tar was put on CVS at savannah.gnu.org (see

http://savannah.gnu.org/projects/tar), and active development and

maintenance work has started again. Currently GNU tar is being maintained by Paul Eggert, Sergey Poznyakoff and Jeff Bailey.

Support for POSIX archives was added by Sergey Poznyakoff.

1.6 Reporting bugs or suggestions

If you find problems or have suggestions about this program or manual, please report them to ‘bug-tar@gnu.org’.

When reporting a bug, please be sure to include as much detail as possible, in order to reproduce it. .

2 Tutorial Introduction to tar

This chapter guides you through some basic examples of three tar operations: ‘--create’, ‘--list’, and ‘--extract’. If you already know how to use some other version of tar, then you may not need to read this chapter. This chapter omits most complicated details about how tar works.

2.1 Assumptions this Tutorial Makes

This chapter is paced to allow beginners to learn about tar slowly. At the same time, we will try to cover all the basic aspects of these three operations. In order to accomplish both of these tasks, we have made certain assumptions about your knowledge before reading this manual, and the hardware you will be using:

• Before you start to work through this tutorial, you should understand what the terms “archive” and “archive member” mean (see Section 1.2 [Definitions], page 1). In addition, you should understand something about how Unix-type operating systems work, and you should know how to use some basic utilities. For example, you should know how to create, list, copy, rename, edit, and delete files and directories; how to change between directories; and how to figure out where you are in the file system. You should have some basic understanding of directory structure and how files are named according to which directory they are in. You should understand concepts such as standard output and standard input, what various definitions of the term “argument” mean, and the differences between relative and absolute path names.

• This manual assumes that you are working from your own home directory (unless we state otherwise). In this tutorial, you will create a directory to practice tar commands in. When we show path names, we will assume that those paths are relative to your home directory. For example, my home directory path is ‘/home/fsf/melissa’. All of my examples are in a subdirectory of the directory named by that path name; the subdirectory is called ‘practice’.

• In general, we show examples of archives which exist on (or can be written to, or worked with from) a directory on a hard disk. In most cases, you could write those archives to, or work with them on any other device, such as a tape drive. However, some of the later examples in the tutorial and next chapter will not work on tape drives. Additionally, working with tapes is much more complicated than working with hard disks. For these reasons, the tutorial does not cover working with tape drives. See Chapter 9 [Media], page 137, for complete information on using tar archives with tape drives.

2.2 Stylistic Conventions

In the examples, ‘$’ represents a typical shell prompt. It precedes lines you should type; to make this more clear, those lines are shown in this font, as opposed to lines which represent the computer’s response; those lines are shown in this font, or sometimes ‘like this’.

2.3 Basic tar Operations and Options

tar can take a wide variety of arguments which specify and define the actions it will have on the particular set of files or the archive. The main types of arguments to tar fall into one of two classes: operations, and options.

Some arguments fall into a class called operations; exactly one of these is both allowed and required for any instance of using tar; you may not specify more than one. People sometimes speak of operating modes. You are in a particular operating mode when you have specified the operation which specifies it; there are eight operations in total, and thus there are eight operating modes.

The other arguments fall into the class known as options. You are not required to specify any options, and you are allowed to specify more than one at a time (depending on the way you are using tar at that time). Some options are used so frequently, and are so useful for helping you type commands more carefully that they are effectively “required”. We will discuss them in this chapter.

You can write most of the tar operations and options in any of three forms: long (mnemonic) form, short form, and old style. Some of the operations and options have no short or “old” forms; however, the operations and options which we will cover in this tutorial have corresponding abbreviations. We will indicate those abbreviations appropriately to get you used to seeing them. (Note that the “old style” option forms exist in GNU tar for compatibility with Unix tar. In this book we present a full discussion of this way of writing options and operations (see Section 3.3.3 [Old Options], page 26), and we discuss the other two styles of writing options (See Section 3.3.1 [Long Options], page 24, and see Section 3.3.2 [Short Options], page 25).

In the examples and in the text of this tutorial, we usually use the long forms of operations and options; but the “short” forms produce the same result and can make typing long tar commands easier. For example, instead of typing

     tar --create --verbose --file=afiles.tar apple angst aspic

you can type

     tar -c -v -f afiles.tar apple angst aspic

or even

     tar -cvf afiles.tar apple angst aspic

For more information on option syntax, see Section 4.2 [Advanced tar], page 52. In discussions in the text, when we name an option by its long form, we also give the corresponding short option in parentheses.

The term, “option”, can be confusing at times, since “operations” are often lumped in with the actual, optional “options” in certain general class statements. For example, we just talked about “short and long forms of options and operations”. However, experienced tar users often refer to these by shorthand terms such as, “short and long options”. This term assumes that the “operations” are included, also. Context will help you determine which definition of “options” to use.

Similarly, the term “command” can be confusing, as it is often used in two different ways. People sometimes refer to tar “commands”. A tar command is the entire command line of user input which tells tar what to do — including the operation, options, and any arguments (file names, pipes, other commands, etc.). However, you will also sometimes hear the term “the tar command”. When the word “command” is used specifically like this, a person is usually referring to the tar operation, not the whole line. Again, use context to figure out which of the meanings the speaker intends.

2.4 The Three Most Frequently Used Operations

Here are the three most frequently used operations (both short and long forms), as well as a brief description of their meanings. The rest of this chapter will cover how to use these operations in detail. We will present the rest of the operations in the next chapter.

‘--create’

‘-c’ Create a new tar archive.

‘--list’

‘-t’ List the contents of an archive.

‘--extract’

‘-x’ Extract one or more members from an archive.

2.5 Two Frequently Used Options

To understand how to run tar in the three operating modes listed previously, you also need to understand how to use two of the options to tar: ‘--file’ (which takes an archive file as an argument) and ‘--verbose’. (You are usually not required to specify either of these options when you run tar, but they can be very useful in making things more clear and helping you avoid errors.)

The ‘--file’ Option

‘--file=archive-name’
‘-f archive-name’

Specify the name of an archive file. You can specify an argument for the ‘--file=archive-name’ (‘-f

archive-name’) option whenever you use tar; this option determines the name of the archive file that tar will work on.

If you don’t specify this argument, then tar will examine the environment variable TAPE. If it is set, its value will be used as the archive name. Otherwise, tar will use the default archive, determined at the compile time. Usually it is standard output or some physical tape drive attached to your machine (you can verify what the default is by running tar --show-defaults, see Section 3.6 [defaults], page 46). If there is no tape drive attached, or the default is not meaningful, then tar will print an error message. The error message might look roughly like one of the following:

     tar: can't open /dev/rmt8 : No such device or address
     tar: can't open /dev/rsmt0 : I/O error

To avoid confusion, we recommend that you always specify an archive file

name by using ‘--file=archive-name’ (‘-f archive-name’) when writing

your tar commands. For more information on using the ‘--file=archive-name’ (‘-f archive-name’) option, see Section 6.1 [file], page 87.

The ‘--verbose’ Option

‘--verbose’

‘-v’ Show the files being worked on as tar is running.

‘--verbose’ (‘-v’) shows details about the results of running tar. This can be especially useful when the results might not be obvious. For example, if you want to see the progress of tar as it writes files into the archive, you can use the ‘--verbose’ option. In the beginning, you may find it useful to use ‘--verbose’ at all times; when you are more accustomed to tar, you will likely want to use it at certain times but not at others. We will use ‘--verbose’ at times to help make something clear, and we will give many examples both using and not using ‘--verbose’ to show the differences.

Each instance of ‘--verbose’ on the command line increases the verbosity level by one, so if you need more details on the output, specify it twice.

When reading archives (‘--list’, ‘--extract’, ‘--diff’), tar by default prints only the names of the members being extracted. Using ‘--verbose’ will show a full, ls style member listing.

In contrast, when writing archives (‘--create’, ‘--append’, ‘--update’), tar does not print file names by default. So, a single ‘--verbose’ option shows the file names being added to the archive, while two ‘--verbose’ options enable the full listing.

For example, to create an archive in verbose mode:

     $ tar -cvf afiles.tar apple angst aspic
     apple

angst
aspic

Creating the same archive with the verbosity level 2 could give:

$ tar -cvvf afiles.tar apple angst aspic
-rw-r--r-- gray/staff    62373 2006-06-09 12:06 apple
-rw-r--r-- gray/staff    11481 2006-06-09 12:06 angst
-rw-r--r-- gray/staff    23152 2006-06-09 12:06 aspic

This works equally well using short or long forms of options. Using long forms, you would simply write out the mnemonic form of the option twice, like this:

$ tar --create --verbose --verbose ...

Note that you must double the hyphens properly each time.

Later in the tutorial, we will give examples using ‘--verbose --verbose’. The full output consists of six fields:

• File type and permissions in symbolic form. These are displayed in the same format as the first column of ls -l output (see section “Verbose listing” in GNU file utilities).

• Owner name and group separated by a slash character. If these data are not available (for example, when listing a ‘v7’ format archive), numeric ID values are printed instead.

• Size of the file, in bytes.

• File modification date in ISO 8601 format.

• File modification time.

• File name. If the name contains any special characters (white space, newlines, etc.) these are displayed in an unambiguous form using so called quoting style. For the detailed discussion of available styles and on how to use them, see Section 6.6 [quoting styles], page 96. Depending on the file type, the name can be followed by some additional information, described in the following table:

‘-> link-name’

The file or archive member is a symbolic link and link-name is the name of file it links to.

‘link to link-name’

The file or archive member is a hard link and link-name is the name of file it links to.

‘--Long Link--’

The archive member is an old GNU format long link. You will normally not encounter this.

‘--Long Name--’

The archive member is an old GNU format long name. You will normally not encounter this.

‘--Volume Header--’

The archive member is a GNU volume header (see Section 9.6.2 [Tape Files], page 154).

‘--Continued at byte n--’

Encountered only at the beginning of a multi-volume archive (see Section 9.6 [Using Multiple Tapes], page 150). This archive member is a continuation from the previous volume. The number n gives the offset where the original file was split.

‘unknown file type c’

An archive member of unknown type. c is the type character from the archive header. If you encounter such a message, it means that either your archive contains proprietary member types GNU tar is not able to handle, or the archive is corrupted.

For example, here is an archive listing containing most of the special suffixes explained above:

V--------- 0/0          1536 2006-06-09 13:07 MyVolume--Volume Header--
-rw-r--r-- gray/staff 456783 2006-06-09 12:06 aspic--Continued at
byte 32456--
-rw-r--r-- gray/staff  62373 2006-06-09 12:06 apple
lrwxrwxrwx gray/staff      0 2006-06-09 13:01 angst -> apple
-rw-r--r-- gray/staff  35793 2006-06-09 12:06 blues
hrw-r--r-- gray/staff      0 2006-06-09 12:06 music link to blues

Getting Help: Using the ‘--help’ Option

‘--help’

The ‘--help’ option to tar prints out a very brief list of all operations and option available for the current version of tar available on your system.

2.6 How to Create Archives

(This message will disappear, once this node revised.) One of the basic operations of tar is ‘--create’ (‘-c’), which you use to

create a tar archive. We will explain ‘--create’ first because, in order to learn about the other operations, you will find it useful to have an archive available to practice on.

To make this easier, in this section you will first create a directory containing three files. Then, we will show you how to create an archive (inside the new directory). Both the directory, and the archive are specifically for you to practice on. The rest of this chapter and the next chapter will show many examples using this directory and the files you will create: some of those files may be other directories and other archives.

The three files you will archive in this example are called ‘blues’, ‘folk’, and ‘jazz’. The archive is called ‘collection.tar’.

This section will proceed slowly, detailing how to use ‘--create’ in verbose mode, and showing examples using both short and long forms. In the rest of the tutorial, and in the examples in the next chapter, we will proceed at a slightly quicker pace. This section moves more slowly to allow beginning users to understand how tar works.

2.6.1 Preparing a Practice Directory for Examples

To follow along with this and future examples, create a new directory called ‘practice’ containing files called ‘blues’, ‘folk’ and ‘jazz’. The files can contain any information you like: ideally, they should contain information which relates to their names, and be of different lengths. Our examples assume that ‘practice’ is a subdirectory of your home directory.

Now cd to the directory named ‘practice’; ‘practice’ is now your working directory. (Please note: Although the full path name of this directory is ‘/homedir/practice’, in our examples we will refer to this directory as ‘practice’; the homedir is presumed.

In general, you should check that the files to be archived exist where you think they do (in the working directory) by running ls. Because you just created the directory and the files and have changed to that directory, you probably don’t need to do that this time.

It is very important to make sure there isn’t already a file in the working directory with the archive name you intend to use (in this case, ‘collection.tar’), or that you don’t care about its contents. Whenever you use ‘create’, tar will erase the current contents of the file named by

‘--file=archive-name’ (‘-f archive-name’) if it exists. tar will not tell

you if you are about to overwrite an archive unless you specify an option which does this (see Section 4.5 [backup], page 69, for the information on how to do so). To add files to an existing archive, you need to use a different option, such as ‘--append’ (‘-r’); see Section 4.2.2 [append], page 53 for information on how to do this.

2.6.2 Creating the Archive

To place the files ‘blues’, ‘folk’, and ‘jazz’ into an archive named ‘collection.tar’, use the following command:

     $ tar --create --file=collection.tar blues folk jazz

The order of the arguments is not very important, when using long option forms. You could also say:

     $ tar blues --create folk --file=collection.tar jazz

However, you can see that this order is harder to understand; this is why we will list the arguments in the order that makes the commands easiest to understand (and we encourage you to do the same when you use tar, to avoid errors).

Note that the sequence ‘--file=collection.tar’ is considered to be one argument. If you substituted any other string of characters for collection.tar, then that string would become the name of the archive file you create.

The order of the options becomes more important when you begin to use short forms. With short forms, if you type commands in the wrong order (even if you type them correctly in all other ways), you may end up with results you don’t expect. For this reason, it is a good idea to get into the habit of typing options in the order that makes inherent sense. See Section 2.6.4 [short create], page 13, for more information on this.

In this example, you type the command as shown above: ‘--create’ is the operation which creates the new archive (‘collection.tar’), and ‘--file’ is the option which lets you give it the name you chose. The files, ‘blues’, ‘folk’, and ‘jazz’, are now members of the archive, ‘collection.tar’ (they are file name arguments to the ‘--create’ operation. See Chapter 6 [Choosing], page 87, for the detailed discussion on these.) Now that they are in the archive, they are called archive members, not files. (see Section 1.2 [Definitions], page 1).

When you create an archive, you must specify which files you want placed in the archive. If you do not specify any archive members, GNU tar will complain.

If you now list the contents of the working directory (ls), you will find the archive file listed as well as the files you saw previously:

     blues   folk   jazz   collection.tar

Creating the archive ‘collection.tar’ did not destroy the copies of the files in the directory.

Keep in mind that if you don’t indicate an operation, tar will not run and will prompt you for one. If you don’t name any files, tar will complain. You must have write access to the working directory, or else you will not be able to create an archive in that directory.

Caution: Do not attempt to use ‘--create’ (‘-c’) to add files to an existing archive; it will delete the archive and write a new one. Use ‘--append’ (‘-r’) instead. See Section 4.2.2 [append], page 53.

2.6.3 Running ‘--create’ with ‘--verbose’

If you include the ‘--verbose’ (‘-v’) option on the command line, tar will list the files it is acting on as it is working. In verbose mode, the create example above would appear as:

     $ tar --create --verbose --file=collection.tar blues folk jazz
     blues
     folk
     jazz

This example is just like the example we showed which did not use ‘--verbose’, except that tar generated the remaining lines (note the different font styles).

In the rest of the examples in this chapter, we will frequently use verbose mode so we can show actions or tar responses that you would otherwise not see, and which are important for you to understand.

2.6.4 Short Forms with ‘create’

As we said before, the ‘--create’ (‘-c’) operation is one of the most basic uses of tar, and you will use it countless times. Eventually, you will probably want to use abbreviated (or “short”) forms of options. A full discussion of the three different forms that options can take appears in Section 3.3 [Styles], page 24; for now, here is what the previous example (including the ‘--verbose’ (‘-v’) option) looks like using short option forms:

     $ tar -cvf collection.tar blues folk jazz
     blues
     folk
     jazz

As you can see, the system responds the same no matter whether you use long or short option forms.

One difference between using short and long option forms is that, although the exact placement of arguments following options is no more specific when using short forms, it is easier to become confused and make a mistake when using short forms. For example, suppose you attempted the above example in the following way:

     $ tar -cfv collection.tar blues folk jazz

In this case, tar will make an archive file called ‘v’, containing the files ‘blues’, ‘folk’, and ‘jazz’, because the ‘v’ is the closest “file name” to the ‘-f’ option, and is thus taken to be the chosen archive file name. tar will try to add a file called ‘collection.tar’ to the ‘v’ archive file; if the file ‘collection.tar’ did not already exist, tar will report an error indicating that this file does not exist. If the file ‘collection.tar’ does already exist (e.g., from a previous command you may have run), then tar will add this file to the archive. Because the ‘-v’ option did not get registered, tar will not run under ‘verbose’ mode, and will not report its progress.

The end result is that you may be quite confused about what happened, and possibly overwrite a file. To illustrate this further, we will show you how an example we showed previously would look using short forms.

This example,

     $ tar blues --create folk --file=collection.tar jazz

is confusing as it is. When shown using short forms, however, it becomes much more so:

     $ tar blues -c folk -f collection.tar jazz

It would be very easy to put the wrong string of characters immediately following the ‘-f’, but doing that could sacrifice valuable data.

For this reason, we recommend that you pay very careful attention to the order of options and placement of file and archive names, especially when using short option forms. Not having the option name written out mnemonically can affect how well you remember which option does what, and therefore where different names have to be placed.

2.6.5 Archiving Directories

You can archive a directory by specifying its directory name as a file name argument to tar. The files in the directory will be archived relative to the working directory, and the directory will be re-created along with its contents when the archive is extracted.

To archive a directory, first move to its superior directory. If you have followed the previous instructions in this tutorial, you should type:

     $ cd ..
     $

This will put you into the directory which contains ‘practice’, i.e., your home directory. Once in the superior directory, you can specify the subdirectory, ‘practice’, as a file name argument. To store ‘practice’ in the new archive file ‘music.tar’, type:

     $ tar --create --verbose --file=music.tar practice

tar should output:

     practice/
     practice/blues
     practice/folk
     practice/jazz
     practice/collection.tar

Note that the archive thus created is not in the subdirectory ‘practice’, but rather in the current working directory—the directory from which tar was invoked. Before trying to archive a directory from its superior directory, you should make sure you have write access to the superior directory itself, not only the directory you are trying archive with tar. For example, you will probably not be able to store your home directory in an archive by invoking tar from the root directory; See Section 6.10.2 [absolute], page 106. (Note also that ‘collection.tar’, the original archive file, has itself been archived. tar will accept any file as a file to be archived, regardless of its content. When ‘music.tar’ is extracted, the archive file ‘collection.tar’ will be re-written into the file system).

If you give tar a command such as

     $ tar --create --file=foo.tar .
tar will report ‘tar: ./foo.tar is the archive; not dumped’. This hap-

pens because tar creates the archive ‘foo.tar’ in the current directory before putting any files into it. Then, when tar attempts to add all the files in the directory ‘.’ to the archive, it notices that the file ‘./foo.tar’ is the same as the archive ‘foo.tar’, and skips it. (It makes no sense to put an archive into itself.) GNU tar will continue in this case, and create the archive normally, except for the exclusion of that one file. (Please note: Other implementations of tar may not be so clever; they will enter an infinite loop when this happens, so you should not depend on this behavior unless you are certain you are running GNU tar. In general, it is wise to always place the archive outside of the directory being dumped.

2.7 How to List Archives

Frequently, you will find yourself wanting to determine exactly what a particular archive contains. You can use the ‘--list’ (‘-t’) operation to get the member names as they currently appear in the archive, as well as various attributes of the files at the time they were archived. For example, you can examine the archive ‘collection.tar’ that you created in the last section with the command,

     $ tar --list --file=collection.tar

The output of tar would then be:

     blues
     folk
     jazz

The archive ‘bfiles.tar’ would list as follows:

     ./birds
     baboon
     ./box
Be sure to use a ‘--file=archive-name’ (‘-f archive-name’) option just

as with ‘--create’ (‘-c’) to specify the name of the archive.

If you use the ‘--verbose’ (‘-v’) option with ‘--list’, then tar will print out a listing reminiscent of ‘ls -l’, showing owner, file size, and so forth. This output is described in detail in [verbose member listing], page 9.

If you had used ‘--verbose’ (‘-v’) mode, the example above would look like:

     $ tar --list --verbose --file=collection.tar folk
     -rw-r--r-- myself user 62 1990-05-23 10:55 folk

It is important to notice that the output of tar --list --verbose does not necessarily match that produced by tar --create --verbose while creating the archive. It is because GNU tar, unless told explicitly not to do so, removes some directory prefixes from file names before storing them in the archive (See Section 6.10.2 [absolute], page 106, for more information). In other words, in verbose mode GNU tar shows file names when creating an archive and member names when listing it. Consider this example:

     $ tar cfv archive /etc/mail
     tar: Removing leading `/' from member names
     /etc/mail/
     /etc/mail/sendmail.cf
     /etc/mail/aliases
     $ tar tf archive
     etc/mail/
     etc/mail/sendmail.cf
     etc/mail/aliases

This default behavior can sometimes be inconvenient. You can force GNU tar show member names when creating archive by supplying

‘--show-stored-names’ option.
‘--show-stored-names’

Print member (as opposed to file) names when creating the archive.

You can specify one or more individual member names as arguments when using ‘list’. In this case, tar will only list the names of members you

identify. For example, tar --list --file=afiles.tar apple would only

print ‘apple’.

Because tar preserves paths, file names must be specified as they appear in the archive (i.e., relative to the directory from which the archive was created). Therefore, it is essential when specifying member names to tar that you give the exact member names. For example,

tar --list --file=bfiles.tar birds would produce an error message

something like ‘tar: birds: Not found in archive’, because there is no member named ‘birds’, only one named ‘./birds’. While the names ‘birds’ and ‘./birds’ name the same file, member names by default are compared verbatim.

   However, tar --list --file=bfiles.tar baboon would respond with

‘baboon’, because this exact member name is in the archive file ‘bfiles.tar’. If you are not sure of the exact file name, use globbing patterns, for example:

     $ tar --list --file=bfiles.tar --wildcards '*b*'

will list all members whose name contains ‘b’. See Section 6.5 [wildcards], page 93, for a detailed discussion of globbing patterns and related tar command line options.

Listing the Contents of a Stored Directory

To get information about the contents of an archived directory, use the directory name as a file name argument in conjunction with ‘--list’ (‘-t’). To find out file attributes, include the ‘--verbose’ (‘-v’) option.

For example, to find out about files in the directory ‘practice’, in the archive file ‘music.tar’, type:

     $ tar --list --verbose --file=music.tar practice

tar responds:

     drwxrwxrwx myself user 0 1990-05-31 21:49 practice/
     -rw-r--r-- myself user 42 1990-05-21 13:29 practice/blues
     -rw-r--r-- myself user 62 1990-05-23 10:55 practice/folk
     -rw-r--r-- myself user 40 1990-05-21 13:30 practice/jazz
     -rw-r--r-- myself user 10240 1990-05-31 21:49 practice/collection.tar

When you use a directory name as a file name argument, tar acts on all the files (including sub-directories) in that directory.

2.8 How to Extract Members from an Archive

(This message will disappear, once this node revised.)

Creating an archive is only half the job—there is no point in storing files in an archive if you can’t retrieve them. The act of retrieving members from an archive so they can be used and manipulated as unarchived files again is called extraction. To extract files from an archive, use the ‘--extract’ (‘--get’ or ‘-x’) operation. As with ‘--create’, specify the name of the archive with ‘--file’ (‘-f’) option. Extracting an archive does not modify the archive in any way; you can extract it multiple times if you want or need to.

Using ‘--extract’, you can extract an entire archive, or specific files. The files can be directories containing other files, or not. As with ‘--create’ (‘-c’) and ‘--list’ (‘-t’), you may use the short or the long form of the operation without affecting the performance.

2.8.1 Extracting an Entire Archive

To extract an entire archive, specify the archive file name only, with no individual file names as arguments. For example,

     $ tar -xvf collection.tar

produces this:

     -rw-r--r-- me user     28 1996-10-18 16:31 jazz
     -rw-r--r-- me user     21 1996-09-23 16:44 blues
     -rw-r--r-- me user     20 1996-09-23 16:44 folk

2.8.2 Extracting Specific Files

To extract specific archive members, give their exact member names as arguments, as printed by ‘--list’ (‘-t’). If you had mistakenly deleted one of the files you had placed in the archive ‘collection.tar’ earlier (say, ‘blues’), you can extract it from the archive without changing the archive’s structure. Its contents will be identical to the original file ‘blues’ that you deleted.

First, make sure you are in the ‘practice’ directory, and list the files in the directory. Now, delete the file, ‘blues’, and list the files in the directory again.

You can now extract the member ‘blues’ from the archive file

‘collection.tar’ like this:
     $ tar --extract --file=collection.tar blues

If you list the files in the directory again, you will see that the file ‘blues’ has been restored, with its original permissions, data modification times, and owner.^{1} (These parameters will be identical to those which the file had when you originally placed it in the archive; any changes you may have made before deleting the file from the file system, however, will not have been made to the archive member.) The archive file, ‘collection.tar’, is the same as it was before you extracted ‘blues’. You can confirm this by running tar with ‘--list’ (‘-t’).

Remember that as with other operations, specifying the exact member

name is important. tar --extract --file=bfiles.tar birds will fail, be-

cause there is no member named ‘birds’. To extract the member named

‘./birds’, you must specify tar --extract --file=bfiles.tar ./birds.

If you don’t remember the exact member names, use ‘--list’ (‘-t’) option (see Section 2.7 [list], page 15). You can also extract those members that match a specific globbing pattern. For example, to extract from ‘bfiles.tar’ all files that begin with ‘b’, no matter their directory prefix, you could type:

     $ tar -x -f bfiles.tar --wildcards --no-anchored 'b*'

Here, ‘--wildcards’ instructs tar to treat command line arguments as globbing patterns and ‘--no-anchored’ informs it that the patterns apply to member names after any ‘/’ delimiter. The use of globbing patterns is discussed in detail in See Section 6.5 [wildcards], page 93.

You can extract a file to standard output by combining the above options with the ‘--to-stdout’ (‘-O’) option (see [Writing to Standard Output], page 66).

If you give the ‘--verbose’ option, then ‘--extract’ will print the names of the archive members as it extracts them.

2.8.3 Extracting Files that are Directories

Extracting directories which are members of an archive is similar to extracting other files. The main difference to be aware of is that if the extracted directory has the same name as any directory already in the working directory, then files in the extracted directory will be placed into the directory of the same name. Likewise, if there are files in the pre-existing directory with the same names as the members which you extract, the files from the extracted archive will replace the files already in the working directory (and possible subdirectories). This will happen regardless of whether or not the files in the working directory were more recent than those extracted (there exist, however, special options that alter this behavior see Section 4.4.2 [Writing], page 62).

1 This is only accidentally true, but not in general. Whereas modification times are

always restored, in most cases, one has to be root for restoring the owner, and use a

special option for restoring permissions. Here, it just happens that the restoring user

is also the owner of the archived members, and that the current umask is compatible

with original permissions.

However, if a file was stored with a directory name as part of its file name, and that directory does not exist under the working directory when the file is extracted, tar will create the directory.

We can demonstrate how to use ‘--extract’ to extract a directory file with an example. Change to the ‘practice’ directory if you weren’t there, and remove the files ‘folk’ and ‘jazz’. Then, go back to the parent directory and extract the archive ‘music.tar’. You may either extract the entire archive, or you may extract only the files you just deleted. To extract the entire archive, don’t give any file names as arguments after the archive name ‘music.tar’. To extract only the files you deleted, use the following command:

     $ tar -xvf music.tar practice/folk practice/jazz
     practice/folk
     practice/jazz

If you were to specify two ‘--verbose’ (‘-v’) options, tar would have displayed more detail about the extracted files, as shown in the example below:

     $ tar -xvvf music.tar practice/folk practice/jazz
     -rw-r--r-- me user     28 1996-10-18 16:31 practice/jazz
     -rw-r--r-- me user     20 1996-09-23 16:44 practice/folk

Because you created the directory with ‘practice’ as part of the file names of each of the files by archiving the ‘practice’ directory as ‘practice’, you must give ‘practice’ as part of the file names when you extract those files from the archive.

2.8.4 Extracting Archives from Untrusted Sources

Extracting files from archives can overwrite files that already exist. If you receive an archive from an untrusted source, you should make a new directory and extract into that directory, so that you don’t have to worry about the extraction overwriting one of your existing files. For example, if ‘untrusted.tar’ came from somewhere else on the Internet, and you don’t necessarily trust its contents, you can extract it as follows:

     $ mkdir newdir
     $ cd newdir
     $ tar -xvf ../untrusted.tar

It is also a good practice to examine contents of the archive before extracting it, using ‘--list’ (‘-t’) option, possibly combined with ‘--verbose’ (‘-v’).

2.8.5 Commands That Will Fail

Here are some sample commands you might try which will not work, and why they won’t work.

If you try to use this command,

     $ tar -xvf music.tar folk jazz

you will get the following response:

     tar: folk: Not found in archive

     tar: jazz: Not found in archive
     $

This is because these files were not originally in the parent directory ‘..’, where the archive is located; they were in the ‘practice’ directory, and their file names reflect this:

     $ tar -tvf music.tar
     practice/folk
     practice/jazz
     practice/rock

Likewise, if you try to use this command,

     $ tar -tvf music.tar folk jazz

you would get a similar response. Members with those names are not in the archive. You must use the correct member names, or wildcards, in order to extract the files from the archive.

If you have forgotten the correct names of the files in the archive, use tar --list --verbose to list them correctly.

2.9 Going Further Ahead in this Manual

3 Invoking GNU tar

(This message will disappear, once this node revised.)

This chapter is about how one invokes the GNU tar command, from the command synopsis (see Section 3.1 [Synopsis], page 21). There are numerous options, and many styles for writing them. One mandatory option specifies the operation tar should perform (see Section 3.4.1 [Operation Summary], page 28), other options are meant to detail how this operation should be performed (see Section 3.4.2 [Option Summary], page 29). Non-option arguments are not always interpreted the same way, depending on what the operation is.

You will find in this chapter everything about option styles and rules for writing them (see Section 3.3 [Styles], page 24). On the other hand, operations and options are fully described elsewhere, in other chapters. Here, you will find only synthetic descriptions for operations and options, together with pointers to other parts of the tar manual.

Some options are so special they are fully described right in this chapter. They have the effect of inhibiting the normal operation of tar or else, they globally alter the amount of feedback the user receives about what is going on. These are the ‘--help’ and ‘--version’ (see Section 3.5 [help], page 44), ‘--verbose’ (see Section 3.7 [verbose], page 46) and ‘--interactive’ options (see Section 3.8 [interactive], page 48).

3.1 General Synopsis of tar

The GNU tar program is invoked as either one of:

     tar option... [name]...
     tar letter... [argument]... [option]... [name]...

The second form is for when old options are being used.

You can use tar to store files in an archive, to extract them from an archive, and to do other types of archive manipulation. The primary argument to tar, which is called the operation, specifies which action to take. The other arguments to tar are either options, which change the way tar performs an operation, or file names or archive members, which specify the files or members tar is to act on.

You can actually type in arguments in any order, even if in this manual the options always precede the other arguments, to make examples easier to understand. Further, the option stating the main operation mode (the tar main command) is usually given first.

Each name in the synopsis above is interpreted as an archive member name when the main command is one of ‘--compare’ (‘--diff’, ‘-d’),

‘--delete’, ‘--extract’ (‘--get’, ‘-x’), ‘--list’ (‘-t’) or ‘--update’ (‘-u’).

When naming archive members, you must give the exact name of the member in the archive, as it is printed by ‘--list’. For ‘--append’ (‘-r’) and ‘--create’ (‘-c’), these name arguments specify the names of either files or directory hierarchies to place in the archive. These files or hierarchies should already exist in the file system, prior to the execution of the tar command.

tar interprets relative file names as being relative to the working directory. tar will make all file names relative (by removing leading slashes when archiving or restoring files), unless you specify otherwise (using the ‘--absolute-names’ option). See Section 6.10.2 [absolute], page 106, for more information about ‘--absolute-names’.

If you give the name of a directory as either a file name or a member name, then tar acts recursively on all the files and directories beneath that directory. For example, the name ‘/’ identifies all the files in the file system

to tar.

The distinction between file names and archive member names is especially important when shell globbing is used, and sometimes a source of confusion for newcomers. See Section 6.5 [wildcards], page 93, for more information about globbing. The problem is that shells may only glob using existing files in the file system. Only tar itself may glob on archive members, so when needed, you must ensure that wildcard characters reach tar without being interpreted by the shell first. Using a backslash before ‘*’ or ‘?’, or putting the whole argument between quotes, is usually sufficient for this.

Even if names are often specified on the command line, they can also be read from a text file in the file system, using the ‘--files-from=file-of-

names’ (‘-T file-of-names’) option.

If you don’t use any file name arguments, ‘--append’ (‘-r’), ‘--delete’

and ‘--concatenate’ (‘--catenate’, ‘-A’) will do nothing, while ‘--create’

(‘-c’) will usually yield a diagnostic and inhibit tar execution. The other

operations of tar (‘--list’, ‘--extract’, ‘--compare’, and ‘--update’) will

act on the entire contents of the archive.

Besides successful exits, GNU tar may fail for many reasons. Some reasons correspond to bad usage, that is, when the tar command is improperly written. Errors may be encountered later, while encountering an error processing the archive or the files. Some errors are recoverable, in which case the failure is delayed until tar has completed all its work. Some errors are such that it would not meaningful, or at least risky, to continue processing: tar then aborts processing immediately. All abnormal exits, whether immediate or delayed, should always be clearly diagnosed on stderr, after a line stating the nature of the error.

Possible exit codes of GNU tar are summarized in the following table:

0  ‘Successful termination’.

1 ‘Some files differ’. If tar was invoked with ‘--compare’ (‘--diff’, ‘-d’) command line option, this means that some files in the archive differ from their disk counterparts (see Section 4.2.6 [compare], page 59). If tar was given ‘--create’, ‘--append’ or ‘--update’ option, this exit code means that some files were changed while being archived and so the resulting archive does not contain the exact copy of the file set.

2 ‘Fatal error’. This means that some fatal, unrecoverable error occurred.

If tar has invoked a subprocess and that subprocess exited with a nonzero exit code, tar exits with that code as well. This can happen, for example, if tar was given some compression option (see Section 8.1.1 [gzip], page 118) and the external compressor program failed. Another example is rmt failure during backup to the remote device (see Section 9.2 [Remote Tape Server], page 139).

3.2 Using tar Options

GNU tar has a total of eight operating modes which allow you to perform a variety of tasks. You are required to choose one operating mode each time you employ the tar program by specifying one, and only one operation as an argument to the tar command (two lists of four operations each may be found at Section 2.4 [frequent operations], page 7 and Section 4.2.1 [Operations], page 52). Depending on circumstances, you may also wish to customize how the chosen operating mode behaves. For example, you may wish to change the way the output looks, or the format of the files that you wish to archive may require you to do something special in order to make the archive look right.

You can customize and control tar’s performance by running tar with one or more options (such as ‘--verbose’ (‘-v’), which we used in the tutorial). As we said in the tutorial, options are arguments to tar which are (as their name suggests) optional. Depending on the operating mode, you may specify one or more options. Different options will have different effects, but in general they all change details of the operation, such as archive format, archive name, or level of user interaction. Some options make sense with all operating modes, while others are meaningful only with particular modes. You will likely use some options frequently, while you will only use others infrequently, or not at all. (A full list of options is available in see Section 3.4 [All Options], page 28.)

The TAR_OPTIONS environment variable specifies default options to be placed in front of any explicit options. For example, if TAR_OPTIONS is ‘-v --unlink-first’, tar behaves as if the two options ‘-v’ and ‘--unlink-first’ had been specified before any explicit options. Option specifications are separated by whitespace. A backslash escapes the next character, so it can be used to specify an option containing whitespace or a backslash.

Note that tar options are case sensitive. For example, the options ‘-T’ and ‘-t’ are different; the first requires an argument for stating the name of a file providing a list of names, while the second does not require an argument and is another way to write ‘--list’ (‘-t’).

In addition to the eight operations, there are many options to tar, and three different styles for writing both: long (mnemonic) form, short form, and old style. These styles are discussed below. Both the options and the operations can be written in any of these three styles.

3.3 The Three Option Styles

There are three styles for writing operations and options to the command line invoking tar. The different styles were developed at different times during the history of tar. These styles will be presented below, from the most recent to the oldest.

Some options must take an argument. (For example, ‘--file’ (‘-f’)) takes the name of an archive file as an argument. If you do not supply an archive file name, tar will use a default, but this can be confusing; thus, we recommend that you always supply a specific archive file name.) Where you place the arguments generally depends on which style of options you choose. We will detail specific information relevant to each option style in the sections on the different option styles, below. The differences are subtle, yet can often be very important; incorrect option placement can cause you to overwrite a number of important files. We urge you to note these differences, and only use the option style(s) which makes the most sense to you until you feel comfortable with the others.

Some options may take an argument. Such options may have at most long and short forms, they do not have old style equivalent. The rules for specifying an argument for such options are stricter than those for specifying mandatory arguments. Please, pay special attention to them.

3.3.1 Long Option Style

Each option has at least one long (or mnemonic) name starting with two dashes in a row, e.g., ‘--list’. The long names are more clear than their corresponding short or old names. It sometimes happens that a single long option has many different names which are synonymous, such as ‘--compare’ and ‘--diff’. In addition, long option names can be given unique abbreviations. For example, ‘--cre’ can be used in place of ‘--create’ because there is no other long option which begins with ‘cre’. (One way to find this out is by trying it and seeing what happens; if a particular abbreviation could represent more than one option, tar will tell you that that abbreviation is ambiguous and you’ll know that that abbreviation won’t work. You may also choose to run ‘tar --help’ to see a list of options. Be aware that if you run tar with a unique abbreviation for the long name of an option you didn’t want to use, you are stuck; tar will perform the command as ordered.)

Long options are meant to be obvious and easy to remember, and their meanings are generally easier to discern than those of their corresponding short options (see below). For example:

     $ tar --create --verbose --blocking-factor=20 --file=/dev/rmt0

gives a fairly good set of hints about what the command does, even for those not fully acquainted with tar.

Long options which require arguments take those arguments immediately following the option name. There are two ways of specifying a mandatory argument. It can be separated from the option name either by an equal sign, or by any amount of white space characters. For example, the ‘--file’ option (which tells the name of the tar archive) is given a file such as ‘archive.tar’ as argument by using any of the following notations: ‘--file=archive.tar’

or ‘--file archive.tar’.

In contrast, optional arguments must always be introduced using an equal sign. For example, the ‘--backup’ option takes an optional argument specifying backup type. It must be used as ‘--backup=backup-type’.

3.3.2 Short Option Style

Most options also have a short option name. Short options start with a single dash, and are followed by a single character, e.g., ‘-t’ (which is equivalent to ‘--list’). The forms are absolutely identical in function; they are interchangeable.

The short option names are faster to type than long option names.

Short options which require arguments take their arguments immediately following the option, usually separated by white space. It is also possible to stick the argument right after the short option name, using no intervening space. For example, you might write ‘-f archive.tar’ or ‘-farchive.tar’

instead of using ‘--file=archive.tar’. Both ‘--file=archive-name’ and

‘-f archive-name’ denote the option which indicates a specific archive, here

named ‘archive.tar’.

Short options which take optional arguments take their arguments immediately following the option letter, without any intervening white space characters.

Short options’ letters may be clumped together, but you are not required to do this (as compared to old options; see below). When short options are clumped as a set, use one (single) dash for them all, e.g., ‘tar -cvf’. Only the last option in such a set is allowed to have an argument^{1}.

When the options are separated, the argument for each option which requires an argument directly follows that option, as is usual for Unix programs. For example:

     $ tar -c -v -b 20 -f /dev/rmt0

If you reorder short options’ locations, be sure to move any arguments that belong to them. If you do not move the arguments properly, you may end up overwriting files.

1 Clustering many options, the last of which has an argument, is a rather opaque way to

write options. Some wonder if GNU getopt should not even be made helpful enough

for considering such usages as invalid.

3.3.3 Old Option Style

(This message will disappear, once this node revised.)

Like short options, old options are single letters. However, old options must be written together as a single clumped set, without spaces separating them or dashes preceding them^{2}. This set of letters must be the first to appear on the command line, after the tar program name and some white space; old options cannot appear anywhere else. The letter of an old option is exactly the same letter as the corresponding short option. For example, the old option ‘t’ is the same as the short option ‘-t’, and consequently, the same as the long option ‘--list’. So for example, the command ‘tar cv’ specifies the option ‘-v’ in addition to the operation ‘-c’.

When options that need arguments are given together with the command, all the associated arguments follow, in the same order as the options. Thus, the example given previously could also be written in the old style as follows:

     $ tar cvbf 20 /dev/rmt0

Here, ‘20’ is the argument of ‘-b’ and ‘/dev/rmt0’ is the argument of ‘-f’.

On the other hand, this old style syntax makes it difficult to match option letters with their corresponding arguments, and is often confusing. In the command ‘tar cvbf 20 /dev/rmt0’, for example, ‘20’ is the argument for ‘-b’, ‘/dev/rmt0’ is the argument for ‘-f’, and ‘-v’ does not have a corresponding argument. Even using short options like in ‘tar -c -v -b 20 -f /dev/rmt0’ is clearer, putting all arguments next to the option they pertain to.

If you want to reorder the letters in the old option argument, be sure to reorder any corresponding argument appropriately.

This old way of writing tar options can surprise even experienced users. For example, the two commands:

     tar cfz archive.tar.gz file
     tar -cfz archive.tar.gz file

are quite different. The first example uses ‘archive.tar.gz’ as the value for option ‘f’ and recognizes the option ‘z’. The second example, however, uses ‘z’ as the value for option ‘f’ — probably not what was intended.

Old options are kept for compatibility with old versions of tar.

This second example could be corrected in many ways, among which the following are equivalent:

     tar -czf archive.tar.gz file
     tar -cf archive.tar.gz -z file
     tar cf archive.tar.gz -z file

As far as we know, all tar programs, GNU and non-GNU, support old options. GNU tar supports them not only for historical reasons, but also because many people are used to them. For compatibility with Unix tar, the first argument is always treated as containing command and option letters even if it doesn’t start with ‘-’. Thus, ‘tar c’ is equivalent to ‘tar -c’: both of them specify the ‘--create’ (‘-c’) command to create an archive.

2 Beware that if you precede options with a dash, you are announcing the short option

style instead of the old option style; short options are decoded differently.

3.3.4 Mixing Option Styles

All three styles may be intermixed in a single tar command, so long as the rules for each style are fully respected^{3}. Old style options and either of the modern styles of options may be mixed within a single tar command. However, old style options must be introduced as the first arguments only, following the rule for old options (old options must appear directly after the tar command and some white space). Modern options may be given only after all arguments to the old options have been collected. If this rule is not respected, a modern option might be falsely interpreted as the value of the argument to one of the old style options.

For example, all the following commands are wholly equivalent, and illustrate the many combinations and orderings of option styles.

     tar --create --file=archive.tar
     tar --create -f archive.tar
     tar --create -farchive.tar
     tar --file=archive.tar --create
     tar --file=archive.tar -c
     tar -c --file=archive.tar
     tar -c -f archive.tar
     tar -c -farchive.tar
     tar -cf archive.tar
     tar -cfarchive.tar
     tar -f archive.tar --create
     tar -f archive.tar -c
     tar -farchive.tar --create
     tar -farchive.tar -c
     tar c --file=archive.tar
     tar c -f archive.tar
     tar c -farchive.tar
     tar cf archive.tar
     tar f archive.tar --create
     tar f archive.tar -c
     tar fc archive.tar

On the other hand, the following commands are not equivalent to the previous set:

     tar -f -c archive.tar
     tar -fc archive.tar
     tar -fcarchive.tar
     tar -farchive.tarc
     tar cfarchive.tar

3 Before GNU tar version 1.11.6, a bug prevented intermixing old style options with long

options in some cases.

These last examples mean something completely different from what the user intended (judging based on the example in the previous set which uses long options, whose intent is therefore very clear). The first four specify that the tar archive would be a file named ‘-c’, ‘c’, ‘carchive.tar’ or ‘archive.tarc’, respectively. The first two examples also specify a single non-option, name argument having the value ‘archive.tar’. The last example contains only old style option letters (repeating option ‘c’ twice), not all of which are meaningful (eg., ‘.’, ‘h’, or ‘i’), with no argument value.

3.4 All tar Options

The coming manual sections contain an alphabetical listing of all tar operations and options, with brief descriptions and cross references to more in-depth explanations in the body of the manual. They also contain an alphabetically arranged table of the short option forms with their corresponding long option. You can use this table as a reference for deciphering tar commands in scripts.

3.4.1 Operations

‘--append’
‘-r’

Appends files to the end of the archive. See Section 4.2.2 [append], page 53.

‘--catenate’
‘-A’

Same as ‘--concatenate’. See Section 4.2.4 [concatenate], page 57.

‘--compare’
‘-d’

Compares archive members with their counterparts in the file system, and reports differences in file size, mode, owner, modification date and contents. See Section 4.2.6 [compare], page 59.

‘--concatenate’
‘-A’

Appends other tar archives to the end of the archive. See Section 4.2.4 [concatenate], page 57.

‘--create’
‘-c’

Creates a new tar archive. See Section 2.6 [create], page 10.

‘--delete’

Deletes members from the archive. Don’t try this on a archive on a tape! See Section 4.2.5 [delete], page 58.

‘--diff’
‘-d’

Same ‘--compare’. See Section 4.2.6 [compare], page 59.

‘--extract’
‘-x’

Extracts members from the archive into the file system. See Section 2.8 [extract], page 17.

‘--get’
‘-x’

Same as ‘--extract’. See Section 2.8 [extract], page 17.

‘--list’
‘-t’

Lists the members in an archive. See Section 2.7 [list], page 15.

‘--update’
‘-u’

Adds files to the end of the archive, but only if they are newer than their counterparts already in the archive, or if they do not already exist in the archive. See Section 4.2.3 [update], page 56.

3.4.2 tar Options

‘--absolute-names’
‘-P’

Normally when creating an archive, tar strips an initial ‘/’ from member names. This option disables that behavior. See Section 6.10.2 [absolute], page 106.

‘--after-date’

(See ‘--newer’, see Section 6.8 [after], page 102)

‘--anchored’

A pattern must match an initial subsequence of the name’s components. See [controlling pattern-matching], page 94.

‘--atime-preserve’
‘--atime-preserve=replace’
‘--atime-preserve=system’

Attempt to preserve the access time of files when reading them. This option currently is effective only on files that you own, unless you have superuser privileges. ‘--atime-preserve=replace’ remembers the access time of a file before reading it, and then restores the access time afterwards. This may cause problems if other programs are reading the file at the same time, as the times of their accesses will be lost. On most platforms restoring the access time also requires tar to restore the data modification time too, so this option may also cause problems if other programs are writing the file at the same time. (Tar attempts to detect this situation, but cannot do so reliably due to race conditions.) Worse, on most platforms restoring the access time also updates the status change time, which means that this option is incompatible with incremental backups. ‘--atime-preserve=system’ avoids changing time stamps on files, without interfering with time stamp updates caused by other programs, so it works better with incremental backups. However, it requires a special O_NOATIME option from the underlying operating and file system implementation, and it also requires that searching directories does not update their access times. As of this writing (November 2005) this works only with Linux, and only with Linux kernels 2.6.8 and later. Worse, there is currently no reliable way to know whether this feature actually works. Sometimes tar knows that it does not work, and if you

use ‘--atime-preserve=system’ then tar complains and exits

right away. But other times tar might think that the option works when it actually does not. Currently ‘--atime-preserve’ with no operand defaults to ‘--atime-preserve=replace’, but this may change in the future as support for ‘--atime-preserve=system’ improves. If your operating system does not support ‘--atime-preserve= system’, you might be able to preserve access times reliably by by using the mount command. For example, you can mount the file system read-only, or access the file system via a read-only loopback mount, or use the ‘noatime’ mount option available on some systems. However, mounting typically requires superuser privileges and can be a pain to manage.

‘--backup=backup-type’

Rather than deleting files from the file system, tar will back them up using simple or numbered backups, depending upon backup-type. See Section 4.5 [backup], page 69.

‘--block-number’
‘-R’

With this option present, tar prints error messages for read errors with the block number in the archive file. See [blocknumber], page 48.

‘--blocking-factor=blocking’
‘-b blocking’

Sets the blocking factor tar uses to blocking x 512 bytes per record. See Section 9.4.2 [Blocking Factor], page 143.

‘--bzip2’
‘-j’

This option tells tar to read or write archives through bzip2. See Section 8.1.1 [gzip], page 118.

‘--checkpoint[=number]’

This option directs tar to print periodic checkpoint messages as it reads through the archive. It is intended for when you want a visual indication that tar is still running, but don’t want to see ‘--verbose’ output. For a detailed description, see [Progress information], page 47.

‘--check-links’

‘-l’ If this option was given, tar will check the number of links dumped for each processed file. If this number does not match the total number of hard links for the file, a warning message will be output^{4}.

‘--compress’
‘--uncompress’
‘-Z’

tar will use the compress program when reading or writing the archive. This allows you to directly act on archives while saving space. See Section 8.1.1 [gzip], page 118.

‘--confirmation’

(See ‘--interactive’.) See Section 3.8 [interactive], page 48.

‘--delay-directory-restore’

Delay setting modification times and permissions of extracted directories until the end of extraction. See [Directory Modification Times and Permissions], page 65.

‘--dereference’
‘-h’

When creating a tar archive, tar will archive the file that a symbolic link points to, rather than archiving the symlink. See Section 8.3.2 [dereference], page 125.

‘--directory=dir’
‘-C dir’

When this option is specified, tar will change its current directory to dir before performing any operations. When this option is used during archive creation, it is order sensitive. See Section 6.10.1 [directory], page 105.

4 Earlier versions of GNU tar understood ‘-l’ as a synonym for ‘--one-file-system’.

The current semantics, which complies to UNIX98, was introduced with version 1.15.91. See Appendix A [Changes], page 159, for more information.

‘--exclude=pattern’

When performing operations, tar will skip files that match pattern. See Section 6.4 [exclude], page 91.

‘--exclude-from=file’
‘-X file’

Similar to ‘--exclude’, except tar will use the list of patterns in the file file. See Section 6.4 [exclude], page 91.

‘--exclude-caches’

Automatically excludes all directories containing a cache directory tag. See Section 6.4 [exclude], page 91.

‘--exclude-tag=file’

Exclude all directories, containing file named file. See Section 6.4 [exclude], page 91.

‘--file=archive’
‘-f archive’

tar will use the file archive as the tar archive it performs operations on, rather than tar’s compilation dependent default. See [file tutorial], page 7.

‘--files-from=file’
‘-T file’

tar will use the contents of file as a list of archive members or files to operate on, in addition to those specified on the command-line. See Section 6.3 [files], page 89.

‘--force-local’

Forces tar to interpret the filename given to ‘--file’ as a local file, even if it looks like a remote tape drive name. See [local and remote archives], page 88.

‘--format=format’
‘-H format’

Selects output archive format. Format may be one of the following:

‘v7’ Creates an archive that is compatible with Unix V7

          tar.

‘oldgnu’ Creates an archive that is compatible with GNU tar version 1.12 or earlier.

‘gnu’ Creates archive in GNU tar 1.13 format. Basically it is the same as ‘oldgnu’ with the only difference in the way it handles long numeric fields.

‘ustar’ Creates a POSIX.1-1988 compatible archive.

‘posix’ Creates a POSIX.1-2001 archive.

See Chapter 8 [Formats], page 117, for a detailed discussion of these formats.

‘--group=group’

Files added to the tar archive will have a group id of group, rather than the group from the source file. group is first decoded as a group symbolic name, but if this interpretation fails, it has to be a decimal numeric group ID. See Section 4.3.1 [override], page 59. Also see the comments for the ‘--owner=user’ option.

‘--gzip’
‘--gunzip’
‘--ungzip’
‘-z’

This option tells tar to read or write archives through gzip, allowing tar to directly operate on several kinds of compressed archives transparently. See Section 8.1.1 [gzip], page 118.

‘--help’
‘-?’

tar will print out a short message summarizing the operations and options to tar and exit. See Section 3.5 [help], page 44.

‘--ignore-case’

Ignore case when matching member or file names with patterns. See [controlling pattern-matching], page 94.

‘--ignore-command-error’

Ignore exit codes of subprocesses. See [Writing to an External Program], page 67.

‘--ignore-failed-read’

Do not exit unsuccessfully merely because an unreadable file was encountered. See Section 4.4.1 [Reading], page 61.

‘--ignore-zeros’
‘-i’

With this option, tar will ignore zeroed blocks in the archive, which normally signals EOF. See Section 4.4.1 [Reading], page 61.

‘--incremental’
‘-G’

Used to inform tar that it is working with an old GNU-format incremental backup archive. It is intended primarily for backwards compatibility only. See Section 5.2 [Incremental Dumps], page 74, for a detailed discussion of incremental archives.

‘--index-file=file’

Send verbose output to file instead of to standard output.

‘--info-script=script-file’
‘--new-volume-script=script-file’
‘-F script-file’

When tar is performing multi-tape backups, script-file is run at the end of each tape. If script-file exits with nonzero status, tar fails immediately. See [info-script], page 152, for a detailed discussion of script-file.

‘--interactive’
‘--confirmation’
‘-w’

Specifies that tar should ask the user for confirmation before performing potentially destructive options, such as overwriting files. See Section 3.8 [interactive], page 48.

‘--keep-newer-files’

Do not replace existing files that are newer than their archive copies when extracting files from an archive.

‘--keep-old-files’
‘-k’

Do not overwrite existing files when extracting files from an archive. See [Keep Old Files], page 64.

‘--label=name’
‘-V name’

When creating an archive, instructs tar to write name as a name record in the archive. When extracting or listing archives, tar will only operate on archives that have a label matching the pattern specified in name. See Section 9.6.2 [Tape Files], page 154.

‘--listed-incremental=snapshot-file’
‘-g snapshot-file’

During a ‘--create’ operation, specifies that the archive that tar creates is a new GNU-format incremental backup, using snapshot-file to determine which files to backup. With other operations, informs tar that the archive is in incremental format. See Section 5.2 [Incremental Dumps], page 74.

‘--mode=permissions’

When adding files to an archive, tar will use permissions for the archive members, rather than the permissions from the files. permissions can be specified either as an octal number or as symbolic permissions, like with chmod. See Section 4.3.1 [override], page 59.

‘--mtime=date’

When adding files to an archive, tar will use date as the modification time of members when creating archives, instead of their actual modification times. The value of date can be either a textual date representation (see Chapter 7 [Date input formats], page 109) or a name of the existing file, starting with ‘/’ or ‘.’. In the latter case, the modification time of that file is used. See Section 4.3.1 [override], page 59.

‘--multi-volume’
‘-M’

Informs tar that it should create or otherwise operate on a multi-volume tar archive. See Section 9.6 [Using Multiple Tapes], page 150.

‘--new-volume-script’

(see –info-script)

‘--seek’
‘-n’

Assume that the archive media supports seeks to arbitrary locations. Usually tar determines automatically whether the archive can be seeked or not. This option is intended for use in cases when such recognition fails.

‘--newer=date’
‘--after-date=date’
‘-N’

When creating an archive, tar will only add files that have changed since date. If date begins with ‘/’ or ‘.’, it is taken to be the name of a file whose data modification time specifies the date. See Section 6.8 [after], page 102.

‘--newer-mtime=date’

Like ‘--newer’, but add only files whose contents have changed (as opposed to just ‘--newer’, which will also back up files for which any status information has changed). See Section 6.8 [after], page 102.

‘--no-anchored’

An exclude pattern can match any subsequence of the name’s components. See [controlling pattern-matching], page 94.

‘--no-delay-directory-restore’

Setting modification times and permissions of extracted directories when all files from this directory has been extracted. This is the default. See [Directory Modification Times and Permissions], page 65.

‘--no-ignore-case’

Use case-sensitive matching. See [controlling pattern-matching], page 94.

‘--no-ignore-command-error’

Print warnings about subprocesses terminated with a non-zero exit code. See [Writing to an External Program], page 67.

‘--no-overwrite-dir’

Preserve metadata of existing directories when extracting files from an archive. See [Overwrite Old Files], page 63.

‘--no-quote-chars=string’

Remove characters listed in string from the list of quoted characters set by the previous ‘--quote-chars’ option (see Section 6.6 [quoting styles], page 96).

‘--no-recursion’

With this option, tar will not recurse into directories. See Section 6.9 [recurse], page 104.

‘--no-same-owner’
‘-o’

When extracting an archive, do not attempt to preserve the owner specified in the tar archive. This the default behavior for ordinary users.

‘--no-same-permissions’

When extracting an archive, subtract the user’s umask from files from the permissions specified in the archive. This is the default behavior for ordinary users.

‘--no-unquote’

Treat all input file or member names literally, do not interpret escape sequences. See [input name quoting], page 88.

‘--no-wildcards’

Do not use wildcards. See [controlling pattern-matching], page 94.

‘--no-wildcards-match-slash’

Wildcards do not match ‘/’. See [controlling pattern-matching], page 94.

‘--null’

When tar is using the ‘--files-from’ option, this option instructs tar to expect filenames terminated with ‘NUL’, so tar can correctly work with file names that contain newlines. See Section 6.3.1 [nul], page 91.

‘--numeric-owner’

This option will notify tar that it should use numeric user and group IDs when creating a tar file, rather than names. See Section 8.2 [Attributes], page 122.

‘-o’ The function of this option depends on the action tar is performing. When extracting files, ‘-o’ is a synonym for ‘--no-same-owner’, i.e., it prevents tar from restoring ownership of files being extracted. When creating an archive, it is a synonym for ‘--old-archive’. This behavior is for compatibility with previous versions of GNU tar, and will be removed in the future releases. See Appendix A [Changes], page 159, for more information.

‘--occurrence[=number]’

This option can be used in conjunction with one of the sub-

commands ‘--delete’, ‘--diff’, ‘--extract’ or ‘--list’ when

a list of files is given either on the command line or via ‘-T’ option. This option instructs tar to process only the numberth occurrence of each named file. Number defaults to 1, so

     tar -x -f archive.tar --occurrence filename

will extract the first occurrence of the member ‘filename’ from ‘archive.tar’ and will terminate without scanning to the end of the archive.

‘--old-archive’

Synonym for ‘--format=v7’.

‘--one-file-system’

Used when creating an archive. Prevents tar from recursing into directories that are on different file systems from the current directory^{5}.

‘--overwrite’

Overwrite existing files and directory metadata when extracting files from an archive. See [Overwrite Old Files], page 63.

‘--overwrite-dir’

Overwrite the metadata of existing directories when extracting files from an archive. See [Overwrite Old Files], page 63.

‘--owner=user’

Specifies that tar should use user as the owner of members when creating archives, instead of the user associated with the source file. user is first decoded as a user symbolic name, but if this interpretation fails, it has to be a decimal numeric user ID. See Section 4.3.1 [override], page 59. This option does not affect extraction from archives.

5 Earlier versions of GNU tar understood ‘-l’ as a synonym for ‘--one-file-system’.

This has changed in version 1.15.91. See Appendix A [Changes], page 159, for more information.

‘--transform=sed-expr’

Transform file or member names using sed replacement expression sed-expr. For example,

     $ tar cf archive.tar --transform 's,^\./,usr/,' .

will add to ‘archive’ files from the current working directory, replacing initial ‘./’ prefix with ‘usr/’. For the detailed discussion, See Section 6.7 [transform], page 100. To see transformed member names in verbose listings, use ‘--show-transformed-names’ option (see [show-transformed-names], page 101).

‘--quote-chars=string’

Always quote characters from string, even if the selected quoting style would not quote them (see Section 6.6 [quoting styles], page 96).

‘--quoting-style=style’

Set quoting style to use when printing member and file names (see Section 6.6 [quoting styles], page 96). Valid style values

are: literal, shell, shell-always, c, escape, locale, and

clocale. Default quoting style is escape, unless overridden while configuring the package.

‘--pax-option=keyword-list’

This option is meaningful only with POSIX.1-2001 archives (see Section 8.3.6 [posix], page 126). It modifies the way tar handles the extended header keywords. Keyword-list is a commaseparated list of keyword options. See Section 8.3.6.1 [PAX keywords], page 126, for a detailed discussion.

‘--portability’
‘--old-archive’

Synonym for ‘--format=v7’.

‘--posix’ Same as ‘--format=posix’.
‘--preserve’

Synonymous with specifying both ‘--preserve-permissions’ and ‘--same-order’. See [Setting Access Permissions], page 65.

‘--preserve-order’

(See ‘--same-order’; see Section 4.4.1 [Reading], page 61.)

‘--preserve-permissions’
‘--same-permissions’
‘-p’

When tar is extracting an archive, it normally subtracts the users’ umask from the permissions specified in the archive and uses that number as the permissions to create the destination file. Specifying this option instructs tar that it should use the permissions directly from the archive. See [Setting Access Permissions], page 65.

‘--read-full-records’
‘-B’

Specifies that tar should reblock its input, for reading from pipes on systems with buggy implementations. See Section 4.4.1 [Reading], page 61.

‘--record-size=size’

Instructs tar to use size bytes per record when accessing the archive. See Section 9.4.2 [Blocking Factor], page 143.

‘--recursion’

With this option, tar recurses into directories. See Section 6.9 [recurse], page 104.

‘--recursive-unlink’

Remove existing directory hierarchies before extracting directories of the same name from the archive. See [Recursive Unlink], page 64.

‘--remove-files’

Directs tar to remove the source file from the file system after appending it to an archive. See [remove files], page 68.

‘--restrict’

Disable use of some potentially harmful tar options. Currently this option disables shell invocation from multi-volume menu (see Section 9.6 [Using Multiple Tapes], page 150).

‘--rmt-command=cmd’

Notifies tar that it should use cmd instead of the default ‘/usr/libexec/rmt’ (see Section 9.2 [Remote Tape Server], page 139).

‘--rsh-command=cmd’

Notifies tar that is should use cmd to communicate with remote devices. See Section 9.1 [Device], page 137.

‘--same-order’
‘--preserve-order’
‘-s’

This option is an optimization for tar when running on machines with small amounts of memory. It informs tar that the list of file arguments has already been sorted to match the order of files in the archive. See Section 4.4.1 [Reading], page 61.

‘--same-owner’

When extracting an archive, tar will attempt to preserve the owner specified in the tar archive with this option present. This is the default behavior for the superuser; this option has an effect only for ordinary users. See Section 8.2 [Attributes], page 122.

‘--same-permissions’

(See ‘--preserve-permissions’; see [Setting Access Permissions], page 65.)

‘--show-defaults’

Displays the default options used by tar and exits successfully. This option is intended for use in shell scripts. Here is an example of what you can see using this option:

     $ tar --show-defaults
     --format=gnu -f- -b20 --quoting-style=escape \
     --rmt-command=/usr/libexec/rmt --rsh-command=/usr/bin/rsh
‘--show-omitted-dirs’

Instructs tar to mention directories its skipping over when operating on a tar archive. See [show-omitted-dirs], page 48.

‘--show-transformed-names’
‘--show-stored-names’

Display file or member names after applying any transformations (see Section 6.7 [transform], page 100). In particular, when used in conjunction with one of archive creation operations it instructs tar to list the member names stored in the archive, as opposed to the actual file names. See [listing member and file names], page 15.

‘--sparse’
‘-S’

Invokes a GNU extension when adding files to an archive that handles sparse files efficiently. See Section 8.1.2 [sparse], page 121.

‘--sparse-version=version’

Specified the format version to use when archiving sparse files. Implies ‘--sparse’. See Section 8.1.2 [sparse], page 121. For the description of the supported sparse formats, See [Sparse Formats], page 174.

‘--starting-file=name’
‘-K name’

This option affects extraction only; tar will skip extracting files in the archive until it finds one that matches name. See Section 4.4.3 [Scarce], page 68.

‘--strip-components=number’

Strip given number of leading components from file names before extraction.^{6} For example, if archive ‘archive.tar’ contained

‘/some/file/name’, then running
     tar --extract --file archive.tar --strip-components=2

would extract this file to file ‘name’. , summary

‘--suffix=suffix’

Alters the suffix tar uses when backing up files from the default ‘~’. See Section 4.5 [backup], page 69.

‘--tape-length=num’
‘-L num’

Specifies the length of tapes that tar is writing as being num x 1024 bytes long. See Section 9.6 [Using Multiple Tapes], page 150.

‘--test-label’

Reads the volume label. If an argument is specified, test whether it matches the volume label. See [–test-label option], page 155.

‘--to-command=command’

During extraction tar will pipe extracted files to the standard input of command. See [Writing to an External Program], page 67.

‘--to-stdout’
‘-O’

During extraction, tar will extract files to stdout rather than to the file system. See [Writing to Standard Output], page 66.

‘--totals[=signo]’

Displays the total number of bytes transferred when processing an archive. If an argument is given, these data are displayed on request, when signal signo is delivered to tar. See [totals], page 47.

‘--touch’
‘-m’

Sets the data modification time of extracted files to the extraction time, rather than the data modification time stored in the archive. See [Data Modification Times], page 64.

‘--uncompress’

(See ‘--compress’. see Section 8.1.1 [gzip], page 118)

6 This option was called ‘--strip-path’ in version 1.14.

‘--ungzip’

(See ‘--gzip’. see Section 8.1.1 [gzip], page 118)

‘--unlink-first’
‘-U’

Directs tar to remove the corresponding file from the file system before extracting it from the archive. See [Unlink First], page 64.

‘--unquote’

Enable unquoting input file or member names (default). See [input name quoting], page 88.

‘--use-compress-program=prog’

Instructs tar to access the archive through prog, which is presumed to be a compression program of some sort. See Section 8.1.1 [gzip], page 118.

‘--utc’

Display file modification dates in UTC. This option implies

‘--verbose’.
‘--verbose’
‘-v’

Specifies that tar should be more verbose about the operations its performing. This option can be specified multiple times for some operations to increase the amount of information displayed. See Section 3.7 [verbose], page 46.

‘--verify’
‘-W’

Verifies that the archive was correctly written when creating an archive. See Section 9.8 [verify], page 157.

‘--version’

Print information about the program’s name, version, origin and legal status, all on standard output, and then exit successfully. See Section 3.5 [help], page 44.

‘--volno-file=file’

Used in conjunction with ‘--multi-volume’. tar will keep track of which volume of a multi-volume archive its working in file. See [volno-file], page 152.

‘--wildcards’

Use wildcards when matching member names with patterns. See [controlling pattern-matching], page 94.

‘--wildcards-match-slash’

Wildcards match ‘/’. See [controlling pattern-matching], page 94.

Chapter 3: Invoking GNU tar 43

3.4.3 Short Options Cross Reference

Here is an alphabetized list of all of the short option forms, matching them with the equivalent long option.

Short Option Reference
-A  [–concatenate], page 28.
-B  [–read-full-records], page 39.
-C  [–directory], page 31.
-F  [–info-script], page 34.
-G  [–incremental], page 33.
-K  [–starting-file], page 40.
-L  [–tape-length], page 41.
-M  [–multi-volume], page 35.
-N  [–newer], page 35.
-O  [–to-stdout], page 41.
-P  [–absolute-names], page 29.
-R  [–block-number], page 30.
-S  [–sparse], page 40.
-T  [–files-from], page 32.
-U  [–unlink-first], page 42.
-V  [–label], page 34.
-W  [–verify], page 42.
-X  [–exclude-from], page 32.
-Z  [–compress], page 31.
-b  [–blocking-factor], page 30.
-c  [–create], page 28.

-d  [–compare], page 28.
-f  [–file], page 32.
-g  [–listed-incremental], page 34.
-h  [–dereference], page 31.
-i  [–ignore-zeros], page 33.
-j  [–bzip2], page 30.
-k  [–keep-old-files], page 34.
-l  [–check-links], page 31.
-m  [–touch], page 41.

-o When creating, [–no-same-owner], page 36, when extracting — [–portability], page 38. The later usage is deprecated. It is retained for compatibility with the earlier versions of GNU tar. In the future releases ‘-o’ will be equivalent to ‘--no-same-owner’ only.

-p  [–preserve-permissions], page 38.
-r  [–append], page 28.
-s  [–same-order], page 39.
-t  [–list], page 29.
-u  [–update], page 29.
-v  [–verbose], page 42.
-w  [–interactive], page 34.
-x  [–extract], page 29.
-z  [–gzip], page 33.

3.5 GNU tar documentation

Being careful, the first thing is really checking that you are using GNU tar, indeed. The ‘--version’ option causes tar to print information about its name, version, origin and legal status, all on standard output, and then exit successfully. For example, ‘tar --version’ might print:

     tar (GNU tar) 1.16.1
     Copyright (C) 2006 Free Software Foundation, Inc.
     This is free software.  You may redistribute copies of it under the terms
     of the GNU General Public License <http://www.gnu.org/licenses/gpl.html>.
     There is NO WARRANTY, to the extent permitted by law.
     Written by John Gilmore and Jay Fenlason.

The first occurrence of ‘tar’ in the result above is the program name in the package (for example, rmt is another program), while the second occurrence of ‘tar’ is the name of the package itself, containing possibly many programs. The package is currently named ‘tar’, after the name of the main program it contains^{7}.

Another thing you might want to do is checking the spelling or meaning of some particular tar option, without resorting to this manual, for once you have carefully read it. GNU tar has a short help feature, triggerable through the ‘--help’ option. By using this option, tar will print a usage message listing all available options on standard output, then exit successfully, without doing anything else and ignoring all other options. Even if this is only a brief summary, it may be several screens long. So, if you are not using some kind of scrollable window, you might prefer to use something like:

     $ tar --help | less

presuming, here, that you like using less for a pager. Other popular pagers are more and pg. If you know about some keyword which interests you and do not want to read all the ‘--help’ output, another common idiom is doing:

     tar --help | grep keyword

for getting only the pertinent lines. Notice, however, that some tar options have long description lines and the above command will list only the first of them.

The exact look of the option summary displayed by tar --help is configurable. See Appendix B [Configuring Help Summary], page 161, for a detailed description.

If you only wish to check the spelling of an option, running tar --usage may be a better choice. This will display a terse list of tar option without accompanying explanations.

The short help output is quite succinct, and you might have to get back to the full documentation for precise points. If you are reading this paragraph, you already have the tar manual in some form. This manual is available

in a variety of forms from http://www.gnu.org/software/tar/manual. It

may be printed out of the GNU tar distribution, provided you have TEX already installed somewhere, and a laser printer around. Just configure the distribution, execute the command ‘make dvi’, then print ‘doc/tar.dvi’ the usual way (contact your local guru to know how). If GNU tar has been conveniently installed at your place, this manual is also available in interactive, hypertextual form as an Info file. Just call ‘info tar’ or, if you do not have the info program handy, use the Info reader provided within GNU Emacs, calling ‘tar’ from the main Info menu.

7 There are plans to merge the cpio and tar packages into a single one which would be

called paxutils. So, who knows if, one of this days, the ‘--version’ would not output

‘tar (GNU paxutils) 3.2’

There is currently no man page for GNU tar. If you observe such a man page on the system you are running, either it does not belong to GNU tar, or it has not been produced by GNU. Some package maintainers convert tar --help output to a man page, using help2man. In any case, please bear in mind that the authoritative source of information about GNU tar is this Texinfo documentation.

3.6 Obtaining GNU tar default values

GNU tar has some predefined defaults that are used when you do not explicitly specify another values. To obtain a list of such defaults, use ‘--show-defaults’ option. This will output the values in the form of tar command line options:

     tar --show-defaults
     --format=gnu -f- -b20 --quoting-style=escape
     --rmt-command=/etc/rmt --rsh-command=/usr/bin/rsh

Notice, that this option outputs only one line. The example output above has been split to fit page boundaries. The above output shows that this version of GNU tar defaults to using ‘gnu’ archive format (see Chapter 8 [Formats], page 117), it uses standard output as the archive, if no ‘--file’ option has been given (see [file tutorial], page 7), the default blocking factor is 20 (see Section 9.4.2 [Blocking Factor], page 143). It also shows the default locations where tar will look for rmt and rsh binaries.

3.7 Checking tar progress

Typically, tar performs most operations without reporting any information to the user except error messages. When using tar with many options, particularly ones with complicated or difficult-to-predict behavior, it is possible to make serious mistakes. tar provides several options that make observing tar easier. These options cause tar to print information as it progresses in its job, and you might want to use them just for being more careful about what is going on, or merely for entertaining yourself. If you have encountered a problem when operating on an archive, however, you may need more information than just an error message in order to solve the problem. The following options can be helpful diagnostic tools.

Normally, the ‘--list’ (‘-t’) command to list an archive prints just the file names (one per line) and the other commands are silent. When used with most operations, the ‘--verbose’ (‘-v’) option causes tar to print the name of each file or archive member as it is processed. This and the other options which make tar print status information can be useful in monitoring

tar.

With ‘--create’ or ‘--extract’, ‘--verbose’ used once just prints the names of the files or members as they are processed. Using it twice causes tar to print a longer listing (See [verbose member listing], page 9, for the description) for each member. Since ‘--list’ already prints the names of the members, ‘--verbose’ used once with ‘--list’ causes tar to print an ‘ls -l’ type listing of the files in the archive. The following examples both extract members with long list output:

     $ tar --extract --file=archive.tar --verbose --verbose
     $ tar xvvf archive.tar

Verbose output appears on the standard output except when an archive is being written to the standard output, as with ‘tar --create --file=- --verbose’ (‘tar cfv -’, or even ‘tar cv’—if the installer let standard output be the default archive). In that case tar writes verbose output to the standard error stream.

If ‘--index-file=file’ is specified, tar sends verbose output to file rather than to standard output or standard error.

The ‘--totals’ option causes tar to print on the standard error the total amount of bytes transferred when processing an archive. When creating or appending to an archive, this option prints the number of bytes written to the archive and the average speed at which they have been written, e.g.:

     $ tar -c -f archive.tar --totals /home
     Total bytes written: 7924664320 (7.4GiB, 85MiB/s)

When reading an archive, this option displays the number of bytes read:

     $ tar -x -f archive.tar --totals
     Total bytes read: 7924664320 (7.4GiB, 95MiB/s)

Finally, when deleting from an archive, the ‘--totals’ option displays both numbers plus number of bytes removed from the archive:

     $ tar --delete -f foo.tar --totals --wildcards '*~'
     Total bytes read: 9543680 (9.2MiB, 201MiB/s)
     Total bytes written: 3829760 (3.7MiB, 81MiB/s)
     Total bytes deleted: 1474048

You can also obtain this information on request. When ‘--totals’ is used with an argument, this argument is interpreted as a symbolic name of a signal, upon delivery of which the statistics is to be printed:

‘--totals=signo’

Print statistics upon delivery of signal signo. Valid arguments

          are: SIGHUP, SIGQUIT, SIGINT, SIGUSR1 and SIGUSR2. Short-

ened names without ‘SIG’ prefix are also accepted.

Both forms of ‘--totals’ option can be used simultaneously. Thus, tar -x --totals --totals=USR1 instructs tar to extract all members from its default archive and print statistics after finishing the extraction, as well as when receiving signal SIGUSR1.

The ‘--checkpoint’ option prints an occasional message as tar reads or writes the archive. It is designed for those who don’t need the more detailed (and voluminous) output of ‘--block-number’ (‘-R’), but do want visual confirmation that tar is actually making forward progress. By default it prints a message each 10 records read or written. This can be changed by giving it a numeric argument after an equal sign:

     $ tar -c --checkpoint=1000 /var
     tar: Write checkpoint 1000
     tar: Write checkpoint 2000
     tar: Write checkpoint 3000

This example shows the default checkpoint message used by tar. If you place a dot immediately after the equal sign, it will print a ‘.’ at each checkpoint. For example:

     $ tar -c --checkpoint=.1000 /var
     ...

The ‘--show-omitted-dirs’ option, when reading an archive—with ‘--list’ or ‘--extract’, for example—causes a message to be printed for each directory in the archive which is skipped. This happens regardless of the reason for skipping: the directory might not have been named on the command line (implicitly or explicitly), it might be excluded by the use of the ‘--exclude=pattern’ option, or some other reason.

If ‘--block-number’ (‘-R’) is used, tar prints, along with every message it would normally produce, the block number within the archive where the message was triggered. Also, supplementary messages are triggered when reading blocks full of NULs, or when hitting end of file on the archive. As of now, if the archive if properly terminated with a NUL block, the reading of the file may stop before end of file is met, so the position of end of file will not usually show when ‘--block-number’ (‘-R’) is used. Note that GNU tar drains the archive before exiting when reading the archive from a pipe.

This option is especially useful when reading damaged archives, since it helps pinpoint the damaged sections. It can also be used with ‘--list’ (‘-t’) when listing a file-system backup tape, allowing you to choose among several backup tapes when retrieving a file later, in favor of the tape where the file appears earliest (closest to the front of the tape). See Section 4.5 [backup], page 69.

3.8 Asking for Confirmation During Operations

Typically, tar carries out a command without stopping for further instructions. In some situations however, you may want to exclude some files and archive members from the operation (for instance if disk or storage space is tight). You can do this by excluding certain files automatically (see Chapter 6 [Choosing], page 87), or by performing an operation interactively, using

the ‘--interactive’ (‘-w’) option. tar also accepts ‘--confirmation’ for

this option.

When the ‘--interactive’ (‘-w’) option is specified, before reading, writing, or deleting files, tar first prints a message for each such file, telling what operation it intends to take, then asks for confirmation on the terminal. The actions which require confirmation include adding a file to the archive, extracting a file from the archive, deleting a file from the archive, and deleting a file from disk. To confirm the action, you must type a line of input beginning with ‘y’. If your input line begins with anything other than ‘y’, tar skips that file.

If tar is reading the archive from the standard input, tar opens the file ‘/dev/tty’ to support the interactive communications.

Verbose output is normally sent to standard output, separate from other error messages. However, if the archive is produced directly on standard output, then verbose output is mixed with errors on stderr. Producing the archive on standard output may be used as a way to avoid using disk space, when the archive is soon to be consumed by another process reading it, say. Some people felt the need of producing an archive on stdout, still willing to segregate between verbose output and error output. A possible approach would be using a named pipe to receive the archive, and having the consumer process to read from that named pipe. This has the advantage of letting standard output free to receive verbose output, all separate from errors.

50 GNU tar 1.16.1

4 GNU tar Operations

4.1 Basic GNU tar Operations

The basic tar operations, ‘--create’ (‘-c’), ‘--list’ (‘-t’) and ‘--extract’ (‘--get’, ‘-x’), are currently presented and described in the tutorial chapter of this manual. This section provides some complementary notes for these operations.

‘--create’
‘-c’

Creating an empty archive would have some kind of elegance. One can initialize an empty archive and later use ‘--append’ (‘-r’) for adding all members. Some applications would not welcome making an exception in the way of adding the first archive member. On the other hand, many people reported that it is dangerously too easy for tar to destroy a magnetic tape with an empty archive^{1}. The two most common errors are:

1. Mistakingly using create instead of extract, when the intent was to extract the full contents of an archive. This error is likely: keys c and x are right next to each other on the QWERTY keyboard. Instead of being unpacked, the archive then gets wholly destroyed. When users speak about exploding an archive, they usually mean something else :-).

2. Forgetting the argument to file, when the intent was to create an archive with a single file in it. This error is likely because a tired user can easily add the f key to the cluster of option letters, by the mere force of habit, without realizing the full consequence of doing so. The usual consequence is that the single file, which was meant to be saved, is rather destroyed.

So, recognizing the likelihood and the catastrophic nature of these errors, GNU tar now takes some distance from elegance, and cowardly refuses to create an archive when ‘--create’ option is given, there are no arguments besides options, and ‘--files-from’ (‘-T’) option is not used. To get around the cautiousness of GNU tar and nevertheless create an archive with nothing in it, one may still use, as the value for the ‘--files-from’ option, a file with no names in it, as shown in the following commands:

1 This is well described in Unix-haters Handbook, by Simson Garfinkel, Daniel Weise &

Steven Strassmann, IDG Books, ISBN 1-56884-203-1.

                 tar --create --file=empty-archive.tar --files-from=/dev/null
                 tar cfT empty-archive.tar /dev/null
‘--extract’
‘--get’
‘-x’

A socket is stored, within a GNU tar archive, as a pipe.

‘`--list' (`-t')’

GNU tar now shows dates as ‘1996-08-30’, while it used to show them as ‘Aug 30 1996’. Preferably, people should get used to ISO 8601 dates. Local American dates should be made available again with full date localization support, once ready. In the meantime, programs not being localizable for dates should prefer international dates, that’s really the way to go.

          Look up http://www.cl.cam.ac.uk/~mgk25/iso-time.html if

you are curious, it contains a detailed explanation of the ISO 8601 standard.

4.2 Advanced GNU tar Operations

Now that you have learned the basics of using GNU tar, you may want to learn about further ways in which tar can help you.

This chapter presents five, more advanced operations which you probably won’t use on a daily basis, but which serve more specialized functions. We also explain the different styles of options and why you might want to use one or another, or a combination of them in your tar commands. Additionally, this chapter includes options which allow you to define the output from tar more carefully, and provide help and error correction in special circumstances.

4.2.1 The Five Advanced tar Operations

(This message will disappear, once this node revised.)

In the last chapter, you learned about the first three operations to tar. This chapter presents the remaining five operations to tar: ‘--append’,

‘--update’, ‘--concatenate’, ‘--delete’, and ‘--compare’.

You are not likely to use these operations as frequently as those covered in the last chapter; however, since they perform specialized functions, they are quite useful when you do need to use them. We will give examples using the same directory and files that you created in the last chapter. As you may recall, the directory is called ‘practice’, the files are ‘jazz’, ‘blues’, ‘folk’, ‘rock’, and the two archive files you created are ‘collection.tar’

and ‘music.tar’.

We will also use the archive files ‘afiles.tar’ and ‘bfiles.tar’. The archive ‘afiles.tar’ contains the members ‘apple’, ‘angst’, and ‘aspic’; ‘bfiles.tar’ contains the members ‘./birds’, ‘baboon’, and ‘./box’.

Unless we state otherwise, all practicing you do and examples you follow in this chapter will take place in the ‘practice’ directory that you created in the previous chapter; see Section 2.6.1 [prepare for examples], page 11. (Below in this section, we will remind you of the state of the examples where the last chapter left them.)

The five operations that we will cover in this chapter are:

‘--append’

‘-r’ Add new entries to an archive that already exists.

‘--update’

‘-r’ Add more recent copies of archive members to the end of an archive, if they exist.

‘--concatenate’
‘--catenate’

‘-A’ Add one or more pre-existing archives to the end of another archive.

‘--delete’

Delete items from an archive (does not work on tapes).

‘--compare’
‘--diff’

‘-d’ Compare archive members to their counterparts in the file system.

4.2.2 How to Add Files to Existing Archives:

     ‘--append’

(This message will disappear, once this node revised.)

If you want to add files to an existing archive, you don’t need to create a new archive; you can use ‘--append’ (‘-r’). The archive must already exist in order to use ‘--append’. (A related operation is the ‘--update’ operation; you can use this to add newer versions of archive members to an existing archive. To learn how to do this with ‘--update’, see Section 4.2.3 [update], page 56.)

If you use ‘--append’ to add a file that has the same name as an archive member to an archive containing that archive member, then the old member is not deleted. What does happen, however, is somewhat complex. tar allows you to have infinite number of files with the same name. Some operations treat these same-named members no differently than any other set of archive members: for example, if you view an archive with ‘--list’ (‘-t’), you will see all of those members listed, with their data modification times, owners, etc.

Other operations don’t deal with these members as perfectly as you might prefer; if you were to use ‘--extract’ to extract the archive, only the most recently added copy of a member with the same name as four other members would end up in the working directory. This is because ‘--extract’ extracts an archive in the order the members appeared in the archive; the most recently archived members will be extracted last. Additionally, an extracted member will replace a file of the same name which existed in the directory already, and tar will not prompt you about this^{2}. Thus, only the most recently archived member will end up being extracted, as it will replace the one extracted before it, and so on.

There exists a special option that allows you to get around this behavior and extract (or list) only a particular copy of the file. This is ‘--occurrence’ option. If you run tar with this option, it will extract only the first copy of the file. You may also give this option an argument specifying the number of copy to be extracted. Thus, for example if the archive ‘archive.tar’ contained three copies of file ‘myfile’, then the command

     tar --extract --file archive.tar --occurrence=2 myfile

would extract only the second copy. See Section 3.4.2 [Option Summary], page 29, for the description of ‘--occurrence’ option.

If you want to replace an archive member, use ‘--delete’ to delete the member you want to remove from the archive, , and then use ‘--append’ to add the member you want to be in the archive. Note that you can not change the order of the archive; the most recently added member will still appear last. In this sense, you cannot truly “replace” one member with another. (Replacing one member with another will not work on certain types of media, such as tapes; see Section 4.2.5 [delete], page 58 and Chapter 9 [Media], page 137, for more information.)

4.2.2.1 Appending Files to an Archive

(This message will disappear, once this node revised.)

The simplest way to add a file to an already existing archive is the ‘--append’ (‘-r’) operation, which writes specified files into the archive whether or not they are already among the archived files.

When you use ‘--append’, you must specify file name arguments, as there is no default. If you specify a file that already exists in the archive, another copy of the file will be added to the end of the archive. As with other operations, the member names of the newly added files will be exactly the same as their names given on the command line. The ‘--verbose’ (‘-v’) option will print out the names of the files as they are written into the archive.

‘--append’ cannot be performed on some tape drives, unfortunately, due to deficiencies in the formats those tape drives use. The archive must be a valid tar archive, or else the results of using this operation will be unpredictable. See Chapter 9 [Media], page 137.

To demonstrate using ‘--append’ to add a file to an archive, create a file called ‘rock’ in the ‘practice’ directory. Make sure you are in the ‘practice’ directory. Then, run the following tar command to add ‘rock’

2 Unless you give it ‘--keep-old-files’ option, or the disk copy is newer than the the

one in the archive and you invoke tar with ‘--keep-newer-files’ option

to ‘collection.tar’:
     $ tar --append --file=collection.tar rock

If you now use the ‘--list’ (‘-t’) operation, you will see that ‘rock’ has been added to the archive:

     $ tar --list --file=collection.tar
     -rw-r--r-- me user     28 1996-10-18 16:31 jazz
     -rw-r--r-- me user     21 1996-09-23 16:44 blues
     -rw-r--r-- me user     20 1996-09-23 16:44 folk
     -rw-r--r-- me user     20 1996-09-23 16:44 rock

4.2.2.2 Multiple Members with the Same Name

You can use ‘--append’ (‘-r’) to add copies of files which have been updated since the archive was created. (However, we do not recommend doing this since there is another tar option called ‘--update’; See Section 4.2.3 [update], page 56, for more information. We describe this use of ‘--append’ here for the sake of completeness.) When you extract the archive, the older version will be effectively lost. This works because files are extracted from an archive in the order in which they were archived. Thus, when the archive is extracted, a file archived later in time will replace a file of the same name which was archived earlier, even though the older version of the file will remain in the archive unless you delete all versions of the file.

Supposing you change the file ‘blues’ and then append the changed version to ‘collection.tar’. As you saw above, the original ‘blues’ is in the archive ‘collection.tar’. If you change the file and append the new version of the file to the archive, there will be two copies in the archive. When you extract the archive, the older version of the file will be extracted first, and then replaced by the newer version when it is extracted.

You can append the new, changed copy of the file ‘blues’ to the archive in this way:

     $ tar --append --verbose --file=collection.tar blues
     blues

Because you specified the ‘--verbose’ option, tar has printed the name of the file being appended as it was acted on. Now list the contents of the archive:

     $ tar --list --verbose --file=collection.tar
     -rw-r--r-- me user     28 1996-10-18 16:31 jazz
     -rw-r--r-- me user     21 1996-09-23 16:44 blues
     -rw-r--r-- me user     20 1996-09-23 16:44 folk
     -rw-r--r-- me user     20 1996-09-23 16:44 rock
     -rw-r--r-- me user     58 1996-10-24 18:30 blues

The newest version of ‘blues’ is now at the end of the archive (note the different creation dates and file sizes). If you extract the archive, the older version of the file ‘blues’ will be replaced by the newer version. You can confirm this by extracting the archive and running ‘ls’ on the directory.

If you wish to extract the first occurrence of the file ‘blues’ from the archive, use ‘--occurrence’ option, as shown in the following example:

     $ tar --extract -vv --occurrence --file=collection.tar blues
     -rw-r--r-- me user     21 1996-09-23 16:44 blues

See Section 4.4.2 [Writing], page 62, for more information on ‘--extract’ and See Section 3.4.2 [Option Summary], page 29, for the description of

‘--occurrence’ option.

4.2.3 Updating an Archive

(This message will disappear, once this node revised.)

In the previous section, you learned how to use ‘--append’ to add a file to an existing archive. A related operation is ‘--update’ (‘-u’). The ‘--update’ operation updates a tar archive by comparing the date of the specified archive members against the date of the file with the same name. If the file has been modified more recently than the archive member, then the newer version of the file is added to the archive (as with ‘--append’).

Unfortunately, you cannot use ‘--update’ with magnetic tape drives. The operation will fail.

Both ‘--update’ and ‘--append’ work by adding to the end of the archive. When you extract a file from the archive, only the version stored last will wind up in the file system, unless you use the ‘--backup’ option. See Section 4.2.2.2 [multiple], page 55, for a detailed discussion.

4.2.3.1 How to Update an Archive Using ‘--update’

You must use file name arguments with the ‘--update’ (‘-u’) operation. If you don’t specify any files, tar won’t act on any files and won’t tell you that it didn’t do anything (which may end up confusing you).

To see the ‘--update’ option at work, create a new file, ‘classical’, in your practice directory, and some extra text to the file ‘blues’, using any text editor. Then invoke tar with the ‘update’ operation and the ‘--verbose’ (‘-v’) option specified, using the names of all the files in the practice directory as file name arguments:

     $ tar --update -v -f collection.tar blues folk rock classical
     blues
     classical
     $

Because we have specified verbose mode, tar prints out the names of the files it is working on, which in this case are the names of the files that needed to be updated. If you run ‘tar --list’ and look at the archive, you will see ‘blues’ and ‘classical’ at its end. There will be a total of two versions of the member ‘blues’; the one at the end will be newer and larger, since you added text before updating it.

(The reason tar does not overwrite the older file when updating it is because writing to the middle of a section of tape is a difficult process.

Tapes are not designed to go backward. See Chapter 9 [Media], page 137, for more information about tapes.

‘--update’ (‘-u’) is not suitable for performing backups for two reasons: it does not change directory content entries, and it lengthens the archive every time it is used. The GNU tar options intended specifically for backups are more efficient. If you need to run backups, please consult Chapter 5 [Backups], page 73.

4.2.4 Combining Archives with ‘--concatenate’

Sometimes it may be convenient to add a second archive onto the end of an archive rather than adding individual files to the archive. To add one or more archives to the end of another archive, you should use the ‘--concatenate’ (‘--catenate’, ‘-A’) operation.

To use ‘--concatenate’, give the first archive with ‘--file’ option and name the rest of archives to be concatenated on the command line. The members, and their member names, will be copied verbatim from those archives to the first one.^{3} The new, concatenated archive will be called by the same name as the one given with the ‘--file’ option. As usual, if you omit ‘--file’, tar will use the value of the environment variable TAPE, or, if this has not been set, the default archive name.

To demonstrate how ‘--concatenate’ works, create two small archives called ‘bluesrock.tar’ and ‘folkjazz.tar’, using the relevant files from

‘practice’:
     $ tar -cvf bluesrock.tar blues rock
     blues
     rock
     $ tar -cvf folkjazz.tar folk jazz
     folk
     jazz

If you like, You can run ‘tar --list’ to make sure the archives contain what they are supposed to:

     $ tar -tvf bluesrock.tar
     -rw-r--r-- melissa user    105 1997-01-21 19:42 blues
     -rw-r--r-- melissa user     33 1997-01-20 15:34 rock
     $ tar -tvf jazzfolk.tar
     -rw-r--r-- melissa user     20 1996-09-23 16:44 folk
     -rw-r--r-- melissa user     65 1997-01-30 14:15 jazz

We can concatenate these two archives with tar:

     $ cd ..
     $ tar --concatenate --file=bluesrock.tar jazzfolk.tar

If you now list the contents of the ‘bluesrock.tar’, you will see that now it also contains the archive members of ‘jazzfolk.tar’:

     $ tar --list --file=bluesrock.tar

3 This can cause multiple members to have the same name, for information on how this

affects reading the archive, Section 4.2.2.2 [multiple], page 55.

     blues
     rock
     folk
     jazz

When you use ‘--concatenate’, the source and target archives must already exist and must have been created using compatible format parameters. Notice, that tar does not check whether the archives it concatenates have compatible formats, it does not even check if the files are really tar archives.

Like ‘--append’ (‘-r’), this operation cannot be performed on some tape drives, due to deficiencies in the formats those tape drives use.

It may seem more intuitive to you to want or try to use cat to concatenate two archives instead of using the ‘--concatenate’ operation; after all, cat is the utility for combining files.

However, tar archives incorporate an end-of-file marker which must be removed if the concatenated archives are to be read properly as one archive. ‘--concatenate’ removes the end-of-archive marker from the target archive before each new archive is appended. If you use cat to combine the archives, the result will not be a valid tar format archive. If you need to retrieve files from an archive that was added to using the cat utility, use the ‘--ignore-zeros’ (‘-i’) option. See [Ignore Zeros], page 62, for further information on dealing with archives improperly combined using the cat shell utility.

4.2.5 Removing Archive Members Using ‘--delete’

(This message will disappear, once this node revised.)

You can remove members from an archive by using the ‘--delete’ option. Specify the name of the archive with ‘--file’ (‘-f’) and then specify the names of the members to be deleted; if you list no member names, nothing will be deleted. The ‘--verbose’ option will cause tar to print the names of the members as they are deleted. As with ‘--extract’, you must give the exact member names when using ‘tar --delete’. ‘--delete’ will remove all versions of the named file from the archive. The ‘--delete’ operation can run very slowly.

Unlike other operations, ‘--delete’ has no short form.

This operation will rewrite the archive. You can only use ‘--delete’ on an archive if the archive device allows you to write to any point on the media, such as a disk; because of this, it does not work on magnetic tapes. Do not try to delete an archive member from a magnetic tape; the action will not succeed, and you will be likely to scramble the archive and damage your tape. There is no safe way (except by completely re-writing the archive) to delete files from most kinds of magnetic tape. See Chapter 9 [Media], page 137.

To delete all versions of the file ‘blues’ from the archive ‘collection.tar’ in the ‘practice’ directory, make sure you are in that directory, and then,

     $ tar --list --file=collection.tar
     blues
     folk
     jazz
     rock
     $ tar --delete --file=collection.tar blues
     $ tar --list --file=collection.tar
     folk
     jazz
     rock
     $

The ‘--delete’ option has been reported to work properly when tar acts as a filter from stdin to stdout.

4.2.6 Comparing Archive Members with the File System

(This message will disappear, once this node revised.)

The ‘--compare’ (‘-d’), or ‘--diff’ operation compares specified archive members against files with the same names, and then reports differences in file size, mode, owner, modification date and contents. You should only specify archive member names, not file names. If you do not name any members, then tar will compare the entire archive. If a file is represented in the archive but does not exist in the file system, tar reports a difference.

You have to specify the record size of the archive when modifying an archive with a non-default record size.

tar ignores files in the file system that do not have corresponding members in the archive.

The following example compares the archive members ‘rock’, ‘blues’ and ‘funk’ in the archive ‘bluesrock.tar’ with files of the same name in the file system. (Note that there is no file, ‘funk’; tar will report an error message.)

     $ tar --compare --file=bluesrock.tar rock blues funk
     rock
     blues
     tar: funk not found in archive

The spirit behind the ‘--compare’ (‘--diff’, ‘-d’) option is to check whether the archive represents the current state of files on disk, more than validating the integrity of the archive media. For this later goal, See Section 9.8 [verify], page 157.

4.3 Options Used by ‘--create’

The previous chapter described the basics of how to use ‘--create’ (‘-c’) to create an archive from a set of files. See Section 2.6 [create], page 10. This section described advanced options to be used with ‘--create’.

4.3.1 Overriding File Metadata

As described above, a tar archive keeps, for each member it contains, its metadata, such as modification time, mode and ownership of the file. GNU tar allows to replace these data with other values when adding files to the archive. The options described in this section affect creation of archives of any type. For POSIX archives, see also Section 8.3.6.1 [PAX keywords], page 126, for additional ways of controlling metadata, stored in the archive.

‘--mode=permissions’

When adding files to an archive, tar will use permissions for the archive members, rather than the permissions from the files. permissions can be specified either as an octal number or as symbolic permissions, like with chmod (See section “File permissions” in GNU file utilities. This reference also has useful information for those not being overly familiar with the UNIX permission system). Using latter syntax allows for more flexibility. For example, the value ‘a+rw’ adds read and write permissions for everybody, while retaining executable bits on directories or on any other file already marked as executable:

     $ tar -c -f archive.tar --mode='a+rw' .
‘--mtime=date’

When adding files to an archive, tar will use date as the modification time of members when creating archives, instead of their actual modification times. The argument date can be either a textual date representation in almost arbitrary format (see Chapter 7 [Date input formats], page 109) or a name of the existing file, starting with ‘/’ or ‘.’. In the latter case, the modification time of that file will be used. The following example will set the modification date to 00:00:00 UTC, January 1, 1970:

     $ tar -c -f archive.tar --mtime='1970-01-01' .

When used with ‘--verbose’ (see [verbose tutorial], page 8) GNU tar will try to convert the specified date back to its textual representation and compare it with the one given with ‘--mtime’ options. If the two dates differ, tar will print a warning saying what date it will use. This is to help user ensure he is using the right date. For example:

     $ tar -c -f archive.tar -v --mtime=yesterday .
     tar: Option --mtime: Treating date `yesterday' as 2006-06-20
     13:06:29.152478
     ...
‘--owner=user’

Specifies that tar should use user as the owner of members when creating archives, instead of the user associated with the source file. The argument user can be either an existing user symbolic name, or a decimal numeric user ID. There is no value indicating a missing number, and ‘0’ usually means root. Some people like to force ‘0’ as the value to offer in their distributions for the owner of files, because the root user is anonymous anyway, so that might as well be the owner of anonymous archives. For example:

                 $ tar -c -f archive.tar --owner=0 .

# Or:

                 $ tar -c -f archive.tar --owner=root .
‘--group=group’

Files added to the tar archive will have a group id of group, rather than the group from the source file. The argument group can be either an existing group symbolic name, or a decimal numeric group ID.

4.3.2 Ignore Fail Read

‘--ignore-failed-read’

Do not exit with nonzero on unreadable files or directories.

4.4 Options Used by ‘--extract’

(This message will disappear, once this node revised.)

The previous chapter showed how to use ‘--extract’ to extract an archive into the file system. Various options cause tar to extract more information than just file contents, such as the owner, the permissions, the modification date, and so forth. This section presents options to be used with ‘--extract’ when certain special considerations arise. You may review the information presented in Section 2.8 [extract], page 17 for more basic information about the ‘--extract’ operation.

4.4.1 Options to Help Read Archives

(This message will disappear, once this node revised.)

Normally, tar will request data in full record increments from an archive storage device. If the device cannot return a full record, tar will report an error. However, some devices do not always return full records, or do not require the last record of an archive to be padded out to the next record boundary. To keep reading until you obtain a full record, or to accept an incomplete record if it contains an end-of-archive marker, specify the ‘--read-full-records’ (‘-B’) option in conjunction with the ‘--extract’ or ‘--list’ operations. See Section 9.4 [Blocking], page 141.

The ‘--read-full-records’ (‘-B’) option is turned on by default when tar reads an archive from standard input, or from a remote machine. This is because on BSD Unix systems, attempting to read a pipe returns however much happens to be in the pipe, even if it is less than was requested. If this option were not enabled, tar would fail as soon as it read an incomplete record from the pipe.

If you’re not sure of the blocking factor of an archive, you can read the archive by specifying ‘--read-full-records’ (‘-B’) and

‘--blocking-factor=512-size’ (‘-b 512-size’), using a blocking factor

larger than what the archive uses. This lets you avoid having to determine the blocking factor of an archive. See Section 9.4.2 [Blocking Factor], page 143.

Reading Full Records

‘--read-full-records’

‘-B’ Use in conjunction with ‘--extract’ (‘--get’, ‘-x’) to read an archive which contains incomplete records, or one which has a blocking factor less than the one specified.

Ignoring Blocks of Zeros

Normally, tar stops reading when it encounters a block of zeros between file entries (which usually indicates the end of the archive). ‘--ignore-zeros’ (‘-i’) allows tar to completely read an archive which contains a block of zeros before the end (i.e., a damaged archive, or one that was created by concatenating several archives together).

The ‘--ignore-zeros’ (‘-i’) option is turned off by default because many versions of tar write garbage after the end-of-archive entry, since that part of the media is never supposed to be read. GNU tar does not write after the end of an archive, but seeks to maintain compatibility among archiving utilities.

‘--ignore-zeros’

‘-i’ To ignore blocks of zeros (i.e., end-of-archive entries) which may be encountered while reading an archive. Use in conjunction

          with ‘--extract’ or ‘--list’.

4.4.2 Changing How tar Writes Files

(This message will disappear, once this node revised.)

Options Controlling the Overwriting of Existing Files

When extracting files, if tar discovers that the extracted file already exists, it normally replaces the file by removing it before extracting it, to prevent confusion in the presence of hard or symbolic links. (If the existing file is a symbolic link, it is removed, not followed.) However, if a directory cannot be removed because it is nonempty, tar normally overwrites its metadata (ownership, permission, etc.). The ‘--overwrite-dir’ option enables this default behavior. To be more cautious and preserve the metadata of such a directory, use the ‘--no-overwrite-dir’ option.

To be even more cautious and prevent existing files from being replaced, use the ‘--keep-old-files’ (‘-k’) option. It causes tar to refuse to replace or update a file that already exists, i.e., a file with the same name as an archive member prevents extraction of that archive member. Instead, it reports an error.

To be more aggressive about altering existing files, use the ‘--overwrite’ option. It causes tar to overwrite existing files and to follow existing symbolic links when extracting.

Some people argue that GNU tar should not hesitate to overwrite files with other files when extracting. When extracting a tar archive, they expect to see a faithful copy of the state of the file system when the archive was created. It is debatable that this would always be a proper behavior. For example, suppose one has an archive in which ‘usr/local’ is a link to ‘usr/local2’. Since then, maybe the site removed the link and renamed the whole hierarchy from ‘/usr/local2’ to ‘/usr/local’. Such things happen all the time. I guess it would not be welcome at all that GNU tar removes the whole hierarchy just to make room for the link to be reinstated (unless it also simultaneously restores the full ‘/usr/local2’, of course!) GNU tar is indeed able to remove a whole hierarchy to reestablish a symbolic link, for example, but only if ‘--recursive-unlink’ is specified to allow this behavior. In any case, single files are silently removed.

Finally, the ‘--unlink-first’ (‘-U’) option can improve performance in some cases by causing tar to remove files unconditionally before extracting them.

Overwrite Old Files

‘--overwrite’

Overwrite existing files and directory metadata when extracting files from an archive. This causes tar to write extracted files into the file system without regard to the files already on the system; i.e., files with the same names as archive members are overwritten when the archive is extracted. It also causes tar to extract the ownership, permissions, and time stamps onto any preexisting files or directories. If the name of a corresponding file name is a symbolic link, the file pointed to by the symbolic link will be overwritten instead of the symbolic link itself (if this is possible). Moreover, special devices, empty directories and even symbolic links are automatically removed if they are in the way of extraction. Be careful when using the ‘--overwrite’ option, particularly when combined with the ‘--absolute-names’ (‘-P’) option, as this combination can change the contents, ownership or permissions of any file on your system. Also, many systems do not take kindly to overwriting files that are currently being executed.

‘--overwrite-dir’

Overwrite the metadata of directories when extracting files from an archive, but remove other files before extracting.

Keep Old Files

‘--keep-old-files’

‘-k’  Do not replace existing files from archive.  The

‘--keep-old-files’ (‘-k’) option prevents tar from replacing existing files with files with the same name from the archive. The ‘--keep-old-files’ option is meaningless with ‘--list’ (‘-t’). Prevents tar from replacing files in the file system during extraction.

Keep Newer Files

‘--keep-newer-files’

Do not replace existing files that are newer than their archive copies. This option is meaningless with ‘--list’ (‘-t’).

Unlink First

‘--unlink-first’

‘-U’ Remove files before extracting over them. This can make tar run a bit faster if you know in advance that the extracted files all need to be removed. Normally this option slows tar down slightly, so it is disabled by default.

Recursive Unlink

‘--recursive-unlink’

When this option is specified, try removing files and directory hierarchies before extracting over them. This is a dangerous option!

If you specify the ‘--recursive-unlink’ option, tar removes anything that keeps you from extracting a file as far as current permissions will allow it. This could include removal of the contents of a full directory hierarchy.

Setting Data Modification Times

Normally, tar sets the data modification times of extracted files to the corresponding times recorded for the files in the archive, but limits the permissions of extracted files by the current umask setting.

To set the data modification times of extracted files to the time when the files were extracted, use the ‘--touch’ (‘-m’) option in conjunction with

‘--extract’ (‘--get’, ‘-x’).

‘--touch’

‘-m’ Sets the data modification time of extracted archive members to the time they were extracted, not the time recorded for them in the archive. Use in conjunction with ‘--extract’ (‘--get’, ‘-x’).

Setting Access Permissions

To set the modes (access permissions) of extracted files to those recorded for those files in the archive, use ‘--same-permissions’ in conjunction with the ‘--extract’ (‘--get’, ‘-x’) operation.

‘--preserve-permissions’
‘--same-permissions’

‘-p’ Set modes of extracted archive members to those recorded in the archive, instead of current umask settings. Use in conjunction

          with ‘--extract’ (‘--get’, ‘-x’).

Directory Modification Times and Permissions

After successfully extracting a file member, GNU tar normally restores its permissions and modification times, as described in the previous sections. This cannot be done for directories, because after extracting a directory tar will almost certainly extract files into that directory and this will cause the directory modification time to be updated. Moreover, restoring that directory permissions may not permit file creation within it. Thus, restoring directory permissions and modification times must be delayed at least until all files have been extracted into that directory. GNU tar restores directories using the following approach.

The extracted directories are created with the mode specified in the archive, as modified by the umask of the user, which gives sufficient permissions to allow file creation. The meta-information about the directory is recorded in the temporary list of directories. When preparing to extract next archive member, GNU tar checks if the directory prefix of this file contains the remembered directory. If it does not, the program assumes that all files have been extracted into that directory, restores its modification time and permissions and removes its entry from the internal list. This approach allows to correctly restore directory meta-information in the majority of cases, while keeping memory requirements sufficiently small. It is based on the fact, that most tar archives use the predefined order of members: first the directory, then all the files and subdirectories in that directory.

However, this is not always true. The most important exception are incremental archives (see Section 5.2 [Incremental Dumps], page 74). The member order in an incremental archive is reversed: first all directory members are stored, followed by other (non-directory) members. So, when extracting from incremental archives, GNU tar alters the above procedure. It remembers all restored directories, and restores their meta-data only after the entire archive has been processed. Notice, that you do not need to specify any special options for that, as GNU tar automatically detects archives in incremental format.

There may be cases, when such processing is required for normal archives too. Consider the following example:

     $ tar --no-recursion -cvf archive \
         foo foo/file1 bar bar/file foo/file2
     foo/
     foo/file1
     bar/
     bar/file
     foo/file2

During the normal operation, after encountering ‘bar’ GNU tar will assume that all files from the directory ‘foo’ were already extracted and will therefore restore its timestamp and permission bits. However, after extracting ‘foo/file2’ the directory timestamp will be offset again.

To correctly restore directory meta-information in such cases, use

‘delay-directory-restore’ command line option:
‘--delay-directory-restore’

Delays restoring of the modification times and permissions of extracted directories until the end of extraction. This way, correct meta-information is restored even if the archive has unusual member ordering.

‘--no-delay-directory-restore’

Cancel the effect of the previous ‘--delay-directory-restore’. Use this option if you have used ‘--delay-directory-restore’ in TAR_OPTIONS variable (see [TAR OPTIONS], page 23) and wish to temporarily disable it.

Writing to Standard Output

To write the extracted files to the standard output, instead of creating the files on the file system, use ‘--to-stdout’ (‘-O’) in conjunction with ‘--extract’ (‘--get’, ‘-x’). This option is useful if you are extracting files to send them through a pipe, and do not need to preserve them in the file system. If you extract multiple members, they appear on standard output concatenated, in the order they are found in the archive.

‘--to-stdout’

‘-O’ Writes files to the standard output. Use only in conjunction with ‘--extract’ (‘--get’, ‘-x’). When this option is used, instead of creating the files specified, tar writes the contents of the files extracted to its standard output. This may be useful if you are only extracting the files in order to send them through a pipe. This option is meaningless with ‘--list’ (‘-t’).

This can be useful, for example, if you have a tar archive containing a big file and don’t want to store the file on disk before processing it. You can use a command like this:

tar -xOzf foo.tgz bigfile | process

or even like this if you want to process the concatenation of the files:

tar -xOzf foo.tgz bigfile1 bigfile2 | process

However, ‘--to-command’ may be more convenient for use with multiple files. See the next section.

Writing to an External Program

You can instruct tar to send the contents of each extracted file to the standard input of an external program:

‘--to-command=command’

Extract files and pipe their contents to the standard input of command. When this option is used, instead of creating the files specified, tar invokes command and pipes the contents of the files to its standard output. Command may contain command line arguments. The program is executed via sh -c. Notice, that command is executed once for each regular file extracted. Non-regular files (directories, etc.) are ignored when this option is used.

The command can obtain the information about the file it processes from the following environment variables: TAR FILETYPE

Type of the file. It is a single letter with the following meaning:

f  Regular file
d  Directory
l  Symbolic link
h  Hard link
b  Block device
c  Character device

Currently only regular files are supported.

TAR MODE File mode, an octal number.

TAR FILENAME The name of the file.

TAR REALNAME Name of the file as stored in the archive.

TAR UNAME Name of the file owner.

TAR GNAME Name of the file owner group.

TAR ATIME Time of last access. It is a decimal number, representing seconds since the epoch. If the archive provides times with nanosecond precision, the nanoseconds are appended to the timestamp after a decimal point.

TAR MTIME Time of last modification.

TAR CTIME Time of last status change.

TAR SIZE Size of the file.

TAR UID UID of the file owner.

TAR GID GID of the file owner.

In addition to these variables, TAR_VERSION contains the GNU tar version number.

If command exits with a non-0 status, tar will print an error message similar to the following:

tar: 2345: Child returned status 1

Here, ‘2345’ is the PID of the finished process. If this behavior is not wanted, use ‘--ignore-command-error’:

‘--ignore-command-error’

Ignore exit codes of subprocesses. Notice that if the program exits on signal or otherwise terminates abnormally, the error message will be printed even if this option is used.

‘--no-ignore-command-error’

Cancel the effect of any previous ‘--ignore-command-error’ option. This option is useful if you have set

‘--ignore-command-error’  in  TAR_OPTIONS  (see

[TAR OPTIONS], page 23) and wish to temporarily cancel it.

Removing Files

‘--remove-files’

Remove files after adding them to the archive.

4.4.3 Coping with Scarce Resources

(This message will disappear, once this node revised.)

Starting File

‘--starting-file=name’

‘-K name’ Starts an operation in the middle of an archive. Use in conjunction with ‘--extract’ (‘--get’, ‘-x’) or ‘--list’ (‘-t’).

If a previous attempt to extract files failed due to lack of disk space, you can use ‘--starting-file=name’ (‘-K name’) to start extracting only after member name of the archive. This assumes, of course, that there is now free space, or that you are now extracting into a different file system. (You could also choose to suspend tar, remove unnecessary files from the file system, and then restart the same tar operation. In this case, ‘--starting-file’ is not necessary. See Section 5.2 [Incremental Dumps], page 74, See Section 3.8 [interactive], page 48, and Section 6.4 [exclude], page 91.)

Same Order

‘--same-order’
‘--preserve-order’

‘-s’ To process large lists of file names on machines with small amounts of memory. Use in conjunction with ‘--compare’

          (‘--diff’, ‘-d’), ‘--list’ (‘-t’) or ‘--extract’ (‘--get’, ‘-x’).
   The ‘--same-order’ (‘--preserve-order’, ‘-s’) option tells tar that the

list of file names to be listed or extracted is sorted in the same order as the files in the archive. This allows a large list of names to be used, even on a small machine that would not otherwise be able to hold all the names in memory at the same time. Such a sorted list can easily be created by running ‘tar -t’ on the archive and editing its output.

This option is probably never needed on modern computer systems.

4.5 Backup options

GNU tar offers options for making backups of files before writing new versions. These options control the details of these backups. They may apply to the archive itself before it is created or rewritten, as well as individual extracted members. Other GNU programs (cp, install, ln, and mv, for example) offer similar options.

Backup options may prove unexpectedly useful when extracting archives containing many members having identical name, or when extracting archives on systems having file name limitations, making different members appear has having similar names through the side-effect of name truncation. (This is true only if we have a good scheme for truncated backup names, which I’m not sure at all: I suspect work is needed in this area.) When any existing file is backed up before being overwritten by extraction, then clashing files are automatically be renamed to be unique, and the true name is kept for only the last file of a series of clashing files. By using verbose mode, users may track exactly what happens.

At the detail level, some decisions are still experimental, and may change in the future, we are waiting comments from our users. So, please do not learn to depend blindly on the details of the backup features. For example, currently, directories themselves are never renamed through using these options, so, extracting a file over a directory still has good chances to fail. Also, backup options apply to created archives, not only to extracted members. For created archives, backups will not be attempted when the archive is a block or character device, or when it refers to a remote file.

For the sake of simplicity and efficiency, backups are made by renaming old files prior to creation or extraction, and not by copying. The original name is restored if the file creation fails. If a failure occurs after a partial extraction of a file, both the backup and the partially extracted file are kept.

‘--backup[=method]’

Back up files that are about to be overwritten or removed. Without this option, the original versions are destroyed. Use method to determine the type of backups made. If method is not specified, use the value of the VERSION_CONTROL environment variable. And if VERSION_CONTROL is not set, use the ‘existing’ method. This option corresponds to the Emacs variable ‘version-control’; the same values for method are accepted as in Emacs. This option also allows more descriptive names. The valid methods are:

‘t’

‘numbered’

Always make numbered backups.

‘nil’
‘existing’

Make numbered backups of files that already have them, simple backups of the others.

‘never’

‘simple’ Always make simple backups.

‘--suffix=suffix’

Append suffix to each backup file made with ‘--backup’. If this option is not specified, the value of the SIMPLE_BACKUP_SUFFIX environment variable is used. And if SIMPLE_BACKUP_SUFFIX is not set, the default is ‘~’, just as in Emacs.

4.6 Notable tar Usages

(This message will disappear, once this node revised.) You can easily use archive files to transport a group of files from one

system to another: put all relevant files into an archive on one computer system, transfer the archive to another system, and extract the contents there. The basic transfer medium might be magnetic tape, Internet FTP, or even electronic mail (though you must encode the archive with uuencode in order to transport it properly by mail). Both machines do not have to use the same operating system, as long as they both support the tar program.

For example, here is how you might copy a directory’s contents from one disk to another, while preserving the dates, modes, owners and link-structure of all the files therein. In this case, the transfer medium is a pipe, which is one a Unix redirection mechanism:

     $ (cd sourcedir; tar -cf - .) | (cd targetdir; tar -xf -)

You can avoid subshells by using ‘-C’ option:

     $ tar -C sourcedir -cf - . | tar -C targetdir -xf -

The command also works using short option forms:

     $ (cd sourcedir; tar --create --file=- . ) \
            | (cd targetdir; tar --extract --file=-)
     # Or:
     $ tar --directory sourcedir --create --file=- . ) \
            | tar --directory targetdir --extract --file=-

This is one of the easiest methods to transfer a tar archive.

4.7 Looking Ahead: The Rest of this Manual

You have now seen how to use all eight of the operations available to tar, and a number of the possible options. The next chapter explains how to choose and change file and archive names, how to use files to store names of other files which you can then call as arguments to tar (this can help you save time if you expect to archive the same list of files a number of times), and so forth.

If there are too many files to conveniently list on the command line, you can list the names in a file, and tar will read that file. See Section 6.3 [files], page 89.

There are various ways of causing tar to skip over some files, and not archive them. See Chapter 6 [Choosing], page 87.

72 GNU tar 1.16.1

5 Performing Backups and Restoring Files

(This message will disappear, once this node revised.)

GNU tar is distributed along with the scripts which the Free Software Foundation uses for performing backups. There is no corresponding scripts available yet for doing restoration of files. Even if there is a good chance those scripts may be satisfying to you, they are not the only scripts or methods available for doing backups and restore. You may well create your own, or use more sophisticated packages dedicated to that purpose.

Some users are enthusiastic about Amanda (The Advanced Maryland Automatic Network Disk Archiver), a backup system developed by James da Silva ‘jds@cs.umd.edu’ and available on many Unix systems. This is free software, and it is available at these places:

     http://www.cs.umd.edu/projects/amanda/amanda.html
     ftp://ftp.cs.umd.edu/pub/amanda

This chapter documents both the provided shell scripts and tar options which are more specific to usage as a backup tool.

To back up a file system means to create archives that contain all the files in that file system. Those archives can then be used to restore any or all of those files (for instance if a disk crashes or a file is accidentally deleted). File system backups are also called dumps.

5.1 Using tar to Perform Full Dumps

(This message will disappear, once this node revised.)

Full dumps should only be made when no other people or programs are modifying files in the file system. If files are modified while tar is making the backup, they may not be stored properly in the archive, in which case you won’t be able to restore them if you have to. (Files not being modified are written with no trouble, and do not corrupt the entire archive.)

   You will want to use the ‘--label=archive-label’ (‘-V archive-

label’) option to give the archive a volume label, so you can tell what this archive is even if the label falls off the tape, or anything like that.

Unless the file system you are dumping is guaranteed to fit on one volume, you will need to use the ‘--multi-volume’ (‘-M’) option. Make sure you have enough tapes on hand to complete the backup.

If you want to dump each file system separately you will need to use the ‘--one-file-system’ option to prevent tar from crossing file system boundaries when storing (sub)directories.

The ‘--incremental’ (‘-G’) (see Section 5.2 [Incremental Dumps], page 74) option is not needed, since this is a complete copy of everything in the file system, and a full restore from this backup would only be done onto a completely empty disk.

Unless you are in a hurry, and trust the tar program (and your tapes), it is a good idea to use the ‘--verify’ (‘-W’) option, to make sure your files really made it onto the dump properly. This will also detect cases where the file was modified while (or just after) it was being archived. Not all media (notably cartridge tapes) are capable of being verified, unfortunately.

5.2 Using tar to Perform Incremental Dumps

Incremental backup is a special form of GNU tar archive that stores additional metadata so that exact state of the file system can be restored when extracting the archive.

GNU tar currently offers two options for handling incremental back-

ups: ‘--listed-incremental=snapshot-file’ (‘-g snapshot-file’) and
‘--incremental’ (‘-G’).

The option ‘--listed-incremental’ instructs tar to operate on an incremental archive with additional metadata stored in a standalone file, called a snapshot file. The purpose of this file is to help determine which files have been changed, added or deleted since the last backup, so that the next incremental backup will contain only modified files. The name of the snapshot file is given as an argument to the option:

‘--listed-incremental=file’

‘-g file’ Handle incremental backups with snapshot data in file.

To create an incremental backup, you would use ‘--listed-incremental’ together with ‘--create’ (see Section 2.6 [create], page 10). For example:

     $ tar --create \
                --file=archive.1.tar \
                --listed-incremental=/var/log/usr.snar \
                /usr

This will create in ‘archive.1.tar’ an incremental backup of the ‘/usr’ file system, storing additional metadata in the file ‘/var/log/usr.snar’. If this file does not exist, it will be created. The created archive will then be a level 0 backup; please see the next section for more on backup levels.

Otherwise, if the file ‘/var/log/usr.snar’ exists, it determines which files are modified. In this case only these files will be stored in the archive. Suppose, for example, that after running the above command, you delete file ‘/usr/doc/old’ and create directory ‘/usr/local/db’ with the following contents:

     $ ls /usr/local/db
     /usr/local/db/data
     /usr/local/db/index

Some time later you create another incremental backup. You will then see:

     $ tar --create \
                --file=archive.2.tar \

                --listed-incremental=/var/log/usr.snar \
                /usr
     tar: usr/local/db: Directory is new
     usr/local/db/
     usr/local/db/data
     usr/local/db/index

The created archive ‘archive.2.tar’ will contain only these three members. This archive is called a level 1 backup. Notice that ‘/var/log/usr.snar’ will be updated with the new data, so if you plan to create more ‘level 1’ backups, it is necessary to create a working copy of the snapshot file before running tar. The above example will then be modified as follows:

     $ cp /var/log/usr.snar /var/log/usr.snar-1
     $ tar --create \
                --file=archive.2.tar \
                --listed-incremental=/var/log/usr.snar-1 \
                /usr

Incremental dumps depend crucially on time stamps, so the results are unreliable if you modify a file’s time stamps during dumping (e.g., with the ‘--atime-preserve=replace’ option), or if you set the clock backwards.

Metadata stored in snapshot files include device numbers, which, obviously is supposed to be a non-volatile value. However, it turns out that NFS devices have undependable values when an automounter gets in the picture. This can lead to a great deal of spurious redumping in incremental dumps, so it is somewhat useless to compare two NFS devices numbers over time. The solution implemented currently is to considers all NFS devices as being equal when it comes to comparing directories; this is fairly gross, but there does not seem to be a better way to go.

Note that incremental archives use tar extensions and may not be readable by non-GNU versions of the tar program.

To extract from the incremental dumps, use ‘--listed-incremental’ together with ‘--extract’ option (see Section 2.8.2 [extracting files], page 17). In this case, tar does not need to access snapshot file, since all the data necessary for extraction are stored in the archive itself. So, when extracting, you can give whatever argument to ‘--listed-incremental’, the usual practice is to use ‘--listed-incremental=/dev/null’. Alternatively, you can use ‘--incremental’, which needs no arguments. In general, ‘--incremental’ (‘-G’) can be used as a shortcut for ‘--listed-incremental’ when listing or extracting incremental backups (for more information, regarding this option, see [incremental-op], page 76).

When extracting from the incremental backup GNU tar attempts to restore the exact state the file system had when the archive was created. In particular, it will delete those files in the file system that did not exist in their directories when the archive was created. If you have created several levels of incremental files, then in order to restore the exact contents the file system had when the last level was created, you will need to restore from all backups in turn. Continuing our example, to restore the state of ‘/usr’ file system, one would do^{1}:

     $ tar --extract \
                --listed-incremental=/dev/null \
                --file archive.1.tar
     $ tar --extract \
                --listed-incremental=/dev/null \
                --file archive.2.tar

To list the contents of an incremental archive, use ‘--list’ (see Section 2.7 [list], page 15), as usual. To obtain more information about the

archive, use ‘--listed-incremental’ or ‘--incremental’ combined with

two ‘--verbose’ options^{2}:

     tar --list --incremental --verbose --verbose archive.tar

This command will print, for each directory in the archive, the list of files in that directory at the time the archive was created. This information is put out in a format which is both human-readable and unambiguous for a program: each file name is printed as

     x file

where x is a letter describing the status of the file: ‘Y’ if the file is present in the archive, ‘N’ if the file is not included in the archive, or a ‘D’ if the file is a directory (and is included in the archive). See [Dumpdir], page 178, for the detailed description of dumpdirs and status codes. Each such line is terminated by a newline character. The last line is followed by an additional newline to indicate the end of the data.

The option ‘--incremental’ (‘-G’) gives the same behavior as

‘--listed-incremental’ when used with ‘--list’ and ‘--extract’

options. When used with ‘--create’ option, it creates an incremental archive without creating snapshot file. Thus, it is impossible to create several levels of incremental backups with ‘--incremental’ option.

5.3 Levels of Backups

An archive containing all the files in the file system is called a full backup or full dump. You could insure your data by creating a full dump every day. This strategy, however, would waste a substantial amount of archive media and user time, as unchanged files are daily re-archived.

1 Notice, that since both archives were created without ‘-P’ option (see Section 6.10.2

[absolute], page 106), these commands should be run from the root file system.

2 Two ‘--verbose’ options were selected to avoid breaking usual verbose listing output

(‘--list --verbose’) when using in scripts.

Versions of GNU tar up to 1.15.1 used to dump verbatim binary contents

of the DUMPDIR header (with terminating nulls) when ‘--incremental’ or

‘--listed-incremental’ option was given, no matter what the verbosity level. This

behavior, and, especially, the binary output it produced were considered inconvenient

and were changed in version 1.16

It is more efficient to do a full dump only occasionally. To back up files between full dumps, you can use incremental dumps. A level one dump archives all the files that have changed since the last full dump.

A typical dump strategy would be to perform a full dump once a week, and a level one dump once a day. This means some versions of files will in fact be archived more than once, but this dump strategy makes it possible to restore a file system to within one day of accuracy by only extracting two archives—the last weekly (full) dump and the last daily (level one) dump. The only information lost would be in files changed or created since the last daily backup. (Doing dumps more than once a day is usually not worth the trouble).

GNU tar comes with scripts you can use to do full and level-one (actually, even level-two and so on) dumps. Using scripts (shell programs) to perform backups and restoration is a convenient and reliable alternative to typing out file name lists and tar commands by hand.

Before you use these scripts, you need to edit the file ‘backup-specs’, which specifies parameters used by the backup scripts and by the restore script. This file is usually located in ‘/etc/backup’ directory. See Section 5.4 [Backup Parameters], page 77, for its detailed description. Once the backup parameters are set, you can perform backups or restoration by running the appropriate script.

The name of the backup script is backup. The name of the restore script is restore. The following sections describe their use in detail.

Please Note: The backup and restoration scripts are designed to be used together. While it is possible to restore files by hand from an archive which was created using a backup script, and to create an archive by hand which could then be extracted using the restore script, it is easier to use the scripts. See Section 5.2 [Incremental Dumps], page 74, before making such an attempt.

5.4 Setting Parameters for Backups and Restoration

The file ‘backup-specs’ specifies backup parameters for the backup and restoration scripts provided with tar. You must edit ‘backup-specs’ to fit your system configuration and schedule before using these scripts.

Syntactically, ‘backup-specs’ is a shell script, containing mainly variable assignments. However, any valid shell construct is allowed in this file. Particularly, you may wish to define functions within that script (e.g., see RESTORE_BEGIN below). For more information about shell script syntax, please refer to the definition of the Shell Command Language

(http://www.opengroup.org/onlinepubs/009695399/utilities/xcu_

chap02.html#ta g_02). See also section “Bash Features” in Bash Reference Manual.

The shell variables controlling behavior of backup and restore are described in the following subsections.

5.4.1 General-Purpose Variables

[Backup variable] ADMINISTRATOR

The user name of the backup administrator. Backup scripts sends a backup report to this address.

[Backup variable] BACKUP_HOUR

The hour at which the backups are done. This can be a number from 0 to 23, or the time specification in form hours:minutes, or the string ‘now’. This variable is used by backup. Its value may be overridden using ‘--time’ option (see Section 5.5 [Scripted Backups], page 82).

[Backup variable] TAPE_FILE

The device tar writes the archive to. If TAPE FILE is a remote archive (see [remote-dev], page 88), backup script will suppose that your mt is able to access remote devices. If RSH (see [RSH], page 79) is set, ‘--rsh-command’ option will be added to invocations of mt.

[Backup variable] BLOCKING

The blocking factor tar will use when writing the dump archive. See Section 9.4.2 [Blocking Factor], page 143.

[Backup variable] BACKUP_DIRS

A list of file systems to be dumped (for backup), or restored (for restore). You can include any directory name in the list — subdirectories on that file system will be included, regardless of how they may look to other networked machines. Subdirectories on other file systems will be ignored. The host name specifies which host to run tar on, and should normally be the host that actually contains the file system. However, the host machine must have GNU tar installed, and must be able to access the directory containing the backup scripts and their support files using the same file name that is used on the machine where the scripts are run (i.e., what pwd will print when in that directory on that machine). If the host that contains the file system does not have this capability, you can specify another host as long as it can access the file system through NFS. If the list of file systems is very long you may wish to put it in a separate file. This file is usually named ‘/etc/backup/dirs’, but this name may be overridden in ‘backup-specs’ using DIRLIST variable.

[Backup variable] DIRLIST

A path to the file containing the list of the file systems to backup or restore. By default it is ‘/etc/backup/dirs’.

[Backup variable] BACKUP_FILES

A list of individual files to be dumped (for backup), or restored (for restore). These should be accessible from the machine on which the backup script is run. If the list of file systems is very long you may wish to store it in a separate file. This file is usually named ‘/etc/backup/files’, but this name may be overridden in ‘backup-specs’ using FILELIST variable.

[Backup variable] FILELIST

A path to the file containing the list of the individual files to backup or restore. By default it is ‘/etc/backup/files’.

[Backup variable] MT

Full file name of mt binary.

[Backup variable] RSH

Full file name of rsh binary or its equivalent. You may wish to set it to ssh, to improve security. In this case you will have to use public key authentication.

[Backup variable] RSH_COMMAND

Full file name of rsh binary on remote machines. This will be passed via ‘--rsh-command’ option to the remote invocation of GNU tar.

[Backup variable] VOLNO_FILE

Name of temporary file to hold volume numbers. This needs to be accessible by all the machines which have file systems to be dumped.

[Backup variable] XLIST

Name of exclude file list. An exclude file list is a file located on the remote machine and containing the list of files to be excluded from the backup. Exclude file lists are searched in /etc/tar-backup directory. A common use for exclude file lists is to exclude files containing security-sensitive information (e.g., ‘/etc/shadow’ from backups). This variable affects only backup.

[Backup variable] SLEEP_TIME

Time to sleep between dumps of any two successive file systems This variable affects only backup.

[Backup variable] DUMP_REMIND_SCRIPT

Script to be run when it’s time to insert a new tape in for the next volume. Administrators may want to tailor this script for their site. If this variable isn’t set, GNU tar will display its built-in prompt, and will expect confirmation from the console. For the description of the default prompt, see [change volume prompt], page 151.

[Backup variable] SLEEP_MESSAGE

Message to display on the terminal while waiting for dump time. Usually this will just be some literal text.

[Backup variable] TAR

Full file name of the GNU tar executable. If this is not set, backup scripts will search tar in the current shell path.

5.4.2 Magnetic Tape Control

Backup scripts access tape device using special hook functions. These functions take a single argument – the name of the tape device. Their names are kept in the following variables:

[Backup variable] MT_BEGIN

The name of begin function. This function is called before accessing the drive. By default it retensions the tape:

     MT_BEGIN=mt_begin
     mt_begin() {
         mt -f "$1" retension
     }

[Backup variable] MT_REWIND

The name of rewind function. The default definition is as follows:

     MT_REWIND=mt_rewind
     mt_rewind() {
         mt -f "$1" rewind
     }

[Backup variable] MT_OFFLINE

The name of the function switching the tape off line. By default it is defined as follows:

     MT_OFFLINE=mt_offline
     mt_offline() {
         mt -f "$1" offl
     }

[Backup variable] MT_STATUS

The name of the function used to obtain the status of the archive device, including error count. Default definition:

     MT_STATUS=mt_status
     mt_status() {
         mt -f "$1" status
     }

5.4.3 User Hooks

User hooks are shell functions executed before and after each tar invocation. Thus, there are backup hooks, which are executed before and after dumping each file system, and restore hooks, executed before and after restoring a file system. Each user hook is a shell function taking four arguments:

[User Hook Function] hook level host fs fsname

Its arguments are:

level Current backup or restore level.

host Name or IP address of the host machine being dumped or restored.

fs Full path name to the file system being dumped or restored.

fsname File system name with directory separators replaced with colons. This is useful, e.g., for creating unique files.

Following variables keep the names of user hook functions

[Backup variable] DUMP_BEGIN

Dump begin function. It is executed before dumping the file system.

[Backup variable] DUMP_END

Executed after dumping the file system.

[Backup variable] RESTORE_BEGIN

Executed before restoring the file system.

[Backup variable] RESTORE_END

Executed after restoring the file system.

5.4.4 An Example Text of ‘Backup-specs’

The following is an example of ‘backup-specs’:

  # site-specific parameters for file system backup.
  ADMINISTRATOR=friedman
  BACKUP_HOUR=1
  TAPE_FILE=/dev/nrsmt0
  # Use ssh instead of the less secure rsh
  RSH=/usr/bin/ssh
  RSH_COMMAND=/usr/bin/ssh
  # Override MT_STATUS function:
  my_status() {
        mts -t $TAPE_FILE
  }
  MT_STATUS=my_status
  # Disable MT_OFFLINE function
  MT_OFFLINE=:
  BLOCKING=124
  BACKUP_DIRS="
          albert:/fs/fsf
          apple-gunkies:/gd

             albert:/fs/gd2
             albert:/fs/gp
             geech:/usr/jla
             churchy:/usr/roland
             albert:/
             albert:/usr
             apple-gunkies:/
             apple-gunkies:/usr
             gnu:/hack
             gnu:/u
             apple-gunkies:/com/mailer/gnu
             apple-gunkies:/com/archive/gnu"
     BACKUP_FILES="/com/mailer/aliases /com/mailer/league*[a-z]"

5.5 Using the Backup Scripts

The syntax for running a backup script is:

     backup --level=level --time=time

The ‘level’ option requests the dump level. Thus, to produce a full dump, specify --level=0 (this is the default, so ‘--level’ may be omitted if its value is 0).^{3}

The ‘--time’ option determines when should the backup be run. Time may take three forms:

hh:mm The dump must be run at hh hours mm minutes.

hh The dump must be run at hh hours

now The dump must be run immediately.

You should start a script with a tape or disk mounted. Once you start a script, it prompts you for new tapes or disks as it needs them. Media volumes don’t have to correspond to archive files — a multi-volume archive can be started in the middle of a tape that already contains the end of another multi-volume archive. The restore script prompts for media by its archive volume, so to avoid an error message you should keep track of which tape (or disk) contains which volume of the archive (see Section 5.6 [Scripted Restoration], page 83).

3 For backward compatibility, the backup will also try to deduce the requested dump

level from the name of the script itself. If the name consists of a string ‘level-’ followed

by a single decimal digit, that digit is taken as the dump level number. Thus, you may

create a link from backup to level-1 and then run level-1 whenever you need to

create a level one dump.

The backup scripts write two files on the file system. The first is a record file in ‘/etc/tar-backup/’, which is used by the scripts to store and retrieve information about which files were dumped. This file is not meant to be read by humans, and should not be deleted by them. See [Snapshot Files], page 177, for a more detailed explanation of this file.

The second file is a log file containing the names of the file systems and files dumped, what time the backup was made, and any error messages that were generated, as well as how much space was left in the media volume after the last volume of the archive was written. You should check this log file after every backup. The file name is ‘log-mm-dd-yyyy-level-n’, where mm-dd-yyyy represents current date, and n represents current dump level number.

The script also prints the name of each system being dumped to the standard output.

Following is the full list of options accepted by backup script:

‘-l level’
‘--level=level’

Do backup level level (default 0).

‘-f’

‘--force’ Force backup even if today’s log file already exists.

‘-v[level]’
‘--verbose[=level]’

Set verbosity level. The higher the level is, the more debugging information will be output during execution. Default level is 100, which means the highest debugging level.

‘-t start-time’
‘--time=start-time’

Wait till time, then do backup.

‘-h’

‘--help’ Display short help message and exit.

‘-V’
‘--version’

Display information about the program’s name, version, origin and legal status, all on standard output, and then exit successfully.

5.6 Using the Restore Script

To restore files that were archived using a scripted backup, use the restore script. Its usage is quite straightforward. In the simplest form, invoke restore --all, it will then restore all the file systems and files specified in ‘backup-specs’ (see Section 5.4.1 [General-Purpose Variables], page 78).

You may select the file systems (and/or files) to restore by giving restore list of patterns in its command line. For example, running

     restore 'albert:*'

will restore all file systems on the machine ‘albert’. A more complicated example:

     restore 'albert:*' '*:/var'

This command will restore all file systems on the machine ‘albert’ as well as ‘/var’ file system on all machines.

By default restore will start restoring files from the lowest available dump level (usually zero) and will continue through all available dump levels. There may be situations where such a thorough restore is not necessary. For example, you may wish to restore only files from the recent level one backup. To do so, use ‘--level’ option, as shown in the example below:

     restore --level=1

The full list of options accepted by restore follows:

‘-a’

‘--all’ Restore all file systems and files specified in ‘backup-specs’

‘-l level’
‘--level=level’

Start restoring from the given backup level, instead of the default 0.

‘-v[level]’
‘--verbose[=level]’

Set verbosity level. The higher the level is, the more debugging information will be output during execution. Default level is 100, which means the highest debugging level.

‘-h’

‘--help’ Display short help message and exit.

‘-V’
‘--version’

Display information about the program’s name, version, origin and legal status, all on standard output, and then exit successfully.

You should start the restore script with the media containing the first volume of the archive mounted. The script will prompt for other volumes as they are needed. If the archive is on tape, you don’t need to rewind the tape to to its beginning—if the tape head is positioned past the beginning of the archive, the script will rewind the tape as needed. See Section 9.5.1 [Tape Positioning], page 149, for a discussion of tape positioning.

Warning: The script will delete files from the active file system if they were not in the file system when the archive was made.

Chapter 5: Performing Backups and Restoring Files 85

See Section 5.2 [Incremental Dumps], page 74, for an explanation of how the script makes that determination.

86 GNU tar 1.16.1

6 Choosing Files and Names for tar

(This message will disappear, once this node revised.)

Certain options to tar enable you to specify a name for your archive. Other options let you decide which files to include or exclude from the archive, based on when or whether files were modified, whether the file names do or don’t match specified patterns, or whether files are in specified directories.

This chapter discusses these options in detail.

6.1 Choosing and Naming Archive Files

(This message will disappear, once this node revised.)

By default, tar uses an archive file name that was compiled when it was built on the system; usually this name refers to some physical tape drive on the machine. However, the person who installed tar on the system may not have set the default to a meaningful value as far as most users are concerned. As a result, you will usually want to tell tar where to find (or create) the

archive. The ‘--file=archive-name’ (‘-f archive-name’) option allows

you to either specify or name a file to use as the archive instead of the default archive file location.

‘--file=archive-name’
‘-f archive-name’

Name the archive to create or operate on. Use in conjunction with any operation.

For example, in this tar command,

     $ tar -cvf collection.tar blues folk jazz

‘collection.tar’ is the name of the archive. It must directly follow the ‘-f’ option, since whatever directly follows ‘-f’ will end up naming the archive. If you neglect to specify an archive name, you may end up overwriting a file in the working directory with the archive you create since tar will use this file’s name for the archive name.

An archive can be saved as a file in the file system, sent through a pipe or over a network, or written to an I/O device such as a tape, floppy disk, or CD write drive.

If you do not name the archive, tar uses the value of the environment variable TAPE as the file name for the archive. If that is not available, tar uses a default, compiled-in archive name, usually that for tape unit zero (i.e.,

‘/dev/tu00’).

If you use ‘-’ as an archive-name, tar reads the archive from standard input (when listing or extracting files), or writes it to standard output (when creating an archive). If you use ‘-’ as an archive-name when modifying an archive, tar reads the original archive from its standard input and writes the entire new archive to its standard output.

The following example is a convenient way of copying directory hierarchy

from ‘sourcedir’ to ‘targetdir’.
     $ (cd sourcedir; tar -cf - .) | (cd targetdir; tar -xpf -)

The ‘-C’ option allows to avoid using subshells:

     $ tar -C sourcedir -cf - . | tar -C targetdir -xpf -

In both examples above, the leftmost tar invocation archives the contents of ‘sourcedir’ to the standard output, while the rightmost one reads this archive from its standard input and extracts it. The ‘-p’ option tells it to restore permissions of the extracted files.

To specify an archive file on a device attached to a remote machine, use the following:

     --file=hostname:/dev/file-name

tar will complete the remote connection, if possible, and prompt you for a

username and password. If you use ‘--file=@hostname:/dev/file-name’,

tar will complete the remote connection, if possible, using your username as the username on the remote machine.

If the archive file name includes a colon (‘:’), then it is assumed to be a file on another machine. If the archive file is ‘user@host:file’, then file is used on the host host. The remote host is accessed using the rsh program, with a username of user. If the username is omitted (along with the ‘@’ sign), then your user name will be used. (This is the normal rsh behavior.) It is necessary for the remote machine, in addition to permitting your rsh access, to have the ‘rmt’ program installed (This command is included in the GNU tar distribution and by default is installed under ‘prefix/libexec/rmt’, were prefix means your installation prefix). If you need to use a file whose name includes a colon, then the remote tape drive behavior can be inhibited by using the ‘--force-local’ option.

When the archive is being created to ‘/dev/null’, GNU tar tries to minimize input and output operations. The Amanda backup system, when used with GNU tar, has an initial sizing pass which uses this feature.

6.2 Selecting Archive Members

File Name arguments specify which files in the file system tar operates on, when creating or adding to an archive, or which archive members tar operates on, when reading or deleting from an archive. See Section 4.2.1 [Operations], page 52.

To specify file names, you can include them as the last arguments on the command line, as follows:

     tar operation [option1 option2 ...] [file name-1 file name-2 ...]

If a file name begins with dash (‘-’), precede it with ‘--add-file’ option to prevent it from being treated as an option.

By default GNU tar attempts to unquote each file or member name, replacing escape sequences according to the following table:

Escape  Replaced with
\a  Audible bell (ASCII 7)
\b  Backspace (ASCII 8)
\f  Form feed (ASCII 12)
\n  New line (ASCII 10)
\r  Carriage return (ASCII 13)
\t  Horizontal tabulation (ASCII 9)
\v  Vertical tabulation (ASCII 11)
\?  ASCII 127

\n

ASCII n (n should be an octal number of up to 3 digits)

A backslash followed by any other symbol is retained.

This default behavior is controlled by the following command line option:

‘--unquote’

Enable unquoting input file or member names (default).

‘--no-unquote’

Disable unquoting input file or member names. If you specify a directory name as a file name argument, all the files in

that directory are operated on by tar.

If you do not specify files, tar behavior differs depending on the operation mode as described below:

When tar is invoked with ‘--create’ (‘-c’), tar will stop immediately, reporting the following:

     $ tar cf a.tar
     tar: Cowardly refusing to create an empty archive
     Try `tar --help' or `tar --usage' for more information.

If you specify either ‘--list’ (‘-t’) or ‘--extract’ (‘--get’, ‘-x’), tar operates on all the archive members in the archive.

If run with ‘--diff’ option, tar will compare the archive with the contents of the current working directory.

If you specify any other operation, tar does nothing.

By default, tar takes file names from the command line. However, there are other ways to specify file or member names, or to modify the manner in which tar selects the files or members upon which to operate. In general, these methods work both for specifying the names of files and archive members.

6.3 Reading Names from a File

Instead of giving the names of files or archive members on the command line, you can put the names into a file, and then use the ‘--files-from=file-of-names’ (‘-T file-of-names’) option to tar. Give the name of the file which contains the list of files to include as the argument to ‘--files-from’. In the list, the file names should be separated by newlines. You will frequently use this option when you have generated the list of files to archive with the find utility.

‘--files-from=file-name’
‘-T file-name’

Get names to extract or create from file file-name.

If you give a single dash as a file name for ‘--files-from’, (i.e., you specify either --files-from=- or -T -), then the file names are read from standard input.

Unless you are running tar with ‘--create’, you can not use both --

files-from=- and --file=- (-f -) in the same command.

Any number of ‘-T’ options can be given in the command line.

The following example shows how to use find to generate a list of files smaller than 400K in length and put that list into a file called ‘small-files’. You can then use the ‘-T’ option to tar to specify the files from that file, ‘small-files’, to create the archive ‘little.tgz’. (The ‘-z’ option to tar compresses the archive with gzip; see Section 8.1.1 [gzip], page 118 for more information.)

     $ find .  -size -400 -print > small-files
     $ tar -c -v -z -T small-files -f little.tgz

In the file list given by ‘-T’ option, any file name beginning with ‘-’ character is considered a tar option and is processed accordingly.^{1} For example, the common use of this feature is to change to another directory by specifying ‘-C’ option:

     $ cat list
     -C/etc
     passwd
     hosts
     -C/lib
     libc.a
     $ tar -c -f foo.tar --files-from list

In this example, tar will first switch to ‘/etc’ directory and add files ‘passwd’ and ‘hosts’ to the archive. Then it will change to ‘/lib’ directory and will archive the file ‘libc.a’. Thus, the resulting archive ‘foo.tar’ will contain:

     $ tar tf foo.tar
     passwd
     hosts
     libc.a

Notice that the option parsing algorithm used with ‘-T’ is stricter than the one used by shell. Namely, when specifying option arguments, you should observe the following rules:

1 Versions of GNU tar up to 1.15.1 recognized only ‘-C’ option in file lists, and only if

the option and its argument occupied two consecutive lines.

• When using short (single-letter) option form, its argument must immediately follow the option letter, without any intervening whitespace. For example: -Cdir.

• When using long option form, the option argument must be separated from the option by a single equal sign. No whitespace is allowed on any side of the equal sign. For example: --directory=dir.

• For both short and long option forms, the option argument can be given

on the next line after the option name, e.g.:

         --directory dir

and

         -C dir

If you happen to have a file whose name starts with ‘-’, precede it with ‘--add-file’ option to prevent it from being recognized as an option. For

example: --add-file=--my-file.

6.3.1 NUL Terminated File Names

The ‘--null’ option causes ‘--files-from=file-of-names’ (‘-T file-of-

names’) to read file names terminated by a NUL instead of a newline, so files whose names contain newlines can be archived using ‘--files-from’.

‘--null’ Only consider NUL terminated file names, instead of files that terminate in a newline.

The ‘--null’ option is just like the one in GNU xargs and cpio, and is useful with the ‘-print0’ predicate of GNU find. In tar, ‘--null’ also disables special handling for file names that begin with dash.

This example shows how to use find to generate a list of files larger than 800K in length and put that list into a file called ‘long-files’. The ‘-print0’ option to find is just like ‘-print’, except that it separates files with a NUL rather than with a newline. You can then run tar with both the ‘--null’ and ‘-T’ options to specify that tar get the files from that file, ‘long-files’, to create the archive ‘big.tgz’. The ‘--null’ option to tar will cause tar to recognize the NUL separator between files.

     $ find .  -size +800 -print0 > long-files
     $ tar -c -v --null --files-from=long-files --file=big.tar

6.4 Excluding Some Files

(This message will disappear, once this node revised.)

To avoid operating on files whose names match a particular pattern, use

the ‘--exclude’ or ‘--exclude-from’ options.
‘--exclude=pattern’

Causes tar to ignore files that match the pattern.

The ‘--exclude=pattern’ option prevents any file or member whose name matches the shell wildcard (pattern) from being operated on. For example, to create an archive with all the contents of the directory ‘src’ except for files whose names end in ‘.o’, use the command ‘tar -cf src.tar

--exclude='*.o' src’.

You may give multiple ‘--exclude’ options.

‘--exclude-from=file’

‘-X file’ Causes tar to ignore files that match the patterns listed in file.

Use the ‘--exclude-from’ option to read a list of patterns, one per line, from file; tar will ignore files matching those patterns. Thus if tar is called as ‘tar -c -X foo .’ and the file ‘foo’ contains a single line ‘*.o’, no files whose names end in ‘.o’ will be added to the archive.

‘--exclude-caches’

Causes tar to ignore directories containing a cache directory tag.

When creating an archive, the ‘--exclude-caches’ option causes tar to exclude all directories that contain a cache directory tag. A cache directory tag is a short file with the well-known name ‘CACHEDIR.TAG’ and having a standard header specified in

http://www.brynosaurus.com/cachedir/spec.html. Various applica-

tions write cache directory tags into directories they use to hold regenerable, non-precious data, so that such data can be more easily excluded from backups.

Another option, ‘--exclude-tag’, provides a generalization of this concept. It takes a single argument, a file name to look for. Any directory that contains this file will be excluded from the dump.

‘--exclude-tag=file’

Causes tar to ignore directories containing file. Multiple ‘--exclude-tag’ options can be given.

For example:

     $ find dir
     dir
     dir/blues
     dir/jazz
     dir/folk
     dir/folk/tagfile
     $ tar -cf archive.tar --exclude-tag=tagfile -v
     dir/
     dir/blues
     dir/jazz
     ./tar: dir/folk/: contains a cache directory tag tagfile; not dumped
     $ tar -tf archive.tar
     dir/
     dir/blues
     dir/jazz

Problems with Using the exclude Options

Some users find ‘exclude’ options confusing. Here are some common pitfalls:

• The main operating mode of tar does not act on a path name explicitly listed on the command line if one of its file name components is excluded. In the example above, if you create an archive and exclude files that end with ‘*.o’, but explicitly name the file ‘dir.o/foo’ after all the options have been listed, ‘dir.o/foo’ will be excluded from the archive.

• You can sometimes confuse the meanings of ‘--exclude’ and ‘--exclude-from’. Be careful: use ‘--exclude’ when files to be excluded are given as a pattern on the command line. Use ‘--exclude-from’ to introduce the name of a file which contains a list of patterns, one per line; each of these patterns can exclude zero, one, or many files.

• When you use ‘--exclude=pattern’, be sure to quote the pattern parameter, so GNU tar sees wildcard characters like ‘*’. If you do not do this, the shell might expand the ‘*’ itself using files at hand, so tar might receive a list of files instead of one pattern, or none at all, making the command somewhat illegal. This might not correspond to what you want. For example, write:

     $ tar -c -f archive.tar --exclude '*.o' directory

rather than:

     # Wrong!
     $ tar -c -f archive.tar --exclude *.o directory

• You must use use shell syntax, or globbing, rather than regexp syntax, when using exclude options in tar. If you try to use regexp syntax to describe files to be excluded, your command might fail.

• In earlier versions of tar, what is now the ‘--exclude-from’ option was called ‘--exclude’ instead. Now, ‘--exclude’ applies to patterns listed on the command line and ‘--exclude-from’ applies to patterns listed in a file.

6.5 Wildcards Patterns and Matching

Globbing is the operation by which wildcard characters, ‘*’ or ‘?’ for example, are replaced and expanded into all existing files matching the given pattern. GNU tar can use wildcard patterns for matching (or globbing) archive members when extracting from or listing an archive. Wildcard patterns are also used for verifying volume labels of tar archives. This section has the purpose of explaining wildcard syntax for tar.

A pattern should be written according to shell syntax, using wildcard characters to effect globbing. Most characters in the pattern stand for themselves in the matched string, and case is significant: ‘a’ will match only ‘a’, and not ‘A’. The character ‘?’ in the pattern matches any single character in the matched string. The character ‘*’ in the pattern matches zero, one, or more single characters in the matched string. The character ‘\’ says to take the following character of the pattern literally; it is useful when one needs to match the ‘?’, ‘*’, ‘[’ or ‘\’ characters, themselves.

The character ‘[’, up to the matching ‘]’, introduces a character class. A character class is a list of acceptable characters for the next single character of the matched string. For example, ‘[abcde]’ would match any of the first five letters of the alphabet. Note that within a character class, all of the “special characters” listed above other than ‘\’ lose their special meaning; for example, ‘[-\\[*?]]’ would match any of the characters, ‘-’, ‘\’, ‘[’, ‘*’, ‘?’, or ‘]’. (Due to parsing constraints, the characters ‘-’ and ‘]’ must either come first or last in a character class.)

If the first character of the class after the opening ‘[’ is ‘!’ or ‘^’, then the meaning of the class is reversed. Rather than listing character to match, it lists those characters which are forbidden as the next single character of the matched string.

Other characters of the class stand for themselves. The special construction ‘[a-e]’, using an hyphen between two letters, is meant to represent all characters between a and e, inclusive.

Periods (‘.’) or forward slashes (‘/’) are not considered special for wildcard matches. However, if a pattern completely matches a directory prefix of a matched string, then it matches the full matched string: thus, excluding a directory also excludes all the files beneath it.

Controlling Pattern-Matching

For the purposes of this section, we call exclusion members all member names obtained while processing ‘--exclude’ and ‘--exclude-from’ options, and inclusion members those member names that were given in the command line or read from the file specified with ‘--files-from’ option.

These two pairs of member lists are used in the following operations:

‘--diff’, ‘--extract’, ‘--list’, ‘--update’.

There are no inclusion members in create mode (‘--create’ and ‘--append’), since in this mode the names obtained from the command line refer to files, not archive members.

By default, inclusion members are compared with archive members literally^{2} and exclusion members are treated as globbing patterns. For example:

2 Notice that earlier GNU tar versions used globbing for inclusion members, which

contradicted to UNIX98 specification and was not documented. See Appendix A

[Changes], page 159, for more information on this and other changes.

     $ tar tf foo.tar
     a.c
     b.c
     a.txt
     [remarks]
     # Member names are used verbatim:
     $ tar -xf foo.tar -v '[remarks]'
     [remarks]
     # Exclude member names are globbed:
     $ tar -xf foo.tar -v --exclude '*.c'
     a.txt
     [remarks]

This behavior can be altered by using the following options:

‘--wildcards’

Treat all member names as wildcards.

‘--no-wildcards’

Treat all member names as literal strings.

Thus, to extract files whose names end in ‘.c’, you can use:

     $ tar -xf foo.tar -v --wildcards '*.c'
     a.c
     b.c

Notice quoting of the pattern to prevent the shell from interpreting it.

The effect of ‘--wildcards’ option is canceled by ‘--no-wildcards’. This can be used to pass part of the command line arguments verbatim and other part as globbing patterns. For example, the following invocation:

     $ tar -xf foo.tar --wildcards '*.txt' --no-wildcards '[remarks]'

instructs tar to extract from ‘foo.tar’ all files whose names end in ‘.txt’ and the file named ‘[remarks]’.

Normally, a pattern matches a name if an initial subsequence of the name’s components matches the pattern, where ‘*’, ‘?’, and ‘[...]’ are the usual shell wildcards, ‘\’ escapes wildcards, and wildcards can match ‘/’.

Other than optionally stripping leading ‘/’ from names (see Section 6.10.2 [absolute], page 106), patterns and names are used as-is. For example, trailing ‘/’ is not trimmed from a user-specified name before deciding whether to exclude it.

However, this matching procedure can be altered by the options listed below. These options accumulate. For example:

     --ignore-case --exclude='makefile' --no-ignore-case ---exclude='readme'

ignores case when excluding ‘makefile’, but not when excluding ‘readme’.

‘--anchored’
‘--no-anchored’

If anchored, a pattern must match an initial subsequence of the name’s components. Otherwise, the pattern can match any subsequence. Default is ‘--no-anchored’ for exclusion members and ‘--anchored’ inclusion members.

‘--ignore-case’
‘--no-ignore-case’

When ignoring case, upper-case patterns match lower-case names and vice versa. When not ignoring case (the default), matching is case-sensitive.

‘--wildcards-match-slash’
‘--no-wildcards-match-slash’

When wildcards match slash (the default for exclusion members), a wildcard like ‘*’ in the pattern can match a ‘/’ in the name. Otherwise, ‘/’ is matched only by ‘/’.

The ‘--recursion’ and ‘--no-recursion’ options (see Section 6.9 [recurse], page 104) also affect how member patterns are interpreted. If recursion is in effect, a pattern matches a name if it matches any of the name’s parent directories.

The following table summarizes pattern-matching default values:

Members  Default settings

Inclusion

                     ‘--no-wildcards --anchored
                     --no-wildcards-match-slash’

Exclusion

                     ‘--wildcards --no-anchored
                     --wildcards-match-slash’

6.6 Quoting Member Names

When displaying member names, tar takes care to avoid ambiguities caused by certain characters. This is called name quoting. The characters in question are:

• Non-printable control characters:

Character  ASCII Character name
\a  7  Audible bell
\b  8  Backspace
\f  12  Form feed
\n  10  New line
\r  13  Carriage return
\t  9  Horizontal tabulation
\v  11  Vertical tabulation

• Space (ASCII 32)

• Single and double quotes (‘'’ and ‘"’)

• Backslash (‘\’)

The exact way tar uses to quote these characters depends on the quoting style. The default quoting style, called escape (see below), uses backslash notation to represent control characters, space and backslash. Using this quoting style, control characters are represented as listed in column ‘Character’ in the above table, a space is printed as ‘\ ’ and a backslash as ‘\\’.

GNU tar offers seven distinct quoting styles, which can be selected using

‘--quoting-style’ option:
‘--quoting-style=style’

Sets quoting style. Valid values for style argument are: literal, shell, shell-always, c, escape, locale, clocale.

These styles are described in detail below. To illustrate their effect, we will use an imaginary tar archive ‘arch.tar’ containing the following members:

     # 1. Contains horizontal tabulation character.
     a       tab
     # 2. Contains newline character
     a
     newline
     # 3. Contains a space
     a space
     # 4. Contains double quotes
     a"double"quote
     # 5. Contains single quotes
     a'single'quote
     # 6. Contains a backslash character:
     a\backslash

Here is how usual ls command would have listed them, if they had existed in the current working directory:

     $ ls
     a\ttab
     a\nnewline
     a\ space
     a"double"quote
     a'single'quote
     a\\backslash

Quoting styles:

‘literal’ No quoting, display each character as is:

                 $ tar tf arch.tar --quoting-style=literal
                 ./
                 ./a space
                 ./a'single'quote
                 ./a"double"quote
                 ./a\backslash
                 ./a tab
                 ./a
                 newline

‘shell’ Display characters the same way Bourne shell does: control characters, except ‘\t’ and ‘\n’, are printed using backslash escapes, ‘\t’ and ‘\n’ are printed as is, and a single quote is printed as ‘\'’. If a name contains any quoted characters, it is enclosed in single quotes. In particular, if a name contains single quotes, it is printed as several single-quoted strings:

     $ tar tf arch.tar --quoting-style=shell
     ./
     './a space'
     './a'\''single'\''quote'
     './a"double"quote'
     './a\backslash'
     './a tab'
     './a
     newline'

‘shell-always’

Same as ‘shell’, but the names are always enclosed in single quotes:

     $ tar tf arch.tar --quoting-style=shell-always
     './'
     './a space'
     './a'\''single'\''quote'
     './a"double"quote'
     './a\backslash'
     './a tab'
     './a
     newline'

‘c’ Use the notation of the C programming language. All names are enclosed in double quotes. Control characters are quoted using backslash notations, double quotes are represented as ‘\"’, backslash characters are represented as ‘\\’. Single quotes and spaces are not quoted:

     $ tar tf arch.tar --quoting-style=c
     "./"
     "./a space"
     "./a'single'quote"
     "./a\"double\"quote"
     "./a\\backslash"
     "./a\ttab"
     "./a\nnewline"

‘escape’ Control characters are printed using backslash notation, a space is printed as ‘\ ’ and a backslash as ‘\\’. This is the default quoting style, unless it was changed when configured the package.

     $ tar tf arch.tar --quoting-style=escape
     ./
     ./a space
     ./a'single'quote
     ./a"double"quote
     ./a\\backslash
     ./a\ttab
     ./a\nnewline

‘locale’ Control characters, single quote and backslash are printed using backslash notation. All names are quoted using left and right quotation marks, appropriate to the current locale. If it does not define quotation marks, use ‘`’ as left and ‘'’ as right quotation marks. Any occurrences of the right quotation mark in a name are escaped with ‘\’, for example:

For example:

     $ tar tf arch.tar --quoting-style=locale
     `./'
     `./a space'
     `./a\'single\'quote'
     `./a"double"quote'
     `./a\\backslash'
     `./a\ttab'
     `./a\nnewline'

‘clocale’ Same as ‘locale’, but ‘"’ is used for both left and right quotation marks, if not provided by the currently selected locale:

     $ tar tf arch.tar --quoting-style=clocale
     "./"
     "./a space"
     "./a'single'quote"
     "./a\"double\"quote"
     "./a\\backslash"
     "./a\ttab"
     "./a\nnewline"

You can specify which characters should be quoted in addition to those implied by the current quoting style:

‘--quote-chars=string’

Always quote characters from string, even if the selected quoting style would not quote them.

For example, using ‘escape’ quoting (compare with the usual escape listing above):

     $ tar tf arch.tar --quoting-style=escape --quote-chars=' "'
     ./
     ./a\ space
     ./a'single'quote
     ./a\"double\"quote
     ./a\\backslash
     ./a\ttab
     ./a\nnewline

To disable quoting of such additional characters, use the following option:

‘--no-quote-chars=string’

Remove characters listed in string from the list of quoted characters set by the previous ‘--quote-chars’ option.

This option is particularly useful if you have added ‘--quote-chars’ to your TAR_OPTIONS (see [TAR OPTIONS], page 23) and wish to disable it for the current invocation.

Note, that ‘--no-quote-chars’ does not disable those characters that are quoted by default in the selected quoting style.

6.7 Modifying File and Member Names

Tar archives contain detailed information about files stored in them and full file names are part of that information. When storing file to an archive, its file name is recorded in the archive along with the actual file contents. When restoring from an archive, a file is created on disk with exactly the same name as that stored in the archive. In the majority of cases this is the desired behavior of a file archiver. However, there are some cases when it is not.

First of all, it is often unsafe to extract archive members with absolute file names or those that begin with a ‘../’. GNU tar takes special precautions when extracting such names and provides a special option for handling them, which is described in Section 6.10.2 [absolute], page 106.

Secondly, you may wish to extract file names without some leading directory components, or with otherwise modified names. In other cases it is desirable to store files under differing names in the archive.

GNU tar provides two options for these needs.

‘--strip-components=number’

Strip given number of leading components from file names before extraction.

For example, suppose you have archived whole ‘/usr’ hierarchy to a tar archive named ‘usr.tar’. Among other files, this archive contains ‘usr/include/stdlib.h’, which you wish to extract to the current working directory. To do so, you type:

     $ tar -xf usr.tar --strip=2 usr/include/stdlib.h

The option ‘--strip=2’ instructs tar to strip the two leading components (‘usr/’ and ‘include/’) off the file name.

If you add to the above invocation ‘--verbose’ (‘-v’) option, you will note that the verbose listing still contains the full file name, with the two removed components still in place. This can be inconvenient, so tar provides a special option for altering this behavior:

‘--show-transformed-names’

Display file or member names with all requested transformations applied.

For example:

     $ tar -xf usr.tar -v --strip=2 usr/include/stdlib.h
     usr/include/stdlib.h
     $ tar -xf usr.tar -v --strip=2 --show-transformed usr/include/stdlib.h
     stdlib.h

Notice that in both cases the file is ‘stdlib.h’ extracted to the current working directory, ‘--show-transformed-names’ affects only the way its name is displayed.

This option is especially useful for verifying whether the invocation will have the desired effect. Thus, before running

     $ tar -x --strip=n

it is often advisable to run

     $ tar -t -v --show-transformed --strip=n

to make sure the command will produce the intended results.

In case you need to apply more complex modifications to the file name, GNU tar provides a general-purpose transformation option:

‘--transform=expression’

Modify file names using supplied expression.

The expression is a sed-like replace expression of the form:

     s/regexp/replace/[flags]

where regexp is a regular expression, replace is a replacement for each file name part that matches regexp. Both regexp and replace are described in detail in section “The ‘s’ Command” in GNU sed.

Supported flags are: ‘g’ Apply the replacement to all matches to the regexp, not just

the first.

‘i’ Use case-insensitive matching ‘x’ regexp is an extended regular expression (see section “Extended

regular expressions” in GNU sed).

‘number’ Only replace the numberth match of the regexp.

Note: the posix standard does not specify what should happen when you mix the ‘g’ and number modifiers. GNU tar follows the GNU sed implementation in this regard, so the interaction is defined to be: ignore matches before the numberth, and then match and replace all matches from the numberth on.

Any delimiter can be used in lieue of ‘/’, the only requirement being that it be used consistently throughout the expression. For example, the following two expressions are equivalent:

     s/one/two/ s,one,two,

Changing delimiters is often useful when the regex contains slashes. For example, it is more convenient to write s,/,-, than s/\//-/.

Here are several examples of ‘--transform’ usage:

1. Extract ‘usr/’ hierarchy into ‘usr/local/’:

         $ tar --transform='s,usr/,usr/local/,' -x -f arch.tar

2. Strip two leading directory components (equivalent to

    ‘--strip-components=2’):
         $ tar --transform='s,/*[^/]*/[^/]*/,,' -x -f arch.tar

3. Prepend ‘/prefix/’ to each file name:

         $ tar --transform 's,^,/prefix/,' -x -f arch.tar

4. Convert each file name to lower case:

         $ tar --transform 's/.*/\L&/' -x -f arch.tar
   Unlike ‘--strip-components’, ‘--transform’ can be used in any GNU

tar operation mode. For example, the following command adds files to the archive while replacing the leading ‘usr/’ component with ‘var/’:

     $ tar -cf arch.tar --transform='s,^usr/,var/,' /
   To test ‘--transform’ effect we suggest using ‘--show-transformed-names’

option:

     $ tar -cf arch.tar --transform='s,^usr/,var/,' \
            --verbose --show-transformed-names /

If both ‘--strip-components’ and ‘--transform’ are used together, then ‘--transform’ is applied first, and the required number of components is then stripped from its result.

6.8 Operating Only on New Files

(This message will disappear, once this node revised.)

   The ‘--after-date=date’ (‘--newer=date’, ‘-N date’) option causes

tar to only work on files whose data modification or status change times are newer than the date given. If date starts with ‘/’ or ‘.’, it is taken to be a file name; the data modification time of that file is used as the date. If you use this option when creating or appending to an archive, the archive will only include new files. If you use ‘--after-date’ when extracting an archive, tar will only extract files newer than the date you specify.

If you only want tar to make the date comparison based on modification of the file’s data (rather than status changes), then use the

‘--newer-mtime=date’ option.

You may use these options with any operation. Note that these options differ from the ‘--update’ (‘-u’) operation in that they allow you to specify a particular date against which tar can compare when deciding whether or not to archive the files.

‘--after-date=date’
‘--newer=date’

‘-N date’ Only store files newer than date.

Acts on files only if their data modification or status change times are later than date. Use in conjunction with any operation. If date starts with ‘/’ or ‘.’, it is taken to be a file name; the data modification time of that file is used as the date.

‘--newer-mtime=date’

Acts like ‘--after-date’, but only looks at data modification times.

These options limit tar to operate only on files which have been modified after the date specified. A file’s status is considered to have changed if its contents have been modified, or if its owner, permissions, and so forth, have been changed. (For more information on how to specify a date, see Chapter 7 [Date input formats], page 109; remember that the entire date argument must be quoted if it contains any spaces.)

Gurus would say that ‘--after-date’ tests both the data modification time (mtime, the time the contents of the file were last modified) and the status change time (ctime, the time the file’s status was last changed: owner, permissions, etc.) fields, while ‘--newer-mtime’ tests only the mtime field.

To be precise, ‘--after-date’ checks both mtime and ctime and processes the file if either one is more recent than date, while ‘--newer-mtime’ only checks mtime and disregards ctime. Neither does it use atime (the last time the contents of the file were looked at).

Date specifiers can have embedded spaces. Because of this, you may need to quote date arguments to keep the shell from parsing them as separate arguments. For example, the following command will add to the archive all the files modified less than two days ago:

     $ tar -cf foo.tar --newer-mtime '2 days ago'

When any of these options is used with the option ‘--verbose’ (see [verbose tutorial], page 8) GNU tar will try to convert the specified date back to its textual representation and compare that with the one given with the option. If the two dates differ, tar will print a warning saying what date it will use. This is to help user ensure he is using the right date. For example:

     $ tar -c -f archive.tar --after-date='10 days ago' .
     tar: Option --after-date: Treating date `10 days ago' as 2006-06-11
     13:19:37.232434

Please Note: ‘--after-date’ and ‘--newer-mtime’ should not be used for incremental backups. See Section 5.2 [Incremental Dumps], page 74, for proper way of creating incremental backups.

6.9 Descending into Directories

(This message will disappear, once this node revised.)

Usually, tar will recursively explore all directories (either those given on the command line or through the ‘--files-from’ option) for the various files they contain. However, you may not always want tar to act this way.

The ‘--no-recursion’ option inhibits tar’s recursive descent into specified directories. If you specify ‘--no-recursion’, you can use the find utility for hunting through levels of directories to construct a list of file names which you could then pass to tar. find allows you to be more selective when choosing which files to archive; see Section 6.3 [files], page 89, for more information on using find with tar, or look.

‘--no-recursion’

Prevents tar from recursively descending directories.

‘--recursion’

Requires tar to recursively descend directories. This is the default.

When you use ‘--no-recursion’, GNU tar grabs directory entries themselves, but does not descend on them recursively. Many people use find for locating files they want to back up, and since tar usually recursively descends on directories, they have to use the ‘-not -type d’ test in their find invocation (see section “Type test” in Finding Files), as they usually do not want all the files in a directory. They then use the ‘--files-from’ option to archive the files located via find.

The problem when restoring files archived in this manner is that the directories themselves are not in the archive; so the ‘--same-permissions’ (‘--preserve-permissions’, ‘-p’) option does not affect them—while users might really like it to. Specifying ‘--no-recursion’ is a way to tell tar to grab only the directory entries given to it, adding no new files on its own. To summarize, if you use find to create a list of files to be stored in an archive, use it as follows:

     $ find dir tests | \
       tar -cf archive -T - --no-recursion

The ‘--no-recursion’ option also applies when extracting: it causes tar to extract only the matched directory entries, not the files under those directories.

The ‘--no-recursion’ option also affects how globbing patterns are interpreted (see [controlling pattern-matching], page 94).

The ‘--no-recursion’ and ‘--recursion’ options apply to later options and operands, and can be overridden by later occurrences of

‘--no-recursion’ and ‘--recursion’. For example:
     $ tar -cf jams.tar --no-recursion grape --recursion grape/concord

creates an archive with one entry for ‘grape’, and the recursive contents of ‘grape/concord’, but no entries under ‘grape’ other than ‘grape/concord’.

6.10 Crossing File System Boundaries

(This message will disappear, once this node revised.)

tar will normally automatically cross file system boundaries in order to archive files which are part of a directory tree. You can change this behavior by running tar and specifying ‘--one-file-system’. This option only affects files that are archived because they are in a directory that is being archived; tar will still archive files explicitly named on the command line or through ‘--files-from’, regardless of where they reside.

‘--one-file-system’

Prevents tar from crossing file system boundaries when archiving. Use in conjunction with any write operation.

The ‘--one-file-system’ option causes tar to modify its normal behavior in archiving the contents of directories. If a file in a directory is not on the same file system as the directory itself, then tar will not archive that file. If the file is a directory itself, tar will not archive anything beneath it; in other words, tar will not cross mount points.

This option is useful for making full or incremental archival backups of a file system. If this option is used in conjunction with ‘--verbose’ (‘-v’), files that are excluded are mentioned by name on the standard error.

6.10.1 Changing the Working Directory

To change the working directory in the middle of a list of file names, either on the command line or in a file specified using ‘--files-from’ (‘-T’), use ‘--directory’ (‘-C’). This will change the working directory to the specified directory after that point in the list.

‘--directory=directory’
‘-C directory’

Changes the working directory in the middle of a command line.

For example,

     $ tar -c -f jams.tar grape prune -C food cherry

will place the files ‘grape’ and ‘prune’ from the current directory into the archive ‘jams.tar’, followed by the file ‘cherry’ from the directory ‘food’. This option is especially useful when you have several widely separated files that you want to store in the same archive.

Note that the file ‘cherry’ is recorded in the archive under the precise name ‘cherry’, not ‘food/cherry’. Thus, the archive will contain three files that all appear to have come from the same directory; if the archive is extracted with plain ‘tar --extract’, all three files will be written in the current directory.

Contrast this with the command,

     $ tar -c -f jams.tar grape prune -C food red/cherry

which records the third file in the archive under the name ‘red/cherry’ so that, if the archive is extracted using ‘tar --extract’, the third file will be written in a subdirectory named ‘orange-colored’.

You can use the ‘--directory’ option to make the archive independent of the original name of the directory holding the files. The following command

places the files ‘/etc/passwd’, ‘/etc/hosts’, and ‘/lib/libc.a’ into the

archive ‘foo.tar’:

     $ tar -c -f foo.tar -C /etc passwd hosts -C /lib libc.a

However, the names of the archive members will be exactly what they were on the command line: ‘passwd’, ‘hosts’, and ‘libc.a’. They will not appear to be related by file name to the original directories where those files were located.

Note that ‘--directory’ options are interpreted consecutively. If ‘--directory’ specifies a relative file name, it is interpreted relative to the then current directory, which might not be the same as the original current working directory of tar, due to a previous ‘--directory’ option.

When using ‘--files-from’ (see Section 6.3 [files], page 89), you can put various tar options (including ‘-C’) in the file list. Notice, however, that in this case the option and its argument may not be separated by whitespace. If you use short option, its argument must either follow the option letter immediately, without any intervening whitespace, or occupy the next line. Otherwise, if you use long option, separate its argument by an equal sign.

For instance, the file list for the above example will be:

     -C/etc
     passwd
     hosts
     --directory=/lib
     libc.a

To use it, you would invoke tar as follows:

     $ tar -c -f foo.tar --files-from list

The interpretation of ‘--directory’ is disabled by ‘--null’ option.

6.10.2 Absolute File Names

(This message will disappear, once this node revised.)

‘--absolute-names’

‘-P’ Do not strip leading slashes from file names, and permit file names containing a ‘..’ file name component.

By default, GNU tar drops a leading ‘/’ on input or output, and complains about file names containing a ‘..’ component. This option turns off this behavior.

When tar extracts archive members from an archive, it strips any leading slashes (‘/’) from the member name. This causes absolute member names in the archive to be treated as relative file names. This allows you to have such members extracted wherever you want, instead of being restricted to extracting the member in the exact directory named in the archive. For example, if the archive member has the name ‘/etc/passwd’, tar will extract it as if the name were really ‘etc/passwd’.

File names containing ‘..’ can cause problems when extracting, so tar normally warns you about such files when creating an archive, and rejects attempts to extracts such files.

Other tar programs do not do this. As a result, if you create an archive whose member names start with a slash, they will be difficult for other people with a non-GNU tar program to use. Therefore, GNU tar also strips leading slashes from member names when putting members into the archive. For example, if you ask tar to add the file ‘/bin/ls’ to an archive, it will do so, but the member name will be ‘bin/ls’.^{3}

If you use the ‘--absolute-names’ (‘-P’) option, tar will do none of these transformations.

To archive or extract files relative to the root directory, specify the

‘--absolute-names’ (‘-P’) option.

Normally, tar acts on files relative to the working directory—ignoring superior directory names when archiving, and ignoring leading slashes when extracting.

When you specify ‘--absolute-names’ (‘-P’), tar stores file names including all superior directory names, and preserves leading slashes. If you only invoked tar from the root directory you would never need the ‘--absolute-names’ option, but using this option may be more convenient than switching to root.

‘--absolute-names’

Preserves full file names (including superior directory names) when archiving files. Preserves leading slash when extracting files.

tar prints out a message about removing the ‘/’ from file names. This message appears once per GNU tar invocation. It represents something which ought to be told; ignoring what it means can cause very serious surprises, later.

Some people, nevertheless, do not want to see this message. Wanting to play really dangerously, one may of course redirect tar standard error to the sink. For example, under sh:

     $ tar -c -f archive.tar /home 2> /dev/null

Another solution, both nicer and simpler, would be to change to the ‘/’ directory first, and then avoid absolute notation. For example:

3 A side effect of this is that when ‘--create’ is used with ‘--verbose’ the resulting

output is not, generally speaking, the same as the one you’d get running tar --list

command. This may be important if you use some scripts for comparing both outputs.

See [listing member and file names], page 15, for the information on how to handle this

case.

108 GNU tar 1.16.1

$ (cd / && tar -c -f archive.tar home)
# or:
$ tar -c -f archive.tar -C  / home

7 Date input formats

First, a quote:

Our units of temporal measurement, from seconds on up to months, are so complicated, asymmetrical and disjunctive so as to make coherent mental reckoning in time all but impossible. Indeed, had some tyrannical god contrived to enslave our minds to time, to make it all but impossible for us to escape subjection to sodden routines and unpleasant surprises, he could hardly have done better than handing down our present system. It is like a set of trapezoidal building blocks, with no vertical or horizontal surfaces, like a language in which the simplest thought demands ornate constructions, useless particles and lengthy circumlocutions. Unlike the more successful patterns of language and science, which enable us to face experience boldly or at least level-headedly, our system of temporal calculation silently and persistently encourages our terror of time. … It is as though architects had to measure length in feet, width in meters and height in ells; as though basic instruction manuals demanded a knowledge of five different languages. It is no wonder then that we often look into our own immediate past or future, last Tuesday or a week from Sunday, with feelings of helpless confusion. …

— Robert Grudin, Time and the Art of Living. This section describes the textual date representations that gnu programs

accept. These are the strings you, as a user, can supply as arguments to the various programs. The C interface (via the get_date function) is not described here.

7.1 General date syntax

A date is a string, possibly empty, containing many items separated by whitespace. The whitespace may be omitted when no ambiguity arises. The empty string means the beginning of today (i.e., midnight). Order of the items is immaterial. A date string may contain many flavors of items:

• calendar date items

• time of day items

• time zone items

• day of the week items

• relative items

• pure numbers. We describe each of these item types in turn, below.

A few ordinal numbers may be written out in words in some contexts. This is most useful for specifying day of the week items or relative items (see below). Among the most commonly used ordinal numbers, the word ‘last’ stands for −1, ‘this’ stands for 0, and ‘first’ and ‘next’ both stand for 1. Because the word ‘second’ stands for the unit of time there is no way to write the ordinal number 2, but for convenience ‘third’ stands for 3, ‘fourth’ for 4, ‘fifth’ for 5, ‘sixth’ for 6, ‘seventh’ for 7, ‘eighth’ for 8, ‘ninth’ for 9, ‘tenth’ for 10, ‘eleventh’ for 11 and ‘twelfth’ for 12.

When a month is written this way, it is still considered to be written numerically, instead of being “spelled in full”; this changes the allowed strings.

In the current implementation, only English is supported for words and abbreviations like ‘AM’, ‘DST’, ‘EST’, ‘first’, ‘January’, ‘Sunday’, ‘tomorrow’, and ‘year’.

The output of the date command is not always acceptable as a date string, not only because of the language problem, but also because there is no standard meaning for time zone items like ‘IST’. When using date to generate a date string intended to be parsed later, specify a date format that is independent of language and that does not use time zone items other than ‘UTC’ and ‘Z’. Here are some ways to do this:

    $ LC_ALL=C TZ=UTC0 date
    Mon Mar  1 00:21:42 UTC 2004
    $ TZ=UTC0 date +'%Y-%m-%d %H:%M:%SZ'
    2004-03-01 00:21:42Z
    $ date --iso-8601=ns | tr T ' '  # --iso-8601 is a GNU extension.
    2004-02-29 16:21:42,692722128-0800
    $ date --rfc-2822  # a GNU extension
    Sun, 29 Feb 2004 16:21:42 -0800
    $ date +'%Y-%m-%d %H:%M:%S %z'  # %z is a GNU extension.
    2004-02-29 16:21:42 -0800
    $ date +'@%s.%N'  # %s and %N are GNU extensions.
    @1078100502.692722128

Alphabetic case is completely ignored in dates. Comments may be introduced between round parentheses, as long as included parentheses are properly nested. Hyphens not followed by a digit are currently ignored. Leading zeros on numbers are ignored.

Invalid dates like ‘2005-02-29’ or times like ‘24:00’ are rejected. In the typical case of a host that does not support leap seconds, a time like ‘23:59:60’ is rejected even if it corresponds to a valid leap second.

7.2 Calendar date items

A calendar date item specifies a day of the year. It is specified differently, depending on whether the month is specified numerically or literally. All these strings specify the same calendar date:

    1972-09-24     # iso 8601.
    72-9-24        # Assume 19xx for 69 through 99,
                   # 20xx for 00 through 68.

72-09-24       # Leading zeros are ignored.
9/24/72        # Common U.S. writing.
24 September 1972
24 Sept 72     # September has a special abbreviation.
24 Sep 72      # Three-letter abbreviations always allowed.
Sep 24, 1972
24-sep-72
24sep72

The year can also be omitted. In this case, the last specified year is used, or the current year if none. For example:

9/24
sep 24

Here are the rules. For numeric months, the iso 8601 format ‘year-month-day’ is allowed,

where year is any positive number, month is a number between 01 and 12, and day is a number between 01 and 31. A leading zero must be present if a number is less than ten. If year is 68 or smaller, then 2000 is added to it; otherwise, if year is less than 100, then 1900 is added to it. The construct ‘month/day/year’, popular in the United States, is accepted. Also ‘month/day’, omitting the year.

Literal months may be spelled out in full: ‘January’, ‘February’, ‘March’,

‘April’, ‘May’, ‘June’, ‘July’, ‘August’, ‘September’, ‘October’, ‘November’

or ‘December’. Literal months may be abbreviated to their first three letters, possibly followed by an abbreviating dot. It is also permitted to write ‘Sept’ instead of ‘September’.

When months are written literally, the calendar date may be given as any of the following:

day month year
day month
month day year
day-month-year

Or, omitting the year:

month day

7.3 Time of day items

A time of day item in date strings specifies the time on a given day. Here are some examples, all of which represent the same time:

20:02:00.000000
20:02
8:02pm
20:02-0500      # In est (U.S. Eastern Standard Time).

More generally, the time of day may be given as ‘hour:minute:second’, where hour is a number between 0 and 23, minute is a number between 0 and 59, and second is a number between 0 and 59 possibly followed by ‘.’ or ‘,’ and a fraction containing one or more digits. Alternatively, ‘:second’ can be omitted, in which case it is taken to be zero. On the rare hosts that support leap seconds, second may be 60.

If the time is followed by ‘am’ or ‘pm’ (or ‘a.m.’ or ‘p.m.’), hour is restricted to run from 1 to 12, and ‘:minute’ may be omitted (taken to be zero). ‘am’ indicates the first half of the day, ‘pm’ indicates the second half of the day. In this notation, 12 is the predecessor of 1: midnight is ‘12am’ while noon is ‘12pm’. (This is the zero-oriented interpretation of ‘12am’ and ‘12pm’, as opposed to the old tradition derived from Latin which uses ‘12m’ for noon and ‘12pm’ for midnight.)

The time may alternatively be followed by a time zone correction, expressed as ‘shhmm’, where s is ‘+’ or ‘-’, hh is a number of zone hours and mm is a number of zone minutes. You can also separate hh from mm with a colon. When a time zone correction is given this way, it forces interpretation of the time relative to Coordinated Universal Time (utc), overriding any previous specification for the time zone or the local time zone. For example, ‘+0530’ and ‘+05:30’ both stand for the time zone 5.5 hours ahead of utc (e.g., India). The minute part of the time of day may not be elided when a time zone correction is used. This is the best way to specify a time zone correction by fractional parts of an hour.

Either ‘am’/‘pm’ or a time zone correction may be specified, but not both.

7.4 Time zone items

A time zone item specifies an international time zone, indicated by a small set of letters, e.g., ‘UTC’ or ‘Z’ for Coordinated Universal Time. Any included periods are ignored. By following a non-daylight-saving time zone by the string ‘DST’ in a separate word (that is, separated by some white space), the corresponding daylight saving time zone may be specified. Alternatively, a non-daylight-saving time zone can be followed by a time zone correction, to add the two values. This is normally done only for ‘UTC’; for example, ‘UTC+05:30’ is equivalent to ‘+05:30’.

Time zone items other than ‘UTC’ and ‘Z’ are obsolescent and are not recommended, because they are ambiguous; for example, ‘EST’ has a different meaning in Australia than in the United States. Instead, it’s better to use unambiguous numeric time zone corrections like ‘-0500’, as described in the previous section.

If neither a time zone item nor a time zone correction is supplied, time stamps are interpreted using the rules of the default time zone (see Section 7.9 [Specifying time zone rules], page 115).

7.5 Day of week items

The explicit mention of a day of the week will forward the date (only if necessary) to reach that day of the week in the future.

Days of the week may be spelled out in full: ‘Sunday’, ‘Monday’, ‘Tuesday’, ‘Wednesday’, ‘Thursday’, ‘Friday’ or ‘Saturday’. Days may be abbreviated to their first three letters, optionally followed by a period. The special abbreviations ‘Tues’ for ‘Tuesday’, ‘Wednes’ for ‘Wednesday’ and ‘Thur’ or ‘Thurs’ for ‘Thursday’ are also allowed.

A number may precede a day of the week item to move forward supplementary weeks. It is best used in expression like ‘third monday’. In this context, ‘last day’ or ‘next day’ is also acceptable; they move one week before or after the day that day by itself would represent.

A comma following a day of the week item is ignored.

7.6 Relative items in date strings

Relative items adjust a date (or the current date if none) forward or backward. The effects of relative items accumulate. Here are some examples:

    1 year
    1 year ago
    3 years
    2 days

The unit of time displacement may be selected by the string ‘year’ or ‘month’ for moving by whole years or months. These are fuzzy units, as years and months are not all of equal duration. More precise units are ‘fortnight’ which is worth 14 days, ‘week’ worth 7 days, ‘day’ worth 24 hours, ‘hour’ worth 60 minutes, ‘minute’ or ‘min’ worth 60 seconds, and ‘second’ or ‘sec’ worth one second. An ‘s’ suffix on these units is accepted and ignored.

The unit of time may be preceded by a multiplier, given as an optionally signed number. Unsigned numbers are taken as positively signed. No number at all implies 1 for a multiplier. Following a relative item by the string ‘ago’ is equivalent to preceding the unit by a multiplier with value −1.

The string ‘tomorrow’ is worth one day in the future (equivalent to ‘day’), the string ‘yesterday’ is worth one day in the past (equivalent to ‘day ago’).

The strings ‘now’ or ‘today’ are relative items corresponding to zerovalued time displacement, these strings come from the fact a zero-valued time displacement represents the current time when not otherwise changed by previous items. They may be used to stress other items, like in ‘12:00 today’. The string ‘this’ also has the meaning of a zero-valued time displacement, but is preferred in date strings like ‘this thursday’.

When a relative item causes the resulting date to cross a boundary where the clocks were adjusted, typically for daylight saving time, the resulting date and time are adjusted accordingly.

The fuzz in units can cause problems with relative items. For example, ‘2003-07-31 -1 month’ might evaluate to 2003-07-01, because 2003-06-31 is an invalid date. To determine the previous month more reliably, you can ask for the month before the 15th of the current month. For example:

    $ date -R
    Thu, 31 Jul 2003 13:02:39 -0700
    $ date --date='-1 month' +'Last month was %B?'
    Last month was July?
    $ date --date="$(date +%Y-%m-15) -1 month" +'Last month was %B!'
    Last month was June!

Also, take care when manipulating dates around clock changes such as daylight saving leaps. In a few cases these have added or subtracted as much as 24 hours from the clock, so it is often wise to adopt universal time by setting the TZ environment variable to ‘UTC0’ before embarking on calendrical calculations.

7.7 Pure numbers in date strings

The precise interpretation of a pure decimal number depends on the context in the date string.

If the decimal number is of the form yyyymmdd and no other calendar date item (see Section 7.2 [Calendar date items], page 110) appears before it in the date string, then yyyy is read as the year, mm as the month number and dd as the day of the month, for the specified calendar date.

If the decimal number is of the form hhmm and no other time of day item appears before it in the date string, then hh is read as the hour of the day and mm as the minute of the hour, for the specified time of day. mm can also be omitted.

If both a calendar date and a time of day appear to the left of a number in the date string, but no relative item, then the number overrides the year.

7.8 Seconds since the Epoch

If you precede a number with ‘@’, it represents an internal time stamp as a count of seconds. The number can contain an internal decimal point (either ‘.’ or ‘,’); any excess precision not supported by the internal representation is truncated toward minus infinity. Such a number cannot be combined with any other date item, as it specifies a complete time stamp.

Internally, computer times are represented as a count of seconds since an epoch—a well-defined point of time. On GNU and POSIX systems, the epoch is 1970-01-01 00:00:00 utc, so ‘@0’ represents this time, ‘@1’ represents 1970-01-01 00:00:01 utc, and so forth. GNU and most other POSIX-compliant systems support such times as an extension to POSIX, using negative counts, so that ‘@-1’ represents 1969-12-31 23:59:59 utc.

Traditional Unix systems count seconds with 32-bit two’s-complement integers and can represent times from 1901-12-13 20:45:52 through 2038-01-19 03:14:07 utc. More modern systems use 64-bit counts of seconds with nanosecond subcounts, and can represent all the times in the known lifetime of the universe to a resolution of 1 nanosecond.

On most hosts, these counts ignore the presence of leap seconds. For example, on most hosts ‘@915148799’ represents 1998-12-31 23:59:59 utc, ‘@915148800’ represents 1999-01-01 00:00:00 utc, and there is no way to represent the intervening leap second 1998-12-31 23:59:60 utc.

7.9 Specifying time zone rules

Normally, dates are interpreted using the rules of the current time zone, which in turn are specified by the TZ environment variable, or by a system default if TZ is not set. To specify a different set of default time zone rules that apply just to one date, start the date with a string of the form ‘TZ="rule"’. The two quote characters (‘"’) must be present in the date, and any quotes or backslashes within rule must be escaped by a backslash.

For example, with the GNU date command you can answer the question “What time is it in New York when a Paris clock shows 6:30am on October 31, 2004?” by using a date beginning with ‘TZ="Europe/Paris"’ as shown in the following shell transcript:

    $ export TZ="America/New_York"
    $ date --date='TZ="Europe/Paris" 2004-10-31 06:30'
    Sun Oct 31 01:30:00 EDT 2004

In this example, the ‘--date’ operand begins with its own TZ setting, so the rest of that operand is processed according to ‘Europe/Paris’ rules, treating the string ‘2004-10-31 06:30’ as if it were in Paris. However, since the output of the date command is processed according to the overall time zone rules, it uses New York time. (Paris was normally six hours ahead of New York in 2004, but this example refers to a brief Halloween period when the gap was five hours.)

A TZ value is a rule that typically names a location in the ‘tz’

database (http://www.twinsun.com/tz/tz-link.htm).  A recent

catalog of location names appears in the TWiki Date and Time

Gateway (http://twiki.org/cgi-bin/xtra/tzdate). A few non-GNU

hosts require a colon before a location name in a TZ setting, e.g.,

‘TZ=":America/New_York"’.

The ‘tz’ database includes a wide variety of locations ranging from

‘Arctic/Longyearbyen’ to ‘Antarctica/South_Pole’, but if you are at sea

and have your own private time zone, or if you are using a non-GNU host that does not support the ‘tz’ database, you may need to use a POSIX rule instead. Simple POSIX rules like ‘UTC0’ specify a time zone without daylight saving time; other rules can specify simple daylight saving regimes. See section “Specifying the Time Zone with TZ” in The GNU C Library.

7.10 Authors of get_date

get_date was originally implemented by Steven M. Bellovin (smb@research.att.com) while at the University of North Carolina at Chapel Hill. The code was later tweaked by a couple of people on Usenet, then completely overhauled by Rich $alz (rsalz@bbn.com) and Jim Berets (jberets@bbn.com) in August, 1990. Various revisions for the gnu system were made by David MacKenzie, Jim Meyering, Paul Eggert and others.

This chapter was originally produced by François Pinard

(pinard@iro.umontreal.ca) from the ‘getdate.y’ source code,

and then edited by K. Berry (kb@cs.umb.edu).

8 Controlling the Archive Format

Due to historical reasons, there are several formats of tar archives. All of them are based on the same principles, but have some subtle differences that often make them incompatible with each other.

GNU tar is able to create and handle archives in a variety of formats. The most frequently used formats are (in alphabetical order):

gnu Format used by GNU tar versions up to 1.13.25. This format derived from an early POSIX standard, adding some improvements such as sparse file handling and incremental archives. Unfortunately these features were implemented in a way incompatible with other archive formats. Archives in ‘gnu’ format are able to hold pathnames of unlimited length.

oldgnu Format used by GNU tar of versions prior to 1.12.

v7 Archive format, compatible with the V7 implementation of tar. This format imposes a number of limitations. The most important of them are:

1. The maximum length of a file name is limited to 99 characters.

2. The maximum length of a symbolic link is limited to 99 characters.

3. It is impossible to store special files (block and character devices, fifos etc.)

4. Maximum value of user or group ID is limited to 2097151 (7777777 octal)

5. V7 archives do not contain symbolic ownership information (user and group name of the file owner).

This format has traditionally been used by Automake when producing Makefiles. This practice will change in the future, in the meantime, however this means that projects containing filenames more than 99 characters long will not be able to use GNU tar 1.16.1 and Automake prior to 1.9.

ustar Archive format defined by POSIX.1-1988 specification. It stores symbolic ownership information. It is also able to store special files. However, it imposes several restrictions as well:

1. The maximum length of a file name is limited to 256 characters, provided that the filename can be split at directory separator in two parts, first of them being at most 155 bytes long. So, in most cases the maximum file name length will be shorter than 256 characters.

2. The maximum length of a symbolic link name is limited to 100 characters.

3. Maximum size of a file the archive is able to accommodate is 8GB

4. Maximum value of UID/GID is 2097151.

5. Maximum number of bits in device major and minor numbers is 21.

star Format used by Jörg Schilling star implementation. GNU tar is able to read ‘star’ archives but currently does not produce them.

posix Archive format defined by POSIX.1-2001 specification. This is the most flexible and feature-rich format. It does not impose any restrictions on file sizes or filename lengths. This format is quite recent, so not all tar implementations are able to handle it properly. However, this format is designed in such a way that any tar implementation able to read ‘ustar’ archives will be able to read most ‘posix’ archives as well, with the only exception that any additional information (such as long file names etc.) will in such case be extracted as plain text files along with the files it refers to. This archive format will be the default format for future versions of GNU tar.

The following table summarizes the limitations of each of these formats:

Format UID  File Size  Path Name Devn
gnu  1.8e19  Unlimited  Unlimited  63
oldgnu 1.8e19  Unlimited  Unlimited  63
v7  2097151  8GB  99  n/a
ustar  2097151  8GB  256  21
posix  Unlimited  Unlimited  Unlimited  Unlimited

The default format for GNU tar is defined at compilation time. You may check it by running tar --help, and examining the last lines of its output. Usually, GNU tar is configured to create archives in ‘gnu’ format, however, future version will switch to ‘posix’.

8.1 Using Less Space through Compression

8.1.1 Creating and Reading Compressed Archives

GNU tar is able to create and read compressed archives. It supports gzip and bzip2 compression programs. For backward compatibility, it also supports compress command, although we strongly recommend against using it, since there is a patent covering the algorithm it uses and you could be sued for patent infringement merely by running compress! Besides, it is less effective than gzip and bzip2.

Creating a compressed archive is simple: you just specify a compression option along with the usual archive creation commands. The compression option is ‘-z’ (‘--gzip’) to create a gzip compressed archive, ‘-j’ (‘--bzip2’) to create a bzip2 compressed archive, and ‘-Z’ (‘--compress’) to use compress program. For example:

     $ tar cfz archive.tar.gz .

Reading compressed archive is even simpler: you don’t need to specify any additional options as GNU tar recognizes its format automatically. Thus, the following commands will list and extract the archive created in previous example:

     # List the compressed archive
     $ tar tf archive.tar.gz
     # Extract the compressed archive
     $ tar xf archive.tar.gz

The only case when you have to specify a decompression option while reading the archive is when reading from a pipe or from a tape drive that does not support random access. However, in this case GNU tar will indicate which option you should use. For example:

     $ cat archive.tar.gz | tar tf -
     tar: Archive is compressed.  Use -z option
     tar: Error is not recoverable: exiting now

If you see such diagnostics, just add the suggested option to the invocation of GNU tar:

     $ cat archive.tar.gz | tar tfz -

Notice also, that there are several restrictions on operations on compressed archives. First of all, compressed archives cannot be modified, i.e., you cannot update (‘--update’ (‘-u’)) them or delete (‘--delete’) members from them. Likewise, you cannot append another tar archive to a compressed archive using ‘--append’ (‘-r’)). Secondly, multi-volume archives cannot be compressed.

The following table summarizes compression options used by GNU tar.

‘-z’
‘--gzip’
‘--ungzip’

Filter the archive through gzip.

You can use ‘--gzip’ and ‘--gunzip’ on physical devices (tape drives, etc.) and remote files as well as on normal files; data to or from such devices or remote files is reblocked by another copy of the tar program to enforce the specified (or default) record size. The default compression parameters are used; if you need to override them, set GZIP environment variable, e.g.:

                 $ GZIP=--best tar cfz archive.tar.gz subdir

Another way would be to avoid the ‘--gzip’ (‘--gunzip’, ‘--ungzip’, ‘-z’) option and run gzip explicitly:

     $ tar cf - subdir | gzip --best -c - > archive.tar.gz

About corrupted compressed archives: gzip’ed files have no redundancy, for maximum compression. The adaptive nature of the compression scheme means that the compression tables are implicitly spread all over the archive. If you lose a few blocks, the dynamic construction of the compression tables becomes unsynchronized, and there is little chance that you could recover later in the archive.

There are pending suggestions for having a per-volume or perfile compression in GNU tar. This would allow for viewing the contents without decompression, and for resynchronizing decompression at every volume or file, in case of corrupted archives. Doing so, we might lose some compressibility. But this would have make recovering easier. So, there are pros and cons. We’ll see!

‘-j’

‘--bzip2’ Filter the archive through bzip2. Otherwise like ‘--gzip’.

‘-Z’
‘--compress’
‘--uncompress’

Filter the archive through compress. Otherwise like ‘--gzip’.

The GNU Project recommends you not use compress, because there is a patent covering the algorithm it uses. You could be sued for patent infringement merely by running compress.

‘--use-compress-program=prog’

Use external compression program prog. Use this option if you have a compression program that GNU tar does not support. There are two requirements to which prog should comply:

First, when called without options, it should read data from standard input, compress it and output it on standard output.

Secondly, if called with ‘-d’ argument, it should do exactly the opposite, i.e., read the compressed data from the standard input and produce uncompressed data on the standard output.

The ‘--use-compress-program’ option, in particular, lets you implement your own filters, not necessarily dealing with compression/decompression. For example, suppose you wish to implement PGP encryption on top of compression, using gpg (see section “gpg —- encryption and signing tool” in GNU Privacy Guard Manual). The following script does that:

     #! /bin/sh
     case $1 in
     -d) gpg --decrypt - | gzip -d -c;;
     '') gzip -c | gpg -s ;;
     *)  echo "Unknown option $1">&2; exit 1;;
     esac

Suppose you name it ‘gpgz’ and save it somewhere in your PATH. Then the following command will create a compressed archive signed with your private key:

     $ tar -cf foo.tar.gpgz --use-compress=gpgz .

Likewise, the following command will list its contents:

     $ tar -tf foo.tar.gpgz --use-compress=gpgz .

8.1.2 Archiving Sparse Files

Files in the file system occasionally have holes. A hole in a file is a section of the file’s contents which was never written. The contents of a hole reads as all zeros. On many operating systems, actual disk storage is not allocated for holes, but they are counted in the length of the file. If you archive such a file, tar could create an archive longer than the original. To have tar attempt to recognize the holes in a file, use ‘--sparse’ (‘-S’). When you use this option, then, for any file using less disk space than would be expected from its length, tar searches the file for consecutive stretches of zeros. It then records in the archive for the file where the consecutive stretches of zeros are, and only archives the “real contents” of the file. On extraction (using ‘--sparse’ is not needed on extraction) any such files have holes created wherever the continuous stretches of zeros were found. Thus, if you use ‘--sparse’, tar archives won’t take more space than the original.

‘-S’
‘--sparse’

This option instructs tar to test each file for sparseness before attempting to archive it. If the file is found to be sparse it is treated specially, thus allowing to decrease the amount of space used by its image in the archive. This option is meaningful only when creating or updating archives. It has no effect on extraction.

Consider using ‘--sparse’ when performing file system backups, to avoid archiving the expanded forms of files stored sparsely in the system.

Even if your system has no sparse files currently, some may be created in the future. If you use ‘--sparse’ while making file system backups as a matter of course, you can be assured the archive will never take more space on the media than the files take on disk (otherwise, archiving a disk filled with sparse files might take hundreds of tapes). See Section 5.2 [Incremental Dumps], page 74.

However, be aware that ‘--sparse’ option presents a serious drawback. Namely, in order to determine if the file is sparse tar has to read it before trying to archive it, so in total the file is read twice. So, always bear in mind that the time needed to process all files with this option is roughly twice the time needed to archive them without it.

When using ‘POSIX’ archive format, GNU tar is able to store sparse files using in three distinct ways, called sparse formats. A sparse format is identified by its number, consisting, as usual of two decimal numbers, delimited by a dot. By default, format ‘1.0’ is used. If, for some reason, you wish to use an earlier format, you can select it using ‘--sparse-version’ option.

‘--sparse-version=version’

Select the format to store sparse files in. Valid version values are: ‘0.0’, ‘0.1’ and ‘1.0’. See [Sparse Formats], page 174, for a detailed description of each format.

Using ‘--sparse-format’ option implies ‘--sparse’.

8.2 Handling File Attributes

(This message will disappear, once this node revised.) When tar reads files, it updates their access times. To avoid this, use

the ‘--atime-preserve[=METHOD]’ option, which can either reset the access time retroactively or avoid changing it in the first place.

Handling of file attributes

‘--atime-preserve’
‘--atime-preserve=replace’
‘--atime-preserve=system’

Preserve the access times of files that are read. This works only for files that you own, unless you have superuser privileges. ‘--atime-preserve=replace’ works on most systems, but it also restores the data modification time and updates the status change time. Hence it doesn’t interact with incremental dumps nicely (see Section 5.2 [Incremental Dumps], page 74), and it can set access or data modification times incorrectly if other programs access the file while tar is running. ‘--atime-preserve=system’ avoids changing the access time in the first place, if the operating system supports this. Unfortunately, this may or may not work on any given operating system or file system. If tar knows for sure it won’t work, it complains right away. Currently ‘--atime-preserve’ with no operand defaults

to ‘--atime-preserve=replace’, but this is intended to

change to ‘--atime-preserve=system’ when the latter is better-supported.

‘-m’

‘--touch’ Do not extract data modification time.

When this option is used, tar leaves the data modification times of the files it extracts as the times when the files were extracted, instead of setting it to the times recorded in the archive. This option is meaningless with ‘--list’ (‘-t’).

‘--same-owner’

Create extracted files with the same ownership they have in the archive. This is the default behavior for the superuser, so this option is meaningful only for non-root users, when tar is executed on those systems able to give files away. This is considered as a security flaw by many people, at least because it makes quite difficult to correctly account users for the disk space they occupy. Also, the suid or sgid attributes of files are easily and silently lost when files are given away. When writing an archive, tar writes the user id and user name separately. If it can’t find a user name (because the user id is not in ‘/etc/passwd’), then it does not write one. When restoring, it tries to look the name (if one was written) up in ‘/etc/passwd’. If it fails, then it uses the user id stored in the archive instead.

‘--no-same-owner’

‘-o’ Do not attempt to restore ownership when extracting. This is the default behavior for ordinary users, so this option has an effect only for the superuser.

‘--numeric-owner’

The ‘--numeric-owner’ option allows (ANSI) archives to be written without user/group name information or such information to be ignored when extracting. It effectively disables the generation and/or use of user/group name information. This option forces extraction using the numeric ids from the archive, ignoring the names. This is useful in certain circumstances, when restoring a backup from an emergency floppy with different passwd/group files for example. It is otherwise impossible to extract files with the right ownerships if the password file in use during the extraction does not match the one belonging to the file system(s) being extracted. This occurs, for example, if you are restoring your files after a major crash and had booted from an emergency floppy with no password file or put your disk into another machine to do the restore. The numeric ids are always saved into tar archives. The identifying names are added at create time when provided by the system, unless ‘--old-archive’ (‘-o’) is used. Numeric ids could be used when moving archives between a collection of machines using a centralized management for attribution of numeric ids to users and groups. This is often made through using the NIS capabilities. When making a tar file for distribution to other sites, it is sometimes cleaner to use a single owner for all files in the distribution, and nicer to specify the write permission bits of the files as stored in the archive independently of their actual value on the file system. The way to prepare a clean distribution is usually to have some Makefile rule creating a directory, copying all needed files in that directory, then setting ownership and permissions as wanted (there are a lot of possible schemes), and only then making a tar archive out of this directory, before cleaning everything out. Of course, we could add a lot of options to GNU tar for fine tuning permissions and ownership. This is not the good way, I think. GNU tar is already crowded with options and moreover, the approach just explained gives you a great deal of control already.

‘-p’
‘--same-permissions’
‘--preserve-permissions’

Extract all protection information. This option causes tar to set the modes (access permissions) of extracted files exactly as recorded in the archive. If this option is not used, the current umask setting limits the permissions on extracted files. This option is by default enabled when tar is executed by a superuser. This option is meaningless with ‘--list’ (‘-t’).

‘--preserve’
Same as both ‘--same-permissions’ and ‘--same-order’.

The ‘--preserve’ option has no equivalent short option name.

It is equivalent to ‘--same-permissions’ plus ‘--same-order’.

8.3 Making tar Archives More Portable

Creating a tar archive on a particular system that is meant to be useful later on many other machines and with other versions of tar is more challenging than you might think. tar archive formats have been evolving since the first versions of Unix. Many such formats are around, and are not always compatible with each other. This section discusses a few problems, and gives some advice about making tar archives more portable.

One golden rule is simplicity. For example, limit your tar archives to contain only regular files and directories, avoiding other kind of special files. Do not attempt to save sparse files or contiguous files as such. Let’s discuss a few more problems, in turn.

8.3.1 Portable Names

Use portable file and member names. A name is portable if it contains only ASCII letters and digits, ‘/’, ‘.’, ‘_’, and ‘-’; it cannot be empty, start with ‘-’ or ‘//’, or contain ‘/-’. Avoid deep directory nesting. For portability to old Unix hosts, limit your file name components to 14 characters or less.

If you intend to have your tar archives to be read under MSDOS, you should not rely on case distinction for file names, and you might use the GNU doschk program for helping you further diagnosing illegal MSDOS names, which are even more limited than System V’s.

8.3.2 Symbolic Links

Normally, when tar archives a symbolic link, it writes a block to the archive naming the target of the link. In that way, the tar archive is a faithful record of the file system contents. ‘--dereference’ (‘-h’) is used with ‘--create’ (‘-c’), and causes tar to archive the files symbolic links point to, instead of the links themselves. When this option is used, when tar encounters a symbolic link, it will archive the linked-to file, instead of simply recording the presence of a symbolic link.

The name under which the file is stored in the file system is not recorded in the archive. To record both the symbolic link name and the file name in the system, archive the file under both names. If all links were recorded automatically by tar, an extracted file might be linked to a file name that no longer exists in the file system.

If a linked-to file is encountered again by tar while creating the same archive, an entire second copy of it will be stored. (This might be considered a bug.)

So, for portable archives, do not archive symbolic links as such, and use ‘--dereference’ (‘-h’): many systems do not support symbolic links, and moreover, your distribution might be unusable if it contains unresolved symbolic links.

8.3.3 Old V7 Archives

Certain old versions of tar cannot handle additional information recorded by newer tar programs. To create an archive in V7 format (not ANSI), which can be read by these old versions, specify the ‘--format=v7’ option in conjunction with the ‘--create’ (‘-c’) (tar also accepts ‘--portability’ or ‘--old-archive’ for this option). When you specify it, tar leaves out information about directories, pipes, fifos, contiguous files, and device files, and specifies file ownership by group and user IDs instead of group and user names.

When updating an archive, do not use ‘--format=v7’ unless the archive was created using this option.

In most cases, a new format archive can be read by an old tar program without serious trouble, so this option should seldom be needed. On the other hand, most modern tars are able to read old format archives, so it might be safer for you to always use ‘--format=v7’ for your distributions. Notice, however, that ‘ustar’ format is a better alternative, as it is free from many of ‘v7’’s drawbacks.

8.3.4 Ustar Archive Format

Archive format defined by POSIX.1-1988 specification is called ustar. Although it is more flexible than the V7 format, it still has many restrictions (See Chapter 8 [Formats], page 117, for the detailed description of ustar format). Along with V7 format, ustar format is a good choice for archives intended to be read with other implementations of tar.

To create archive in ustar format, use ‘--format=ustar’ option in conjunction with the ‘--create’ (‘-c’).

8.3.5 GNU and old GNU tar format

GNU tar was based on an early draft of the POSIX 1003.1 ustar standard. GNU extensions to tar, such as the support for file names longer than 100 characters, use portions of the tar header record which were specified in that POSIX draft as unused. Subsequent changes in POSIX have allocated the same parts of the header record for other purposes. As a result, GNU tar format is incompatible with the current POSIX specification, and with tar programs that follow it.

In the majority of cases, tar will be configured to create this format by default. This will change in the future releases, since we plan to make ‘POSIX’ format the default.

To force creation a GNU tar archive, use option ‘--format=gnu’.

8.3.6 GNU tar and POSIX tar

Starting from version 1.14 GNU tar features full support for POSIX.1-2001 archives.

A POSIX conformant archive will be created if tar was given ‘--format=posix’ (‘--format=pax’) option. No special option is required to read and extract from a POSIX archive.

8.3.6.1 Controlling Extended Header Keywords

‘--pax-option=keyword-list’

Handle keywords in PAX extended headers. This option is equivalent to ‘-o’ option of the pax utility.

Keyword-list is a comma-separated list of keyword options, each keyword option taking one of the following forms:

delete=pattern

When used with one of archive-creation commands, this option instructs tar to omit from extended header records that it produces any keywords matching the string pattern. When used in extract or list mode, this option instructs tar to ignore any keywords matching the given pattern in the extended header records. In both cases, matching is performed using the pattern matching notation described in POSIX 1003.2, 3.13 (see Section 6.5 [wildcards], page 93). For example:

     --pax-option delete=security.*

would suppress security-related information.

exthdr.name=string

This keyword allows user control over the name that is written into the ustar header blocks for the extended headers. The name is obtained from string after making the following substitutions:

Meta-character  Replaced By

%d The directory name of the file, equivalent to the result of the dirname utility on the translated pathname.

%f The filename of the file, equivalent to the result of the basename utility on the translated pathname.

%p  The process ID of the tar process.
%%  A ‘%’ character.

Any other ‘%’ characters in string produce undefined results. If no option ‘exthdr.name=string’ is specified, tar will use the following default value:

     %d/PaxHeaders.%p/%f
globexthdr.name=string

This keyword allows user control over the name that is written into the ustar header blocks for global extended header records. The name is obtained from the contents of string, after making the following substitutions:

Meta-character  Replaced By

%n An integer that represents the sequence number of the global extended header record in the archive, starting at 1.

%p  The process ID of the tar process.
%%  A ‘%’ character.

Any other ‘%’ characters in string produce undefined results.

If no option ‘globexthdr.name=string’ is specified, tar will use the following default value:

     $TMPDIR/GlobalHead.%p.%n

where ‘$TMPDIR’ represents the value of the TMPDIR environment variable. If TMPDIR is not set, tar uses ‘/tmp’.

keyword=value

When used with one of archive-creation commands, these keyword/value pairs will be included at the beginning of the archive in a global extended header record. When used with one of archive-reading commands, tar will behave as if it has encountered these keyword/value pairs at the beginning of the archive in a global extended header record.

keyword:=value

When used with one of archive-creation commands, these keyword/value pairs will be included as records at the beginning of an extended header for each file. This is effectively equivalent to keyword=value form except that it creates no global extended header records. When used with one of archive-reading commands, tar will behave as if these keyword/value pairs were included as records at the end of each extended header; thus, they will override any global or file-specific extended header record keywords of the same names. For example, in the command:

     tar --format=posix --create \
         --file archive --pax-option gname:=user .

the group name will be forced to a new value for all files stored in the archive.

8.3.7 Checksumming Problems

SunOS and HP-UX tar fail to accept archives created using GNU tar and containing non-ASCII file names, that is, file names having characters with the eight bit set, because they use signed checksums, while GNU tar uses unsigned checksums while creating archives, as per POSIX standards. On reading, GNU tar computes both checksums and accept any. It is somewhat worrying that a lot of people may go around doing backup of their files using faulty (or at least non-standard) software, not learning about it until it’s time to restore their missing files with an incompatible file extractor, or vice versa.

GNU tar compute checksums both ways, and accept any on read, so GNU tar can read Sun tapes even with their wrong checksums. GNU tar produces the standard checksum, however, raising incompatibilities with Sun. That is to say, GNU tar has not been modified to produce incorrect archives to be read by buggy tar’s. I’ve been told that more recent Sun tar now read standard archives, so maybe Sun did a similar patch, after all?

The story seems to be that when Sun first imported tar sources on their system, they recompiled it without realizing that the checksums were computed differently, because of a change in the default signing of char’s in their compiler. So they started computing checksums wrongly. When they later realized their mistake, they merely decided to stay compatible with it, and with themselves afterwards. Presumably, but I do not really know, HP-UX has chosen that their tar archives to be compatible with Sun’s. The current standards do not favor Sun tar format. In any case, it now falls on the shoulders of SunOS and HP-UX users to get a tar able to read the good archives they receive.

8.3.8 Large or Negative Values

(This message will disappear, once this node revised.)

The above sections suggest to use ‘oldest possible’ archive format if in doubt. However, sometimes it is not possible. If you attempt to archive a file whose metadata cannot be represented using required format, GNU tar will print error message and ignore such a file. You will than have to switch to a format that is able to handle such values. The format summary table (see Chapter 8 [Formats], page 117) will help you to do so.

In particular, when trying to archive files larger than 8GB or with timestamps not in the range 1970-01-01 00:00:00 through 2242-03-16 12:56:31 utc, you will have to chose between GNU and POSIX archive formats. When considering which format to choose, bear in mind that the GNU format uses two’s-complement base-256 notation to store values that do not fit into standard ustar range. Such archives can generally be read only by a GNU tar implementation. Moreover, they sometimes cannot be correctly restored on another hosts even by GNU tar. For example, using two’s complement representation for negative time stamps that assumes a signed 32-bit time_t generates archives that are not portable to hosts with differing time_t representations.

On the other hand, POSIX archives, generally speaking, can be extracted by any tar implementation that understands older ustar format. The only exception are files larger than 8GB.

8.3.9 How to Extract GNU-Specific Data Using Other tar Implementations

In previous sections you became acquainted with various quirks necessary to make your archives portable. Sometimes you may need to extract archives containing GNU-specific members using some third-party tar implementation or an older version of GNU tar. Of course your best bet is to have GNU tar installed, but if it is for some reason impossible, this section will explain how to cope without it.

When we speak about GNU-specific members we mean two classes of them: members split between the volumes of a multi-volume archive and sparse members. You will be able to always recover such members if the archive is in PAX format. In addition split members can be recovered from archives in old GNU format. The following subsections describe the required procedures in detail.

8.3.9.1 Extracting Members Split Between Volumes

If a member is split between several volumes of an old GNU format archive most third party tar implementation will fail to extract it. To extract it, use tarcat program (see Section 9.6.3 [Tarcat], page 154). This program

is available from GNU tar home page (http://www.gnu.org/software/

tar/utils/tarcat.html). It concatenates several archive volumes into a single valid archive. For example, if you have three volumes named from ‘vol-1.tar’ to ‘vol-3.tar’, you can do the following to extract them using a third-party tar:

     $ tarcat vol-1.tar vol-2.tar vol-3.tar | tar xf -

You could use this approach for most (although not all) PAX format archives as well. However, extracting split members from a PAX archive is a much easier task, because PAX volumes are constructed in such a way that each part of a split member is extracted to a different file by tar implementations that are not aware of GNU extensions. More specifically, the very first part retains its original name, and all subsequent parts are named using the pattern:

     %d/GNUFileParts.%p/%f.%n

where symbols preceeded by ‘%’ are macro characters that have the following meaning:

Meta-character  Replaced By

%d The directory name of the file, equivalent to the result of the dirname utility on its full name.

%f The file name of the file, equivalent to the result of the basename utility on its full name.

%p The process ID of the tar process that created the archive.

%n  Ordinal number of this particular part.

For example, if the file ‘var/longfile’ was split during archive creation between three volumes, and the creator tar process had process ID ‘27962’, then the member names will be:

     var/longfile
     var/GNUFileParts.27962/longfile.1
     var/GNUFileParts.27962/longfile.2

When you extract your archive using a third-party tar, these files will be created on your disk, and the only thing you will need to do to restore your file in its original form is concatenate them in the proper order, for example:

     $ cd var
     $ cat GNUFileParts.27962/longfile.1 \
       GNUFileParts.27962/longfile.2 >> longfile
     $ rm -f GNUFileParts.27962

Notice, that if the tar implementation you use supports PAX format archives, it will probably emit warnings about unknown keywords during extraction. They will look like this:

     Tar file too small
     Unknown extended header keyword 'GNU.volume.filename' ignored.
     Unknown extended header keyword 'GNU.volume.size' ignored.
     Unknown extended header keyword 'GNU.volume.offset' ignored.

You can safely ignore these warnings.

If your tar implementation is not PAX-aware, you will get more warnings and more files generated on your disk, e.g.:

     $ tar xf vol-1.tar
     var/PaxHeaders.27962/longfile: Unknown file type 'x', extracted as
     normal file
     Unexpected EOF in archive
     $ tar xf vol-2.tar
     tmp/GlobalHead.27962.1: Unknown file type 'g', extracted as normal file
     GNUFileParts.27962/PaxHeaders.27962/sparsefile.1: Unknown file type
     'x', extracted as normal file

Ignore these warnings. The ‘PaxHeaders.*’ directories created will contain files with extended header keywords describing the extracted files. You can delete them, unless they describe sparse members. Read further to learn more about them.

8.3.9.2 Extracting Sparse Members

Any tar implementation will be able to extract sparse members from a PAX archive. However, the extracted files will be condensed, i.e., any zero blocks will be removed from them. When we restore such a condensed file to its original form, by adding zero bloks (or holes) back to their original locations, we call this process expanding a compressed sparse file.

To expand a file, you will need a simple auxiliary program called xsparse. It is available in source form from GNU tar home page

(http://www.gnu.org/software/tar/utils/xsparse.html).

Let’s begin with archive members in sparse format version 1.0^{1}, which are the easiest to expand. The condensed file will contain both file map and file data, so no additional data will be needed to restore it. If the original file name was ‘dir/name’, then the condensed file will be named ‘dir/GNUSparseFile.n/name’, where n is a decimal number^{2}.

To expand a version 1.0 file, run xsparse as follows:

1 See [PAX 1], page 176.

2 technically speaking, n is a process ID of the tar process which created the archive

(see Section 8.3.6.1 [PAX keywords], page 126).

     $ xsparse `cond-file'

where ‘cond-file’ is the name of the condensed file. The utility will deduce the name for the resulting expanded file using the following algorithm:

1. If ‘cond-file’ does not contain any directories, ‘../cond-file’ will be used;

2. If ‘cond-file’ has the form ‘dir/t/name’, where both t and name are simple names, with no ‘/’ characters in them, the output file name will

    be ‘dir/name’.

3. Otherwise, if ‘cond-file’ has the form ‘dir/name’, the output file name will be ‘name’.

In the unlikely case when this algorithm does not suite your needs, you can explicitly specify output file name as a second argument to the command:

     $ xsparse `cond-file' `out-file'

It is often a good idea to run xsparse in dry run mode first. In this mode, the command does not actually expand the file, but verbosely lists all actions it would be taking to do so. The dry run mode is enabled by ‘-n’ command line argument:

     $ xsparse -n /home/gray/GNUSparseFile.6058/sparsefile
     Reading v.1.0 sparse map
     Expanding file `/home/gray/GNUSparseFile.6058/sparsefile' to
     `/home/gray/sparsefile'
     Finished dry run

To actually expand the file, you would run:

     $ xsparse /home/gray/GNUSparseFile.6058/sparsefile

The program behaves the same way all UNIX utilities do: it will keep quiet unless it has simething important to tell you (e.g. an error condition or something). If you wish it to produce verbose output, similar to that from the dry run mode, use ‘-v’ option:

     $ xsparse -v /home/gray/GNUSparseFile.6058/sparsefile
     Reading v.1.0 sparse map
     Expanding file `/home/gray/GNUSparseFile.6058/sparsefile' to
     `/home/gray/sparsefile'
     Done

Additionally, if your tar implementation has extracted the extended headers for this file, you can instruct xstar to use them in order to verify the integrity of the expanded file. The option ‘-x’ sets the name of the extended header file to use. Continuing our example:

     $ xsparse -v -x /home/gray/PaxHeaders.6058/sparsefile \
       /home/gray/GNUSparseFile.6058/sparsefile
     Reading extended header file
     Found variable GNU.sparse.major = 1
     Found variable GNU.sparse.minor = 0
     Found variable GNU.sparse.name = sparsefile
     Found variable GNU.sparse.realsize = 217481216
     Reading v.1.0 sparse map
     Expanding file `/home/gray/GNUSparseFile.6058/sparsefile' to
     `/home/gray/sparsefile'
     Done

An extended header is a special tar archive header that precedes an archive member and contains a set of variables, describing the member properties that cannot be stored in the standard ustar header. While optional for expanding sparse version 1.0 members, use of extended headers is mandatory when expanding sparse members in older sparse formats: v.0.0 and v.0.1 (The sparse formats are described in detail in [Sparse Formats], page 174.) So, for this format, the question is: how to obtain extended headers from the archive?

If you use a tar implementation that does not support PAX format, extended headers for each member will be extracted as a separate file. If we represent the member name as ‘dir/name’, then the extended header file will be named ‘dir/PaxHeaders.n/name’, where n is an integer number.

Things become more difficult if your tar implementation does support PAX headers, because in this case you will have to manually extract the headers. We recommend the following algorithm:

1. Consult the documentation of your tar implementation for an option that prints block numbers along with the archive listing (analogous to GNU tar’s ‘-R’ option). For example, star has ‘-block-number’.

2. Obtain verbose listing using the ‘block number’ option, and find block numbers of the sparse member in question and the member immediately following it. For example, running star on our archive we obtain:

         $ star -t -v -block-number -f arc.tar
         ...
         star: Unknown extended header keyword 'GNU.sparse.size' ignored.
         star: Unknown extended header keyword 'GNU.sparse.numblocks' ignored.
         star: Unknown extended header keyword 'GNU.sparse.name' ignored.
         star: Unknown extended header keyword 'GNU.sparse.map' ignored.
         block        56:  425984 -rw-r--r--  gray/users Jun 25 14:46 2006 GNUSparseFile.28124/sparsefile
         block       897:   65391 -rw-r--r--  gray/users Jun 24 20:06 2006 README
         ...

(as usual, ignore the warnings about unknown keywords.)

3. Let size be the size of the sparse member, Bs be its block number and

Bn be the block number of the next member. Compute:

         N = Bs - Bn - size/512 - 2

This number gives the size of the extended header part in tar blocks. In our example, this formula gives: 897 - 56 - 425984 / 512 - 2 = 7.

4. Use dd to extract the headers:

         dd if=archive of=hname bs=512 skip=Bs count=N

where archive is the archive name, hname is a name of the file to store the extended header in, Bs and N are computed in previous steps. In our example, this command will be

         $ dd if=arc.tar of=xhdr bs=512 skip=56 count=7

Finally, you can expand the condensed file, using the obtained header:

     $ xsparse -v -x xhdr GNUSparseFile.6058/sparsefile
     Reading extended header file
     Found variable GNU.sparse.size = 217481216
     Found variable GNU.sparse.numblocks = 208
     Found variable GNU.sparse.name = sparsefile
     Found variable GNU.sparse.map = 0,2048,1050624,2048,...
     Expanding file `GNUSparseFile.28124/sparsefile' to `sparsefile'
     Done

8.4 Comparison of tar and cpio

(This message will disappear, once this node revised.)

The cpio archive formats, like tar, do have maximum pathname lengths. The binary and old ASCII formats have a max path length of 256, and the new ASCII and CRC ASCII formats have a max path length of 1024. GNU cpio can read and write archives with arbitrary pathname lengths, but other cpio implementations may crash unexplainedly trying to read them.

tar handles symbolic links in the form in which it comes in BSD; cpio doesn’t handle symbolic links in the form in which it comes in System V prior to SVR4, and some vendors may have added symlinks to their system without enhancing cpio to know about them. Others may have enhanced it in a way other than the way I did it at Sun, and which was adopted by AT&T (and which is, I think, also present in the cpio that Berkeley picked up from AT&T and put into a later BSD release—I think I gave them my changes).

(SVR4 does some funny stuff with tar; basically, its cpio can handle tar format input, and write it on output, and it probably handles symbolic links. They may not have bothered doing anything to enhance tar as a result.)

cpio handles special files; traditional tar doesn’t.

tar comes with V7, System III, System V, and BSD source; cpio comes only with System III, System V, and later BSD (4.3-tahoe and later).

tar’s way of handling multiple hard links to a file can handle file systems that support 32-bit inumbers (e.g., the BSD file system); cpios way requires you to play some games (in its "binary" format, i-numbers are only 16 bits, and in its "portable ASCII" format, they’re 18 bits—it would have to play games with the "file system ID" field of the header to make sure that the file system ID/i-number pairs of different files were always different), and I don’t know which cpios, if any, play those games. Those that don’t might get confused and think two files are the same file when they’re not, and make hard links between them.

tars way of handling multiple hard links to a file places only one copy of the link on the tape, but the name attached to that copy is the only one you can use to retrieve the file; cpios way puts one copy for every link, but you can retrieve it using any of the names.

What type of check sum (if any) is used, and how is this calculated. See the attached manual pages for tar and cpio format. tar uses a

checksum which is the sum of all the bytes in the tar header for a file; cpio uses no checksum.

If anyone knows why cpio was made when tar was present at the unix scene, It wasn’t. cpio first showed up in PWB/UNIX 1.0; no generally-available

version of UNIX had tar at the time. I don’t know whether any version that was generally available within AT&T had tar, or, if so, whether the people within AT&T who did cpio knew about it.

On restore, if there is a corruption on a tape tar will stop at that point, while cpio will skip over it and try to restore the rest of the files.

The main difference is just in the command syntax and header format. tar is a little more tape-oriented in that everything is blocked to start

on a record boundary.

Is there any differences between the ability to recover crashed archives between the two of them. (Is there any chance of recovering crashed archives at all.) Theoretically it should be easier under tar since the blocking lets you

find a header with some variation of ‘dd skip=nn’. However, modern cpio’s and variations have an option to just search for the next file header after an error with a reasonable chance of resyncing. However, lots of tape driver software won’t allow you to continue past a media error which should be the only reason for getting out of sync unless a file changed sizes while you were writing the archive.

If anyone knows why cpio was made when tar was present at the unix scene, please tell me about this too. Probably because it is more media efficient (by not blocking everything

and using only the space needed for the headers where tar always uses 512 bytes per file header) and it knows how to archive special files.

You might want to look at the freely available alternatives. The major ones are afio, GNU tar, and pax, each of which have their own extensions with some backwards compatibility.

Sparse files were tarred as sparse files (which you can easily test, because the resulting archive gets smaller, and GNU cpio can no longer read it).

136 GNU tar 1.16.1

9 Tapes and Other Archive Media

(This message will disappear, once this node revised.)

A few special cases about tape handling warrant more detailed description. These special cases are discussed below.

Many complexities surround the use of tar on tape drives. Since the creation and manipulation of archives located on magnetic tape was the original purpose of tar, it contains many features making such manipulation easier.

Archives are usually written on dismountable media—tape cartridges, mag tapes, or floppy disks.

The amount of data a tape or disk holds depends not only on its size, but also on how it is formatted. A 2400 foot long reel of mag tape holds 40 megabytes of data when formatted at 1600 bits per inch. The physically smaller EXABYTE tape cartridge holds 2.3 gigabytes.

Magnetic media are re-usable—once the archive on a tape is no longer needed, the archive can be erased and the tape or disk used over. Media quality does deteriorate with use, however. Most tapes or disks should be discarded when they begin to produce data errors. EXABYTE tape cartridges should be discarded when they generate an error count (number of non-usable bits) of more than 10k.

Magnetic media are written and erased using magnetic fields, and should be protected from such fields to avoid damage to stored data. Sticking a floppy disk to a filing cabinet using a magnet is probably not a good idea.

9.1 Device Selection and Switching

(This message will disappear, once this node revised.)

‘-f [hostname:]file’
‘--file=[hostname:]file’

Use archive file or device file on hostname. This option is used to specify the file name of the archive tar works on. If the file name is ‘-’, tar reads the archive from standard input (when

listing or extracting), or writes it to standard output (when creating). If the ‘-’ file name is given when updating an archive, tar will read the original archive from its standard input, and will write the entire new archive to its standard output.

If the file name contains a ‘:’, it is interpreted as ‘hostname:file name’. If the hostname contains an at sign (‘@’), it is treated as ‘user@hostname:file name’. In either case, tar will invoke the command rsh (or remsh) to start up an /usr/libexec/rmt on the remote machine. If you give an alternate login name, it will be given to the rsh. Naturally, the remote machine must have an executable /usr/libexec/rmt. This program is free software from the University of California, and a copy of the source code can be found with the sources for tar; it’s compiled and installed by default. The exact path to this utility is determined when configuring the package. It is ‘prefix/libexec/rmt’, where prefix stands for your installation prefix. This location may also be overridden at runtime by using ‘rmt-command=command’ option (See Section 3.4.2 [Option Summary], page 29, for detailed description of this option. See Section 9.2 [Remote Tape Server], page 139, for the description of rmt command).

If this option is not given, but the environment variable TAPE is set, its value is used; otherwise, old versions of tar used a default archive name (which was picked when tar was compiled). The default is normally set up to be the first tape drive or other transportable I/O medium on the system.

Starting with version 1.11.5, GNU tar uses standard input and standard output as the default device, and I will not try anymore supporting automatic device detection at installation time. This was failing really in too many cases, it was hopeless. This is now completely left to the installer to override standard input and standard output for default device, if this seems preferable. Further, I think most actual usages of tar are done with pipes or disks, not really tapes, cartridges or diskettes.

Some users think that using standard input and output is running after trouble. This could lead to a nasty surprise on your screen if you forget to specify an output file name—especially if you are going through a network or terminal server capable of buffering large amounts of output. We had so many bug reports in that area of configuring default tapes automatically, and so many contradicting requests, that we finally consider the problem to be portably intractable. We could of course use something like ‘/dev/tape’ as a default, but this is also running after various kind of trouble, going from hung processes to accidental destruction of real tapes. After having seen all this mess, using standard input and output as a default really sounds like the only clean choice left, and a very useful one too.

GNU tar reads and writes archive in records, I suspect this is the main reason why block devices are preferred over character devices. Most probably, block devices are more efficient too. The installer could also check for

‘DEFTAPE’ in ‘<sys/mtio.h>’.
‘--force-local’

Archive file is local even if it contains a colon.

‘--rsh-command=command’

Use remote command instead of rsh. This option exists so that people who use something other than the standard rsh (e.g., a Kerberized rsh) can access a remote device. When this command is not used, the shell command found when the tar program was installed is used instead. This is the first

          found of ‘/usr/ucb/rsh’, ‘/usr/bin/remsh’, ‘/usr/bin/rsh’,

‘/usr/bsd/rsh’ or ‘/usr/bin/nsh’. The installer may have overridden this by defining the environment variable RSH at installation time.

‘-[0-7][lmh]’

Specify drive and density.

‘-M’
‘--multi-volume’

Create/list/extract multi-volume archive.

This option causes tar to write a multi-volume archive—one that may be larger than will fit on the medium used to hold it. See Section 9.6.1 [Multi-Volume Archives], page 151.

‘-L num’
‘--tape-length=num’

Change tape after writing num x 1024 bytes.

This option might be useful when your tape drivers do not properly detect end of physical tapes. By being slightly conservative on the maximum tape length, you might avoid the problem entirely.

‘-F file’
‘--info-script=file’
‘--new-volume-script=file’

Execute ‘file’ at end of each tape. This implies ‘--multi-volume’ (‘-M’). See [info-script], page 152, for a detailed description of this option.

9.2 The Remote Tape Server

In order to access the tape drive on a remote machine, tar uses the remote tape server written at the University of California at Berkeley. The remote tape server must be installed as ‘prefix/libexec/rmt’ on any machine whose tape drive you want to use. tar calls rmt by running an rsh or remsh to the remote machine, optionally using a different login name if one is supplied.

A copy of the source for the remote tape server is provided. It is Copyright \begin{matrix}c \\ ◯ \end{matrix} 1983 by the Regents of the University of California, but can be freely distributed. It is compiled and installed by default.

Unless you use the ‘--absolute-names’ (‘-P’) option, GNU tar will not allow you to create an archive that contains absolute file names (a file name beginning with ‘/’.) If you try, tar will automatically remove the leading ‘/’ from the file names it stores in the archive. It will also type a warning message telling you what it is doing.

When reading an archive that was created with a different tar program, GNU tar automatically extracts entries in the archive which have absolute file names as if the file names were not absolute. This is an important feature.

A visitor here once gave a tar tape to an operator to restore; the operator used Sun tar instead of GNU tar, and the result was that it replaced large portions of our ‘/bin’ and friends with versions from the tape; needless to say, we were unhappy about having to recover the file system from backup tapes.

For example, if the archive contained a file ‘/usr/bin/computoy’, GNU tar would extract the file to ‘usr/bin/computoy’, relative to the current directory. If you want to extract the files in an archive to the same absolute names that they had when the archive was created, you should do a ‘cd /’ before extracting the files from the archive, or you should either use the ‘--absolute-names’ option, or use the command ‘tar -C / ...’.

Some versions of Unix (Ultrix 3.1 is known to have this problem), can claim that a short write near the end of a tape succeeded, when it actually failed. This will result in the -M option not working correctly. The best workaround at the moment is to use a significantly larger blocking factor than the default 20.

In order to update an archive, tar must be able to backspace the archive in order to reread or rewrite a record that was just read (or written). This is currently possible only on two kinds of files: normal disk files (or any other file that can be backspaced with ‘lseek’), and industry-standard 9-track magnetic tape (or any other kind of tape that can be backspaced with the

MTIOCTOP ioctl.

This means that the ‘--append’, ‘--concatenate’, and ‘--delete’ commands will not work on any other kind of file. Some media simply cannot be backspaced, which means these commands and options will never be able to work on them. These non-backspacing media include pipes and cartridge tape drives.

Some other media can be backspaced, and tar will work on them once tar is modified to do so.

Archives created with the ‘--multi-volume’, ‘--label’, and ‘--incremental’ (‘-G’) options may not be readable by other version of tar. In particular, restoring a file that was split over a volume boundary will require some careful work with dd, if it can be done at all. Other versions of tar may also create an empty file whose name is that of the volume header. Some versions of tar may create normal files instead of directories archived with the ‘--incremental’ (‘-G’) option.

9.3 Some Common Problems and their Solutions

errors from system: permission denied no such file or directory not owner

errors from tar: directory checksum error header format error

errors from media/system: i/o error device busy

9.4 Blocking

(This message will disappear, once this node revised.)

Block and record terminology is rather confused, and it is also confusing to the expert reader. On the other hand, readers who are new to the field have a fresh mind, and they may safely skip the next two paragraphs, as the remainder of this manual uses those two terms in a quite consistent way.

John Gilmore, the writer of the public domain tar from which GNU tar was originally derived, wrote (June 1995):

The nomenclature of tape drives comes from IBM, where I believe they were invented for the IBM 650 or so. On IBM mainframes, what is recorded on tape are tape blocks. The logical organization of data is into records. There are various ways of putting records into blocks, including F (fixed sized records), V (variable sized records), FB (fixed blocked: fixed size records, n to a block), VB (variable size records, n to a block), VSB (variable spanned blocked: variable sized records that can occupy more than one block), etc. The JCL ‘DD RECFORM=’ parameter specified this to the operating system. The Unix man page on tar was totally confused about this. When I wrote PD TAR, I used the historically correct terminology (tar writes data records, which are grouped into blocks). It appears that the bogus terminology made it into POSIX (no surprise here), and now François has migrated that terminology back into the source code too. The term physical block means the basic transfer chunk from or to a

device, after which reading or writing may stop without anything being lost. In this manual, the term block usually refers to a disk physical block, assuming that each disk block is 512 bytes in length. It is true that some disk devices have different physical blocks, but tar ignore these differences in its own format, which is meant to be portable, so a tar block is always 512 bytes in length, and block always mean a tar block. The term logical block often represents the basic chunk of allocation of many disk blocks as a single entity, which the operating system treats somewhat atomically; this concept is only barely used in GNU tar.

The term physical record is another way to speak of a physical block, those two terms are somewhat interchangeable. In this manual, the term record usually refers to a tape physical block, assuming that the tar archive is kept on magnetic tape. It is true that archives may be put on disk or used with pipes, but nevertheless, tar tries to read and write the archive one record at a time, whatever the medium in use. One record is made up of an integral number of blocks, and this operation of putting many disk blocks into a single tape block is called reblocking, or more simply, blocking. The term logical record refers to the logical organization of many characters into something meaningful to the application. The term unit record describes a small set of characters which are transmitted whole to or by the application, and often refers to a line of text. Those two last terms are unrelated to what we call a record in GNU tar.

When writing to tapes, tar writes the contents of the archive in chunks known as records. To change the default blocking factor, use the

‘--blocking-factor=512-size’ (‘-b 512-size’) option. Each record will

then be composed of 512-size blocks. (Each tar block is 512 bytes. See [Standard], page 165.) Each file written to the archive uses at least one full record. As a result, using a larger record size can result in more wasted space for small files. On the other hand, a larger record size can often be read and written much more efficiently.

Further complicating the problem is that some tape drives ignore the blocking entirely. For these, a larger record size can still improve performance (because the software layers above the tape drive still honor the blocking), but not as dramatically as on tape drives that honor blocking.

When reading an archive, tar can usually figure out the record size on itself. When this is the case, and a non-standard record size was used when the archive was created, tar will print a message about a non-standard blocking factor, and then operate normally. On some tape devices, however, tar cannot figure out the record size itself. On most of those, you can specify a blocking factor (with ‘--blocking-factor’) larger than the actual blocking factor, and then use the ‘--read-full-records’ (‘-B’) option. (If you specify a blocking factor with ‘--blocking-factor’ and don’t use the ‘--read-full-records’ option, then tar will not attempt to figure out the recording size itself.) On some devices, you must always specify the record size exactly with ‘--blocking-factor’ when reading, because tar cannot figure it out. In any case, use ‘--list’ (‘-t’) before doing any extractions to see whether tar is reading the archive correctly.

tar blocks are all fixed size (512 bytes), and its scheme for putting them into records is to put a whole number of them (one or more) into each record. tar records are all the same size; at the end of the file there’s a block containing all zeros, which is how you tell that the remainder of the last record(s) are garbage.

In a standard tar file (no options), the block size is 512 and the record size is 10240, for a blocking factor of 20. What the ‘--blocking-factor’ option does is sets the blocking factor, changing the record size while leaving the block size at 512 bytes. 20 was fine for ancient 800 or 1600 bpi reel-to-reel tape drives; most tape drives these days prefer much bigger records in order to stream and not waste tape. When writing tapes for myself, some tend to use a factor of the order of 2048, say, giving a record size of around one megabyte.

If you use a blocking factor larger than 20, older tar programs might not be able to read the archive, so we recommend this as a limit to use in practice. GNU tar, however, will support arbitrarily large record sizes, limited only by the amount of virtual memory or the physical characteristics of the tape device.

9.4.1 Format Variations

(This message will disappear, once this node revised.)

Format parameters specify how an archive is written on the archive media. The best choice of format parameters will vary depending on the type and number of files being archived, and on the media used to store the archive.

To specify format parameters when accessing or creating an archive, you can use the options described in the following sections. If you do not specify any format parameters, tar uses default parameters. You cannot modify a compressed archive. If you create an archive with the ‘--blocking-factor’ option specified (see Section 9.4.2 [Blocking Factor], page 143), you must specify that blocking-factor when operating on the archive. See Chapter 8 [Formats], page 117, for other examples of format parameter considerations.

9.4.2 The Blocking Factor of an Archive

(This message will disappear, once this node revised.)

The data in an archive is grouped into blocks, which are 512 bytes. Blocks are read and written in whole number multiples called records. The number of blocks in a record (i.e., the size of a record in units of 512 bytes) is called

the blocking factor. The ‘--blocking-factor=512-size’ (‘-b 512-size’)

option specifies the blocking factor of an archive. The default blocking factor is typically 20 (i.e., 10240 bytes), but can be specified at installation. To find out the blocking factor of an existing archive, use ‘tar --list --file=archive-name’. This may not work on some devices.

Records are separated by gaps, which waste space on the archive media. If you are archiving on magnetic tape, using a larger blocking factor (and therefore larger records) provides faster throughput and allows you to fit more data on a tape (because there are fewer gaps). If you are archiving on cartridge, a very large blocking factor (say 126 or more) greatly increases performance. A smaller blocking factor, on the other hand, may be useful when archiving small files, to avoid archiving lots of nulls as tar fills out the archive to the end of the record. In general, the ideal record size depends on the size of the inter-record gaps on the tape you are using, and the average size of the files you are archiving. See Section 2.6 [create], page 10, for information on writing archives.

Archives with blocking factors larger than 20 cannot be read by very old versions of tar, or by some newer versions of tar running on old machines with small address spaces. With GNU tar, the blocking factor of an archive is limited only by the maximum record size of the device containing the archive, or by the amount of available virtual memory.

Also, on some systems, not using adequate blocking factors, as sometimes imposed by the device drivers, may yield unexpected diagnostics. For example, this has been reported:

     Cannot write to /dev/dlt: Invalid argument

In such cases, it sometimes happen that the tar bundled by the system is aware of block size idiosyncrasies, while GNU tar requires an explicit specification for the block size, which it cannot guess. This yields some people to consider GNU tar is misbehaving, because by comparison, the bundle tar works OK. Adding -b 256, for example, might resolve the problem.

If you use a non-default blocking factor when you create an archive, you must specify the same blocking factor when you modify that archive. Some archive devices will also require you to specify the blocking factor when reading that archive, however this is not typically the case. Usually, you can use ‘--list’ (‘-t’) without specifying a blocking factor—tar reports a non-default record size and then lists the archive members as it would normally. To extract files from an archive with a non-standard blocking factor (particularly if you’re not sure what the blocking factor is), you can usually use the ‘--read-full-records’ (‘-B’) option while specifying a blocking factor larger then the blocking factor of the archive (i.e.,

‘tar --extract --read-full-records --blocking-factor=300’. See Sec-

tion 2.7 [list], page 15, for more information on the ‘--list’ (‘-t’) operation. See Section 4.4.1 [Reading], page 61, for a more detailed explanation of that option.

‘--blocking-factor=number’
‘-b number’

Specifies the blocking factor of an archive. Can be used with any operation, but is usually not necessary with ‘--list’ (‘-t’).

Device blocking

‘-b blocks’
‘--blocking-factor=blocks’

Set record size to blocks ∗ 512 bytes.

This option is used to specify a blocking factor for the archive. When reading or writing the archive, tar, will do reads and writes of the archive in records of block ∗ 512 bytes. This is true even when the archive is compressed. Some devices requires that all write operations be a multiple of a certain size, and so, tar pads the archive out to the next record boundary. The default blocking factor is set when tar is compiled, and is typically 20. Blocking factors larger than 20 cannot be read by very old versions of tar, or by some newer versions of tar running on old machines with small address spaces. With a magnetic tape, larger records give faster throughput and fit more data on a tape (because there are fewer inter-record gaps). If the archive is in a disk file or a pipe, you may want to specify a smaller blocking factor, since a large one will result in a large number of null bytes at the end of the archive. When writing cartridge or other streaming tapes, a much larger blocking factor (say 126 or more) will greatly increase performance. However, you must specify the same blocking factor when reading or updating the archive. Apparently, Exabyte drives have a physical block size of 8K bytes. If we choose our blocksize as a multiple of 8k bytes, then the problem seems to disappear. Id est, we are using block size of 112 right now, and we haven’t had the problem since we switched… With GNU tar the blocking factor is limited only by the maximum record size of the device containing the archive, or by the amount of available virtual memory. However, deblocking or reblocking is virtually avoided in a special case which often occurs in practice, but which requires all the following conditions to be simultaneously true:

• the archive is subject to a compression option,

• the archive is not handled through standard input or output, nor redirected nor piped,

• the archive is directly handled to a local disk, instead of any special device,

• ‘--blocking-factor’ is not explicitly specified on the tar invocation.

If the output goes directly to a local disk, and not through stdout, then the last write is not extended to a full record size. Otherwise, reblocking occurs. Here are a few other remarks on this topic:

• gzip will complain about trailing garbage if asked to uncompress a compressed archive on tape, there is an option to turn the message off, but it breaks the regularity of simply having to use ‘prog -d’ for decompression. It would be nice if gzip was silently ignoring any number of trailing zeros. I’ll ask Jean-loup Gailly, by sending a copy of this message to him.

• compress does not show this problem, but as Jean-loup pointed out to Michael, ‘compress -d’ silently adds garbage after the result of decompression, which tar ignores because it already recognized its end-of-file indicator. So this bug may be safely ignored.

• ‘gzip -d -q’ will be silent about the trailing zeros indeed, but will still return an exit status of 2 which tar reports in turn. tar might ignore the exit status returned, but I hate doing that, as it weakens the protection tar offers users against other possible problems at decompression time. If gzip was silently skipping trailing zeros and also avoiding setting the exit status in this innocuous case, that would solve this situation.

• tar should become more solid at not stopping to read a pipe at the first null block encountered. This inelegantly breaks the pipe. tar should rather drain the pipe out before exiting itself.

‘-i’
‘--ignore-zeros’

Ignore blocks of zeros in archive (means EOF). The ‘--ignore-zeros’ (‘-i’) option causes tar to ignore blocks of zeros in the archive. Normally a block of zeros indicates the end of the archive, but when reading a damaged archive, or one which was created by concatenating several archives together, this option allows tar to read the entire archive. This option is not on by default because many versions of tar write garbage after the zeroed blocks. Note that this option causes tar to read to the end of the archive file, which may sometimes avoid problems when multiple files are stored on a single physical tape.

‘-B’
‘--read-full-records’

Reblock as we read (for reading 4.2BSD pipes). If ‘--read-full-records’ is used, tar will not panic if an attempt to read a record from the archive does not return a full record. Instead, tar will keep reading until it has obtained a full record. This option is turned on by default when tar is reading an archive from standard input, or from a remote machine. This is because on BSD Unix systems, a read of a pipe will return however much happens to be in the pipe, even if it is less than tar requested. If this option was not used, tar would fail as soon as it read an incomplete record from the pipe. This option is also useful with the commands for updating an archive.

Tape blocking

When handling various tapes or cartridges, you have to take care of selecting a proper blocking, that is, the number of disk blocks you put together as a single tape block on the tape, without intervening tape gaps. A tape gap is a small landing area on the tape with no information on it, used for decelerating the tape to a full stop, and for later regaining the reading or writing speed. When the tape driver starts reading a record, the record has to be read whole without stopping, as a tape gap is needed to stop the tape motion without loosing information.

Using higher blocking (putting more disk blocks per tape block) will use the tape more efficiently as there will be less tape gaps. But reading such tapes may be more difficult for the system, as more memory will be required to receive at once the whole record. Further, if there is a reading error on a huge record, this is less likely that the system will succeed in recovering the information. So, blocking should not be too low, nor it should be too high. tar uses by default a blocking of 20 for historical reasons, and it does not really matter when reading or writing to disk. Current tape technology would easily accommodate higher blockings. Sun recommends a blocking of 126 for Exabytes and 96 for DATs. We were told that for some DLT drives, the blocking should be a multiple of 4Kb, preferably 64Kb (-b 128) or 256 for decent performance. Other manufacturers may use different recommendations for the same tapes. This might also depends of the buffering techniques used inside modern tape controllers. Some imposes a minimum blocking, or a maximum blocking. Others request blocking to be some exponent of two.

So, there is no fixed rule for blocking. But blocking at read time should ideally be the same as blocking used at write time. At one place I know, with a wide variety of equipment, they found it best to use a blocking of 32 to guarantee that their tapes are fully interchangeable.

I was also told that, for recycled tapes, prior erasure (by the same drive unit that will be used to create the archives) sometimes lowers the error rates observed at rewriting time.

   I might also use ‘--number-blocks’ instead of ‘--block-number’, so

‘--block’ will then expand to ‘--blocking-factor’ unambiguously.

9.5 Many Archives on One Tape

Most tape devices have two entries in the ‘/dev’ directory, or entries that come in pairs, which differ only in the minor number for this device. Let’s take for example ‘/dev/tape’, which often points to the only or usual tape device of a given system. There might be a corresponding ‘/dev/nrtape’ or ‘/dev/ntape’. The simpler name is the rewinding version of the device, while the name having ‘nr’ in it is the no rewinding version of the same device.

A rewinding tape device will bring back the tape to its beginning point automatically when this device is opened or closed. Since tar opens the archive file before using it and closes it afterwards, this means that a simple:

     $ tar cf /dev/tape directory

will reposition the tape to its beginning both prior and after saving directory contents to it, thus erasing prior tape contents and making it so that any subsequent write operation will destroy what has just been saved.

So, a rewinding device is normally meant to hold one and only one file. If you want to put more than one tar archive on a given tape, you will need to avoid using the rewinding version of the tape device. You will also have to pay special attention to tape positioning. Errors in positioning may overwrite the valuable data already on your tape. Many people, burnt by past experiences, will only use rewinding devices and limit themselves to one file per tape, precisely to avoid the risk of such errors. Be fully aware that writing at the wrong position on a tape loses all information past this point and most probably until the end of the tape, and this destroyed information cannot be recovered.

To save directory-1 as a first archive at the beginning of a tape, and leave that tape ready for a second archive, you should use:

     $ mt -f /dev/nrtape rewind
     $ tar cf /dev/nrtape directory-1

Tape marks are special magnetic patterns written on the tape media, which are later recognizable by the reading hardware. These marks are used after each file, when there are many on a single tape. An empty file (that is to say, two tape marks in a row) signal the logical end of the tape, after which no file exist. Usually, non-rewinding tape device drivers will react to the close request issued by tar by first writing two tape marks after your archive, and by backspacing over one of these. So, if you remove the tape at that time from the tape drive, it is properly terminated. But if you write another file at the current position, the second tape mark will be erased by the new information, leaving only one tape mark between files.

So, you may now save directory-2 as a second archive after the first on the same tape by issuing the command:

     $ tar cf /dev/nrtape directory-2

and so on for all the archives you want to put on the same tape.

Another usual case is that you do not write all the archives the same day, and you need to remove and store the tape between two archive sessions. In general, you must remember how many files are already saved on your tape. Suppose your tape already has 16 files on it, and that you are ready to write the 17th. You have to take care of skipping the first 16 tape marks before saving directory-17, say, by using these commands:

     $ mt -f /dev/nrtape rewind
     $ mt -f /dev/nrtape fsf 16
     $ tar cf /dev/nrtape directory-17

In all the previous examples, we put aside blocking considerations, but you should do the proper things for that as well. See Section 9.4 [Blocking], page 141.

9.5.1 Tape Positions and Tape Marks

(This message will disappear, once this node revised.)

Just as archives can store more than one file from the file system, tapes can store more than one archive file. To keep track of where archive files (or any other type of file stored on tape) begin and end, tape archive devices write magnetic tape marks on the archive media. Tape drives write one tape mark between files, two at the end of all the file entries.

If you think of data as a series of records "rrrr"’s, and tape marks as "*"’s, a tape might look like the following:

     rrrr*rrrrrr*rrrrr*rr*rrrrr**-------------------------

Tape devices read and write tapes using a read/write tape head—a physical part of the device which can only access one point on the tape at a time. When you use tar to read or write archive data from a tape device, the device will begin reading or writing from wherever on the tape the tape head happens to be, regardless of which archive or what part of the archive the tape head is on. Before writing an archive, you should make sure that no data on the tape will be overwritten (unless it is no longer needed). Before reading an archive, you should make sure the tape head is at the beginning of the archive you want to read. You can do it manually via mt utility (see Section 9.5.2 [mt], page 149). The restore script does that automatically (see Section 5.6 [Scripted Restoration], page 83).

If you want to add new archive file entries to a tape, you should advance the tape to the end of the existing file entries, backspace over the last tape mark, and write the new archive file. If you were to add two archives to the example above, the tape might look like the following:

     rrrr*rrrrrr*rrrrr*rr*rrrrr*rrr*rrrr**----------------

9.5.2 The mt Utility

(This message will disappear, once this node revised.)

See Section 9.4.2 [Blocking Factor], page 143.

You can use the mt utility to advance or rewind a tape past a specified number of archive files on the tape. This will allow you to move to the beginning of an archive before extracting or reading it, or to the end of all the archives before writing a new one.

The syntax of the mt command is:

     mt [-f tapename] operation [number]

where tapename is the name of the tape device, number is the number of times an operation is performed (with a default of one), and operation is one of the following:

‘eof’

‘weof’ Writes number tape marks at the current position on the tape.

‘fsf’ Moves tape position forward number files.

‘bsf’ Moves tape position back number files.

‘rewind’ Rewinds the tape. (Ignores number).

‘offline’

‘rewoff1’ Rewinds the tape and takes the tape device off-line. (Ignores number).

‘status’ Prints status information about the tape unit.

If you don’t specify a tapename, mt uses the environment variable TAPE; if TAPE is not set, mt will use the default device specified in your ‘sys/mtio.h’ file (DEFTAPE variable). If this is not defined, the program will display a descriptive error message and exit with code 1.

mt returns a 0 exit status when the operation(s) were successful, 1 if the command was unrecognized, and 2 if an operation failed.

9.6 Using Multiple Tapes

Often you might want to write a large archive, one larger than will fit on the actual tape you are using. In such a case, you can run multiple tar commands, but this can be inconvenient, particularly if you are using options like ‘--exclude=pattern’ or dumping entire file systems. Therefore, tar provides a special mode for creating multi-volume archives.

Multi-volume archive is a single tar archive, stored on several media volumes of fixed size. Although in this section we will often call ‘volume’ a tape, there is absolutely no requirement for multi-volume archives to be stored on tapes. Instead, they can use whatever media type the user finds convenient, they can even be located on files.

When creating a multi-volume archive, GNU tar continues to fill current volume until it runs out of space, then it switches to next volume (usually the operator is queried to replace the tape on this point), and continues working on the new volume. This operation continues until all requested files are dumped. If GNU tar detects end of media while dumping a file, such a file is archived in split form. Some very big files can even be split across several volumes.

Each volume is itself a valid GNU tar archive, so it can be read without any special options. Consequently any file member residing entirely on one volume can be extracted or otherwise operated upon without needing the other volume. Sure enough, to extract a split member you would need all volumes its parts reside on.

Multi-volume archives suffer from several limitations. In particular, they cannot be compressed.

GNU tar is able to create multi-volume archives of two formats (see Chapter 8 [Formats], page 117): ‘GNU’ and ‘POSIX’.

9.6.1 Archives Longer than One Tape or Disk

To create an archive that is larger than will fit on a single unit of the media, use the ‘--multi-volume’ (‘-M’) option in conjunction with the ‘--create’ option (see Section 2.6 [create], page 10). A multi-volume archive can be manipulated like any other archive (provided the ‘--multi-volume’ option is specified), but is stored on more than one tape or disk.

When you specify ‘--multi-volume’, tar does not report an error when it comes to the end of an archive volume (when reading), or the end of the media (when writing). Instead, it prompts you to load a new storage volume. If the archive is on a magnetic tape, you should change tapes when you see the prompt; if the archive is on a floppy disk, you should change disks; etc.

‘--multi-volume’

‘-M’ Creates a multi-volume archive, when used in conjunction with ‘--create’ (‘-c’). To perform any other operation on a multivolume archive, specify ‘--multi-volume’ in conjunction with that operation. For example:

                 $ tar --create --multi-volume --file=/dev/tape files

The method tar uses to detect end of tape is not perfect, and fails on some operating systems or on some devices. If tar cannot detect the end of the tape itself, you can use ‘--tape-length’ option to inform it about the capacity of the tape:

‘--tape-length=size’

‘-L size’ Set maximum length of a volume. The size argument should then be the usable size of the tape in units of 1024 bytes. This option selects ‘--multi-volume’ automatically. For example:

                 $ tar --create --tape-length=41943040 --file=/dev/tape files

When GNU tar comes to the end of a storage media, it asks you to change the volume. The built-in prompt for POSIX locale is^{1}:

     Prepare volume #n for `archive' and hit return:

where n is the ordinal number of the volume to be created and archive is archive file or device name.

When prompting for a new tape, tar accepts any of the following responses: ? Request tar to explain possible responses q Request tar to exit immediately.

n file-name

Request tar to write the next volume on the file file-name.

! Request tar to run a subshell. This option can be disabled by giving ‘--restrict’ command line option to tar^{2}. y Request tar to begin writing the next volume.

1 If you run GNU tar under a different locale, the translation to the locale’s language

will be used.

2 See [–restrict], page 39, for more information about this option

(You should only type ‘y’ after you have changed the tape; otherwise tar will write over the volume it just finished.)

The volume number used by tar in its tape-changing prompt can be changed; if you give the ‘--volno-file=file-of-number’ option, then fileof-number should be an non-existing file to be created, or else, a file already containing a decimal number. That number will be used as the volume number of the first volume written. When tar is finished, it will rewrite the file with the now-current volume number. (This does not change the volume number written on a tape label, as per Section 9.7 [label], page 155, it only affects the number used in the prompt.)

If you want more elaborate behavior than this, you can write a special new volume script, that will be responsible for changing the volume, and instruct tar to use it instead of its normal prompting procedure:

‘--info-script=script-name’
‘--new-volume-script=script-name’
‘-F script-name’

Specify the full name of the volume script to use. The script can be used to eject cassettes, or to broadcast messages such as

          ‘Someone please come change my tape’ when performing unat-

tended backups.

The script-name is executed without any command line arguments. It inherits tar’s shell environment. Additional data is passed to it via the following environment variables:

TAR_VERSION

GNU tar version number.

TAR_ARCHIVE

The name of the archive tar is processing.

TAR_VOLUME

Ordinal number of the volume tar is about to start.

TAR_SUBCOMMAND

Short option describing the operation tar is executing See Section 4.2.1 [Operations], page 52, for a complete list of subcommand options.

TAR_FORMAT

Format of the archive being processed. See Chapter 8 [Formats], page 117, for a complete list of archive format names.

The volume script can instruct tar to use new archive name, by writing in to file descriptor 3 (see below for an example).

If the info script fails, tar exits; otherwise, it begins writing the next volume.

If you want tar to cycle through a series of files or tape drives, there are three approaches to choose from. First of all, you can give tar multiple ‘--file’ options. In this case the specified files will be used, in sequence, as the successive volumes of the archive. Only when the first one in the sequence needs to be used again will tar prompt for a tape change (or run the info script). For example, suppose someone has two tape drives on a system named ‘/dev/tape0’ and ‘/dev/tape1’. For having GNU tar to switch to the second drive when it needs to write the second tape, and then back to the first tape, etc., just do either of:

     $ tar --create --multi-volume --file=/dev/tape0 --file=/dev/tape1 files
     $ tar cMff /dev/tape0 /dev/tape1 files

The second method is to use the ‘n’ response to the tape-change prompt. Finally, the most flexible approach is to use a volume script, that writes

new archive name to the file descriptor #3. For example, the following volume script will create a series of archive files, named ‘archive-vol’, where archive is the name of the archive being created (as given by ‘--file’ option) and vol is the ordinal number of the archive being created:

     #! /bin/sh
     echo Preparing volume $TAR_VOLUME of $TAR_ARCHIVE.
     name=`expr $TAR_ARCHIVE : '\(.*\)-.*'`
     case $TAR_SUBCOMMAND in
     -c)       ;;
     -d|-x|-t) test -r ${name:-$TAR_ARCHIVE}-$TAR_VOLUME || exit 1
       ;;
     *)        exit 1
     esac
     echo ${name:-$TAR_ARCHIVE}-$TAR_VOLUME >&3

The same script cant be used while listing, comparing or extracting from the created archive. For example:

# Create a multi-volume archive:

     $ tar -c -L1024 -f archive.tar -F new-volume .

# Extract from the created archive:

     $ tar -x -f archive.tar -F new-volume .

Notice, that the first command had to use ‘-L’ option, since otherwise GNU tar will end up writing everything to file ‘archive.tar’.

You can read each individual volume of a multi-volume archive as if it were an archive by itself. For example, to list the contents of one volume, use ‘--list’, without ‘--multi-volume’ specified. To extract an archive member from one volume (assuming it is described that volume), use ‘--extract’, again without ‘--multi-volume’.

If an archive member is split across volumes (i.e., its entry begins on one volume of the media and ends on another), you need to specify ‘--multi-volume’ to extract it successfully. In this case, you should load the volume where the archive member starts, and use ‘tar --extract --multi-volume’—tar will prompt for later volumes as it needs them. See Section 2.8.1 [extracting archives], page 17, for more information about extracting archives.

Multi-volume archives can be modified like any other archive. To add files to a multi-volume archive, you need to only mount the last volume of the archive media (and new volumes, if needed). For all other operations, you need to use the entire archive.

If a multi-volume archive was labeled using ‘--label=archive-label’ (see Section 9.7 [label], page 155) when it was created, tar will not automatically label volumes which are added later. To label subsequent volumes, specify ‘--label=archive-label’ again in conjunction with the ‘--append’,

‘--update’ or ‘--concatenate’ operation.

Notice that multi-volume support is a GNU extension and the archives created in this mode should be read only using GNU tar. If you absolutely have to process such archives using a third-party tar implementation, read Section 8.3.9.1 [Split Recovery], page 130.

9.6.2 Tape Files

(This message will disappear, once this node revised.)

To give the archive a name which will be recorded in it, use the

‘--label=volume-label’ (‘-V volume-label’) option. This will write a

special block identifying volume-label as the name of the archive to the front of the archive which will be displayed when the archive is listed with ‘--list’. If you are creating a multi-volume archive with ‘--multi-volume’ (see Section 9.6 [Using Multiple Tapes], page 150), then the volume label will have ‘Volume nnn’ appended to the name you give, where nnn is the number of the volume of the archive. (If you use the ‘--label=volume-label’) option when reading an archive, it checks to make sure the label on the tape matches the one you give. See Section 9.7 [label], page 155.

When tar writes an archive to tape, it creates a single tape file. If multiple archives are written to the same tape, one after the other, they each get written as separate tape files. When extracting, it is necessary to position the tape at the right place before running tar. To do this, use the mt command. For more information on the mt command and on the organization of tapes into a sequence of tape files, see Section 9.5.2 [mt], page 149.

People seem to often do:

     --label="some-prefix `date +some-format`"

or such, for pushing a common date in all volumes or an archive set.

9.6.3 Concatenate Volumes into a Single Archive

Sometimes it is necessary to convert existing GNU tar multi-volume archive to a single tar archive. Simply concatenating all volumes into one will not work, since each volume carries an additional information at the beginning. GNU tar is shipped with the shell script tarcat designed for this purpose.

The script takes a list of files comprising a multi-volume archive and creates the resulting archive at the standard output. For example:

     tarcat vol.1 vol.2 vol.3 | tar tf -

The script implements a simple heuristics to determine the format of the first volume file and to decide how to process the rest of the files. However, it makes no attempt to verify whether the files are given in order or even if they are valid tar archives. It uses dd and does not filter its standard error, so you will usually see lots of spurious messages.

9.7 Including a Label in the Archive

(This message will disappear, once this node revised.)

To avoid problems caused by misplaced paper labels on the archive media, you can include a label entry—an archive member which contains the name of the archive—in the archive itself. Use the ‘--label=archive-label’ (‘-V archive-label’) option in conjunction with the ‘--create’ operation to include a label entry in the archive as it is being created.

‘--label=archive-label’
‘-V archive-label’

Includes an archive-label at the beginning of the archive when the archive is being created, when used in conjunction with the ‘--create’ operation. Checks to make sure the archive label matches the one specified (when used in conjunction with any other operation.

If you create an archive using both ‘--label=archive-label’ (‘-V archive-label’) and ‘--multi-volume’ (‘-M’), each volume of the archive will have an archive label of the form ‘archive-label Volume n’, where n is 1 for the first volume, 2 for the next, and so on. See Section 9.6 [Using Multiple Tapes], page 150, for information on creating multiple volume archives.

The volume label will be displayed by ‘--list’ along with the file contents. If verbose display is requested, it will also be explicitly marked as in the example below:

     $ tar --verbose --list --file=iamanarchive
     V--------- 0 0        0 1992-03-07 12:01 iamalabel--Volume Header--
     -rw-r--r-- ringo user 40 1990-05-21 13:30 iamafilename

However, ‘--list’ option will cause listing entire contents of the archive, which may be undesirable (for example, if the archive is stored on a tape). You can request checking only the volume by specifying ‘--test-label’ option. This option reads only the first block of an archive, so it can be used with slow storage devices. For example:

     $ tar --test-label --file=iamanarchive
     iamalabel

If ‘--test-label’ is used with a single command line argument, tar compares the volume label with the argument. It exits with code 0 if the two strings match, and with code 2 otherwise. In this case no output is displayed. For example:

     $ tar --test-label --file=iamanarchive 'iamalable'
     ⇒ 0
     $ tar --test-label --file=iamanarchive 'iamalable' alabel
     ⇒ 1

If you request any operation, other than ‘--create’, along with using ‘--label’ option, tar will first check if the archive label matches the one specified and will refuse to proceed if it does not. Use this as a safety precaution to avoid accidentally overwriting existing archives. For example, if you wish to add files to ‘archive’, presumably labeled with string ‘My volume’, you will get:

     $ tar -rf archive --label 'My volume' .
     tar: Archive not labeled to match `My volume'

in case its label does not match. This will work even if ‘archive’ is not labeled at all.

Similarly, tar will refuse to list or extract the archive if its label doesn’t match the archive-label specified. In those cases, archive-label argument is interpreted as a globbing-style pattern which must match the actual magnetic volume label. See Section 6.4 [exclude], page 91, for a precise description of how match is attempted^{3}. If the switch ‘--multi-volume’ (‘-M’) is being used, the volume label matcher will also suffix archive-label by ‘ Volume [1-9]*’ if the initial match fails, before giving up. Since the volume numbering is automatically added in labels at creation time, it sounded logical to equally help the user taking care of it when the archive is being read.

The ‘--label’ was once called ‘--volume’, but is not available under that name anymore.

You can also use ‘--label’ to get a common information on all tapes of a series. For having this information different in each series created through a single script used on a regular basis, just manage to get some date string as part of the label. For example:

     $ tar cfMV /dev/tape "Daily backup for `date +%Y-%m-%d`"
     $ tar --create --file=/dev/tape --multi-volume \
          --volume="Daily backup for `date +%Y-%m-%d`"

Also note that each label has its own date and time, which corresponds to when GNU tar initially attempted to write it, often soon after the operator launches tar or types the carriage return telling that the next tape is ready. Comparing date labels does give an idea of tape throughput only if the delays for rewinding tapes and the operator switching them were negligible, which is usually not the case.

3 Previous versions of tar used full regular expression matching, or before that, only ex-

act string matching, instead of wildcard matchers. We decided for the sake of simplicity

to use a uniform matching device through tar.

9.8 Verifying Data as It is Stored

‘-W’
‘--verify’

Attempt to verify the archive after writing.

This option causes tar to verify the archive after writing it. Each volume is checked after it is written, and any discrepancies are recorded on the standard error output.

Verification requires that the archive be on a back-space-able medium. This means pipes, some cartridge tape drives, and some other devices cannot be verified.

You can insure the accuracy of an archive by comparing files in the system with archive members. tar can compare an archive to the file system as the archive is being written, to verify a write operation, or can compare a previously written archive, to insure that it is up to date.

To check for discrepancies in an archive immediately after it is written, use the ‘--verify’ (‘-W’) option in conjunction with the ‘--create’ operation. When this option is specified, tar checks archive members against their counterparts in the file system, and reports discrepancies on the standard error.

To verify an archive, you must be able to read it from before the end of the last written entry. This option is useful for detecting data errors on some tapes. Archives written to pipes, some cartridge tape drives, and some other devices cannot be verified.

One can explicitly compare an already made archive with the file system by using the ‘--compare’ (‘--diff’, ‘-d’) option, instead of using the more automatic ‘--verify’ option. See Section 4.2.6 [compare], page 59.

Note that these two options have a slightly different intent. The ‘--compare’ option checks how identical are the logical contents of some archive with what is on your disks, while the ‘--verify’ option is really for checking if the physical contents agree and if the recording media itself is of dependable quality. So, for the ‘--verify’ operation, tar tries to defeat all in-memory cache pertaining to the archive, while it lets the speed optimization undisturbed for the ‘--compare’ option. If you nevertheless use ‘--compare’ for media verification, you may have to defeat the in-memory cache yourself, maybe by opening and reclosing the door latch of your recording unit, forcing some doubt in your operating system about the fact this is really the same volume as the one just written or read.

The ‘--verify’ option would not be necessary if drivers were indeed able to detect dependably all write failures. This sometimes require many magnetic heads, some able to read after the writes occurred. One would not say that drivers unable to detect all cases are necessarily flawed, as long as programming is concerned.

The ‘--verify’ (‘-W’) option will not work in conjunction with the

‘--multi-volume’ (‘-M’) option or the ‘--append’ (‘-r’), ‘--update’ (‘-u’)

and ‘--delete’ operations. See Section 4.2.1 [Operations], page 52, for more information on these operations.

Also, since tar normally strips leading ‘/’ from file names (see Section 6.10.2 [absolute], page 106), a command like ‘tar --verify -cf /tmp/foo.tar /etc’ will work as desired only if the working directory is ‘/’, as tar uses the archive’s relative member names (e.g., ‘etc/motd’) when verifying the archive.

9.9 Write Protection

Almost all tapes and diskettes, and in a few rare cases, even disks can be write protected, to protect data on them from being changed. Once an archive is written, you should write protect the media to prevent the archive from being accidentally overwritten or deleted. (This will protect the archive from being changed with a tape or floppy drive—it will not protect it from magnet fields or other physical hazards).

The write protection device itself is usually an integral part of the physical media, and can be a two position (write enabled/write disabled) switch, a notch which can be popped out or covered, a ring which can be removed from the center of a tape reel, or some other changeable feature.

Appendix A Changes

This appendix lists some important user-visible changes between version GNU tar 1.16.1 and previous versions. An up-to-date version of this document is available at the GNU tar documentation page

(http://www.gnu.org/software/tar/manual/changes.html).

Use of globbing patterns when listing and extracting.

Previous versions of GNU tar assumed shell-style globbing when extracting from or listing an archive. For example:

     $ tar xf foo.tar '*.c'

would extract all files whose names end in ‘.c’. This behavior was not documented and was incompatible with traditional tar implementations. Therefore, starting from version 1.15.91, GNU tar no longer uses globbing by default. For example, the above invocation is now interpreted as a request to extract from the archive the file named ‘*.c’. To facilitate transition to the new behavior for those users who got used to the previous incorrect one, tar will print a warning if it finds out that a requested member was not found in the archive and its name looks like a globbing pattern. For example:

     $ tar xf foo.tar  '*.c'
     tar: Pattern matching characters used in file names. Please,
     tar: use --wildcards to enable pattern matching, or --no-wildcards to
     tar: suppress this warning.
     tar: *.c: Not found in archive
     tar: Error exit delayed from previous errors

To treat member names as globbing patterns, use –wildcards option. If you want to tar to mimic the behavior of versions prior to 1.15.91, add this option to your TAR_OPTIONS variable. See Section 6.5 [wildcards], page 93, for the detailed discussion of the use of globbing patterns by GNU tar.

Use of short option ‘-o’.

Earlier versions of GNU tar understood ‘-o’ command line option as a synonym for ‘--old-archive’. GNU tar starting from version 1.13.90 understands this option as a synonym for ‘--no-same-owner’. This is compatible with UNIX98 tar implementations. However, to facilitate transition, ‘-o’ option retains its old semantics when it is used with one of archive-creation commands. Users are encouraged to use ‘--format=oldgnu’ instead. It is especially important, since versions of GNU Automake up to and including 1.8.4 invoke tar with this option to produce distribution tarballs. See Chapter 8 [Formats], page 117, for the detailed discussion of this issue and its implications.

. See section “Changing Automake’s Behavior” in GNU Automake, for a description on how to use various archive formats

with automake.

Future versions of GNU tar will understand ‘-o’ only as a syn-

onym for ‘--no-same-owner’.

Use of short option ‘-l’ Earlier versions of GNU tar understood ‘-l’ option as a synonym for ‘--one-file-system’. Since such usage contradicted to UNIX98 specification and harmed compatibility with other implementation, it was declared deprecated in version 1.14. However, to facilitate transition to its new semantics, it was supported by versions 1.15 and 1.15.90. The present use of ‘-l’ as a short variant of ‘--check-links’ was introduced in version 1.15.91.

Use of options ‘--portability’ and ‘--old-archive’

These options are deprecated. Please use ‘--format=v7’ instead.

Use of option ‘--posix’ This option is deprecated. Please use ‘--format=posix’ instead.

Appendix B Configuring Help Summary

Running tar --help displays the short tar option summary (see Section 3.5 [help], page 44). This summary is organized by groups of semantically close options. The options within each group are printed in the following order: a short option, eventually followed by a list of corresponding long option names, followed by a short description of the option. For example, here is an excerpt from the actual tar --help output:

  Main operation mode:
   -A, --catenate, --concatenate   append tar files to an archive
   -c, --create               create a new archive
   -d, --diff, --compare      find differences between archive and
                              file system
       --delete               delete from the archive

The exact visual representation of the help output is configurable via ARGP_HELP_FMT environment variable. The value of this variable is a commaseparated list of format variable assignments. There are two kinds of format variables. An offset variable keeps the offset of some part of help output text from the leftmost column on the screen. A boolean variable is a flag that toggles some output feature on or off. Depending on the type of the corresponding variable, there are two kinds of assignments:

Offset assignment

The assignment to an offset variable has the following syntax:

                 variable=value

where variable is the variable name, and value is a numeric value to be assigned to the variable.

Boolean assignment To assign true value to a variable, simply put this variable name. To assign false value, prefix the variable name with ‘no-’. For example:

                 # Assign true value:
                 dup-args
                 # Assign false value:
                 no-dup-args

Following variables are declared:

[Help Output] boolean dup-args

If true, arguments for an option are shown with both short and long options, even when a given option has both forms, for example:

          -f ARCHIVE, --file=ARCHIVE use archive file or device ARCHIVE

If false, then if an option has both short and long forms, the argument is only shown with the long one, for example:

       -f, --file=ARCHIVE         use archive file or device ARCHIVE

and a message indicating that the argument is applicable to both forms is printed below the options. This message can be disabled using dupargs-note (see below). The default is false.

[Help Output] boolean dup-args-note

If this variable is true, which is the default, the following notice is displayed at the end of the help output:

Mandatory or optional arguments to long options are also mandatory or optional for any corresponding short options.

Setting no-dup-args-note inhibits this message. Normally, only one of variables dup-args or dup-args-note should be set.

[Help Output] offset short-opt-col

Column in which short options start. Default is 2.

     $ tar --help|grep ARCHIVE
       -f, --file=ARCHIVE   use archive file or device ARCHIVE
     $ ARGP_HELP_FMT=short-opt-col=6 tar --help|grep ARCHIVE
           -f, --file=ARCHIVE   use archive file or device ARCHIVE
[Help Output] offset long-opt-col

Column in which long options start. Default is 6. For example:

     $ tar --help|grep ARCHIVE
       -f, --file=ARCHIVE   use archive file or device ARCHIVE
     $ ARGP_HELP_FMT=long-opt-col=16 tar --help|grep ARCHIVE
       -f,           --file=ARCHIVE   use archive file or device ARCHIVE
[Help Output] offset doc-opt-col

Column in which doc options start. A doc option isn’t actually an option, but rather an arbitrary piece of documentation that is displayed in much the same manner as the options. For example, in the description of ‘--format’ option:

       -H, --format=FORMAT        create archive of the given format.
      FORMAT is one of the following:
         gnu                      GNU tar 1.13.x format
         oldgnu                   GNU format as per tar <= 1.12
         pax                      POSIX 1003.1-2001 (pax) format
         posix                    same as pax
         ustar                    POSIX 1003.1-1988 (ustar) format
         v7                       old V7 tar format

the format names are doc options. Thus, if you set ARGP_HELP_FMT=docopt-col=6 the above part of the help output will look as follows:

       -H, --format=FORMAT        create archive of the given format.
      FORMAT is one of the following:
             gnu                      GNU tar 1.13.x format
             oldgnu                   GNU format as per tar <= 1.12
             pax                      POSIX 1003.1-2001 (pax) format
             posix                    same as pax
             ustar                    POSIX 1003.1-1988 (ustar) format
             v7                       old V7 tar format
[Help Output] offset opt-doc-col

Column in which option description starts. Default is 29.

     $ tar --help|grep ARCHIVE
       -f, --file=ARCHIVE         use archive file or device ARCHIVE
     $ ARGP_HELP_FMT=opt-doc-col=19 tar --help|grep ARCHIVE
       -f, --file=ARCHIVE   use archive file or device ARCHIVE
     $ ARGP_HELP_FMT=opt-doc-col=9 tar --help|grep ARCHIVE
       -f, --file=ARCHIVE
                use archive file or device ARCHIVE

Notice, that the description starts on a separate line if opt-doc-col value is too small.

[Help Output] offset header-col

Column in which group headers are printed. A group header is a descriptive text preceding an option group. For example, in the following text:

  Main operation mode:
   -A, --catenate, --concatenate   append tar files to
                              an archive
   -c, --create               create a new archive

‘Main operation mode:’ is the group header. The default value is 1.

[Help Output] offset usage-indent

Indentation of wrapped usage lines. Affects ‘--usage’ output. Default is 12.

[Help Output] offset rmargin

Right margin of the text output. Used for wrapping.

164 GNU tar 1.16.1

Appendix C Tar Internals

Basic Tar Format

(This message will disappear, once this node revised.)

While an archive may contain many files, the archive itself is a single ordinary file. Like any other file, an archive file can be written to a storage device such as a tape or disk, sent through a pipe or over a network, saved on the active file system, or even stored in another archive. An archive file is not easy to read or manipulate without using the tar utility or Tar mode in GNU Emacs.

Physically, an archive consists of a series of file entries terminated by an end-of-archive entry, which consists of two 512 blocks of zero bytes. A file entry usually describes one of the files in the archive (an archive member), and consists of a file header and the contents of the file. File headers contain file names and statistics, checksum information which tar uses to detect file corruption, and information about file types.

Archives are permitted to have more than one member with the same member name. One way this situation can occur is if more than one version of a file has been stored in the archive. For information about adding new versions of a file to an archive, see Section 4.2.3 [update], page 56.

In addition to entries describing archive members, an archive may contain entries which tar itself uses to store information. See Section 9.7 [label], page 155, for an example of such an archive entry.

A tar archive file contains a series of blocks. Each block contains BLOCKSIZE bytes. Although this format may be thought of as being on magnetic tape, other media are often used.

Each file archived is represented by a header block which describes the file, followed by zero or more blocks which give the contents of the file. At the end of the archive file there are two 512-byte blocks filled with binary zeros as an end-of-file marker. A reasonable system should write such endof-file marker at the end of an archive, but must not assume that such a block exists when reading an archive. In particular GNU tar always issues a warning if it does not encounter it.

The blocks may be blocked for physical I/O operations. Each record of

n blocks (where n is set by the ‘--blocking-factor=512-size’ (‘-b 512-

size’) option to tar) is written with a single ‘write ()’ operation. On magnetic tapes, the result of such a write is a single record. When writing an archive, the last record of blocks should be written at the full size, with blocks after the zero block containing all zeros. When reading an archive, a reasonable system should properly handle an archive whose last record is shorter than the rest, or which contains garbage records after a zero block.

The header block is defined in C as follows. In the GNU tar distribution, this is part of file ‘src/tar.h’:

/* tar Header Block, from POSIX 1003.1-1990. */

/* POSIX header. */

struct posix_header

{ /* byte offset */

  char name[100];               /* 0 */
  char mode[8];                 /* 100 */
  char uid[8];                  /* 108 */
  char gid[8];                  /* 116 */
  char size[12];                /* 124 */
  char mtime[12];               /* 136 */
  char chksum[8];               /* 148 */
  char typeflag;                /* 156 */
  char linkname[100];           /* 157 */
  char magic[6];                /* 257 */
  char version[2];              /* 263 */
  char uname[32];               /* 265 */
  char gname[32];               /* 297 */
  char devmajor[8];             /* 329 */
  char devminor[8];             /* 337 */
  char prefix[155];             /* 345 */
                                /* 500 */
};
#define TMAGIC   "ustar"        /* ustar and a null */
#define TMAGLEN  6
#define TVERSION "00"           /* 00 and no null */
#define TVERSLEN 2

/* Values used in typeflag field. */

#define REGTYPE  '0'            /* regular file */
#define AREGTYPE '\0'           /* regular file */
#define LNKTYPE  '1'            /* link */
#define SYMTYPE  '2'            /* reserved */
#define CHRTYPE  '3'            /* character special */
#define BLKTYPE  '4'            /* block special */
#define DIRTYPE  '5'            /* directory */
#define FIFOTYPE '6'            /* FIFO special */
#define CONTTYPE '7'            /* reserved */

#define XHDTYPE 'x' /* Extended header referring to the

next file in the archive */

#define XGLTYPE  'g'            /* Global extended header */

/* Bits used in the mode field, values in octal. */

#define TSUID    04000          /* set UID on execution */
#define TSGID    02000          /* set GID on execution */
#define TSVTX    01000          /* reserved */

/* file permissions */

#define TUREAD   00400          /* read by owner */

Appendix C: Tar Internals 167

#define TUWRITE  00200          /* write by owner */
#define TUEXEC   00100          /* execute/search by owner */
#define TGREAD   00040          /* read by group */
#define TGWRITE  00020          /* write by group */
#define TGEXEC   00010          /* execute/search by group */
#define TOREAD   00004          /* read by other */
#define TOWRITE  00002          /* write by other */
#define TOEXEC   00001          /* execute/search by other */

/* tar Header Block, GNU extensions. */

/* In GNU tar, SYMTYPE is for to symbolic links, and CONTTYPE is for contiguous files, so maybe disobeying the ‘reserved’ comment in POSIX header description. I suspect these were meant to be used this way, and should not have really been ‘reserved’ in the published standards. */

/* *BEWARE* *BEWARE* *BEWARE* that the following information is still boiling, and may change. Even if the OLDGNU format description should be accurate, the so-called GNU format is not yet fully decided. It is surely meant to use only extensions allowed by POSIX, but the sketch below repeats some ugliness from the OLDGNU format, which should rather go away. Sparse files should be saved in such a way that they do *not* require two passes at archive creation time. Huge files get some POSIX fields to overflow, alternate solutions have to be sought for this. */

/* Descriptor for a single file hole. */

struct sparse

{ /* byte offset */

  char offset[12];              /* 0 */
  char numbytes[12];            /* 12 */
                                /* 24 */
};

/* Sparse files are not supported in POSIX ustar format. For sparse files with a POSIX header, a GNU extra header is provided which holds overall sparse information and a few sparse descriptors. When an old GNU header replaces both the POSIX header and the GNU extra header, it holds some sparse descriptors too. Whether POSIX or not, if more sparse descriptors are still needed, they are put into as many successive sparse headers as necessary. The following constants tell how many sparse descriptors fit in each kind of header able to hold them. */

#define SPARSES_IN_EXTRA_HEADER  16
#define SPARSES_IN_OLDGNU_HEADER 4
#define SPARSES_IN_SPARSE_HEADER 21

/* Extension header for sparse files, used immediately after the GNU extra header, and used only if all sparse information cannot fit into that extra header. There might even be many such extension headers, one after the other, until all sparse information has been recorded. */

struct sparse_header

{ /* byte offset */

struct sparse sp[SPARSES_IN_SPARSE_HEADER];
                              /* 0 */
char isextended;              /* 504 */
                              /* 505 */
};

/* The old GNU format header conflicts with POSIX format in such a way that POSIX archives may fool old GNU tar’s, and POSIX tar’s might well be fooled by old GNU tar archives. An old GNU format header uses the space used by the prefix field in a POSIX header, and cumulates information normally found in a GNU extra header. With an old GNU tar header, we never see any POSIX header nor GNU extra header. Supplementary sparse headers are allowed, however. */

struct oldgnu_header

{ /* byte offset */

char unused_pad1[345];        /* 0 */

char atime[12]; /* 345 Incr. archive: atime of the file */ char ctime[12]; /* 357 Incr. archive: ctime of the file */ char offset[12]; /* 369 Multivolume archive: the offset of

the start of this volume */

char longnames[4];            /* 381 Not used */
char unused_pad2;             /* 385 */
struct sparse sp[SPARSES_IN_OLDGNU_HEADER];
                              /* 386 */

char isextended; /* 482 Sparse file: Extension sparse header

follows */

char realsize[12];            /* 483 Sparse file: Real size*/
                              /* 495 */
};

/* OLDGNU MAGIC uses both magic and version fields, which are contiguous. Found in an archive, it indicates an old GNU header format, which will be hopefully become obsolescent. With OLDGNU MAGIC, uname and gname are valid, though the header is not truly POSIX conforming. */

#define OLDGNU_MAGIC "ustar  "  /* 7 chars and a null */

/* The standards committee allows only capital A through capital Z for user-defined expansion. Other letters in use include:

’A’ Solaris Access Control List ’E’ Solaris Extended Attribute File ’I’ Inode only, as in ’star’ ’N’ Obsolete GNU tar, for file names that do not fit into the main header. ’X’ POSIX 1003.1-2001 eXtended (VU version) */

/* This is a dir entry that contains the names of files that were in the

dir at the time the dump was made. */

#define GNUTYPE_DUMPDIR 'D'

Appendix C: Tar Internals 169

/* Identifies the *next* file on the tape as having a long linkname. */

#define GNUTYPE_LONGLINK 'K'

/* Identifies the *next* file on the tape as having a long name. */

#define GNUTYPE_LONGNAME 'L'

/* This is the continuation of a file that began on another volume. */

#define GNUTYPE_MULTIVOL 'M'

/* This is for sparse files. */

#define GNUTYPE_SPARSE 'S'

/* This file is a tape/volume header. Ignore it on extraction. */

#define GNUTYPE_VOLHDR 'V'

/* Solaris extended header */

#define SOLARIS_XHDTYPE 'X'

/* Jörg Schilling star header */

struct star_header

{ /* byte offset */

char name[100];               /* 0 */
char mode[8];                 /* 100 */
char uid[8];                  /* 108 */
char gid[8];                  /* 116 */
char size[12];                /* 124 */
char mtime[12];               /* 136 */
char chksum[8];               /* 148 */
char typeflag;                /* 156 */
char linkname[100];           /* 157 */
char magic[6];                /* 257 */
char version[2];              /* 263 */
char uname[32];               /* 265 */
char gname[32];               /* 297 */
char devmajor[8];             /* 329 */
char devminor[8];             /* 337 */
char prefix[131];             /* 345 */
char atime[12];               /* 476 */
char ctime[12];               /* 488 */
                              /* 500 */
};
#define SPARSES_IN_STAR_HEADER      4
#define SPARSES_IN_STAR_EXT_HEADER  21
struct star_in_header
{

char fill[345]; /* 0 Everything that is before t prefix */

char prefix[1];       /* 345 t name prefix */
char fill2;           /* 346 */
char fill3[8];        /* 347 */

       char isextended;      /* 355 */
       struct sparse sp[SPARSES_IN_STAR_HEADER]; /* 356 */
       char realsize[12];    /* 452 Actual size of the file */

char offset[12]; /* 464 Offset of multivolume contents */

       char atime[12];       /* 476 */
       char ctime[12];       /* 488 */
       char mfill[8];        /* 500 */
       char xmagic[4];       /* 508 "tar" */
     };
     struct star_ext_header
     {
       struct sparse sp[SPARSES_IN_STAR_EXT_HEADER];
       char isextended;
     };

All characters in header blocks are represented by using 8-bit characters in the local variant of ASCII. Each field within the structure is contiguous; that is, there is no padding used within the structure. Each character on the archive medium is stored contiguously.

Bytes representing the contents of files (after the header block of each file) are not translated in any way and are not constrained to represent characters in any character set. The tar format does not distinguish text files from binary files, and no translation of file contents is performed.

The name, linkname, magic, uname, and gname are null-terminated character strings. All other fields are zero-filled octal numbers in ASCII. Each numeric field of width w contains w minus 1 digits, and a null.

The name field is the file name of the file, with directory names (if any) preceding the file name, separated by slashes.

The mode field provides nine bits specifying file permissions and three bits to specify the Set UID, Set GID, and Save Text (sticky) modes. Values for these bits are defined above. When special permissions are required to create a file with a given mode, and the user restoring files from the archive does not hold such permissions, the mode bit(s) specifying those special permissions are ignored. Modes which are not supported by the operating system restoring files from the archive will be ignored. Unsupported modes should be faked up when creating or updating an archive; e.g., the group permission could be copied from the other permission.

The uid and gid fields are the numeric user and group ID of the file owners, respectively. If the operating system does not support numeric user or group IDs, these fields should be ignored.

The size field is the size of the file in bytes; linked files are archived with this field specified as zero.

The mtime field is the data modification time of the file at the time it was archived. It is the ASCII representation of the octal value of the last time the file’s contents were modified, represented as an integer number of seconds since January 1, 1970, 00:00 Coordinated Universal Time.

The chksum field is the ASCII representation of the octal value of the simple sum of all bytes in the header block. Each 8-bit byte in the header is added to an unsigned integer, initialized to zero, the precision of which shall be no less than seventeen bits. When calculating the checksum, the chksum field is treated as if it were all blanks.

The typeflag field specifies the type of file archived. If a particular implementation does not recognize or permit the specified type, the file will be extracted as if it were a regular file. As this action occurs, tar issues a warning to the standard error.

The atime and ctime fields are used in making incremental backups; they store, respectively, the particular file’s access and status change times.

The offset is used by the ‘--multi-volume’ (‘-M’) option, when making a multi-volume archive. The offset is number of bytes into the file that we need to restart at to continue the file on the next tape, i.e., where we store the location that a continued file is continued at.

The following fields were added to deal with sparse files. A file is sparse if it takes in unallocated blocks which end up being represented as zeros, i.e., no useful data. A test to see if a file is sparse is to look at the number blocks allocated for it versus the number of characters in the file; if there are fewer blocks allocated for the file than would normally be allocated for a file of that size, then the file is sparse. This is the method tar uses to detect a sparse file, and once such a file is detected, it is treated differently from non-sparse files.

Sparse files are often dbm files, or other database-type files which have data at some points and emptiness in the greater part of the file. Such files can appear to be very large when an ‘ls -l’ is done on them, when in truth, there may be a very small amount of important data contained in the file. It is thus undesirable to have tar think that it must back up this entire file, as great quantities of room are wasted on empty blocks, which can lead to running out of room on a tape far earlier than is necessary. Thus, sparse files are dealt with so that these empty blocks are not written to the tape. Instead, what is written to the tape is a description, of sorts, of the sparse file: where the holes are, how big the holes are, and how much data is found at the end of the hole. This way, the file takes up potentially far less room on the tape, and when the file is extracted later on, it will look exactly the way it looked beforehand. The following is a description of the fields used to handle a sparse file:

The sp is an array of struct sparse. Each struct sparse contains two 12-character strings which represent an offset into the file and a number of bytes to be written at that offset. The offset is absolute, and not relative to the offset in preceding array element.

The header can hold four of these struct sparse at the moment; if more are needed, they are not stored in the header.

The isextended flag is set when an extended_header is needed to deal with a file. Note that this means that this flag can only be set when dealing with a sparse file, and it is only set in the event that the description of the file will not fit in the allotted room for sparse structures in the header. In other words, an extended header is needed.

The extended_header structure is used for sparse files which need more sparse structures than can fit in the header. The header can fit 4 such structures; if more are needed, the flag isextended gets set and the next

block is an extended_header.

Each extended_header structure contains an array of 21 sparse structures, along with a similar isextended flag that the header had. There can be an indeterminate number of such extended_headers to describe a sparse file.

REGTYPE

AREGTYPE These flags represent a regular file. In order to be compatible with older versions of tar, a typeflag value of AREGTYPE should be silently recognized as a regular file. New archives should be created using REGTYPE. Also, for backward compatibility, tar treats a regular file whose name ends with a slash as a directory.

LNKTYPE This flag represents a file linked to another file, of any type, previously archived. Such files are identified in Unix by each file having the same device and inode number. The linked-to name is specified in the linkname field with a trailing null.

SYMTYPE This represents a symbolic link to another file. The linked-to name is specified in the linkname field with a trailing null.

CHRTYPE

BLKTYPE These represent character special files and block special files respectively. In this case the devmajor and devminor fields will contain the major and minor device numbers respectively. Operating systems may map the device specifications to their own local specification, or may ignore the entry.

DIRTYPE This flag specifies a directory or sub-directory. The directory name in the name field should end with a slash. On systems where disk allocation is performed on a directory basis, the size field will contain the maximum number of bytes (which may be rounded to the nearest disk block allocation unit) which the directory may hold. A size field of zero indicates no such limiting. Systems which do not support limiting in this manner should ignore the size field.

FIFOTYPE This specifies a FIFO special file. Note that the archiving of a FIFO file archives the existence of this file and not its contents.

CONTTYPE This specifies a contiguous file, which is the same as a normal file except that, in operating systems which support it, all its space is allocated contiguously on the disk. Operating systems which do not allow contiguous allocation should silently treat this type as a normal file.

A … Z These are reserved for custom implementations. Some of these are used in the GNU modified format, as described below.

Other values are reserved for specification in future revisions of the P1003 standard, and should not be used by any tar program.

The magic field indicates that this archive was output in the P1003 archive format. If this field contains TMAGIC, the uname and gname fields will contain the ASCII representation of the owner and group of the file respectively. If found, the user and group IDs are used rather than the values in the uid and gid fields.

For references, see ISO/IEC 9945-1:1990 or IEEE Std 1003.1-1990, pages 169-173 (section 10.1) for Archive/Interchange File Format; and IEEE Std 1003.2-1992, pages 380-388 (section 4.48) and pages 936-940 (section E.4.48) for pax - Portable archive interchange.

GNU Extensions to the Archive Format

(This message will disappear, once this node revised.) The GNU format uses additional file types to describe new types of files

in an archive. These are listed below.

GNUTYPE_DUMPDIR

'D' This represents a directory and a list of files created by the ‘--incremental’ (‘-G’) option. The size field gives the total size of the associated list of files. Each file name is preceded by either a ‘Y’ (the file should be in this archive) or an ‘N’. (The file is a directory, or is not stored in the archive.) Each file name is terminated by a null. There is an additional null after the last file name.

GNUTYPE_MULTIVOL

'M' This represents a file continued from another volume of a multivolume archive created with the ‘--multi-volume’ (‘-M’) option. The original type of the file is not given here. The size field gives the maximum size of this piece of the file (assuming the volume does not end before the file is written out). The offset field gives the offset from the beginning of the file where this part of the file begins. Thus size plus offset should equal the original size of the file.

GNUTYPE_SPARSE

'S' This flag indicates that we are dealing with a sparse file. Note that archiving a sparse file requires special operations to find holes in the file, which mark the positions of these holes, along with the number of bytes of data to be found after the hole.

GNUTYPE_VOLHDR

'V' This file type is used to mark the volume header that was given

          with the ‘--label=archive-label’ (‘-V archive-label’) op-

tion when the archive was created. The name field contains the

          name given after the ‘--label=archive-label’ (‘-V archive-

label’) option. The size field is zero. Only the first file in each volume of an archive should have this type.

You may have trouble reading a GNU format archive on a non-GNU system if the options ‘--incremental’ (‘-G’), ‘--multi-volume’ (‘-M’),

‘--sparse’ (‘-S’), or ‘--label=archive-label’ (‘-V archive-label’) were

used when writing the archive. In general, if tar does not use the GNU-added fields of the header, other versions of tar should be able to read the archive. Otherwise, the tar program will give an error, the most likely one being a checksum error.

Storing Sparse Files

The notion of sparse file, and the ways of handling it from the point of view of GNU tar user have been described in detail in Section 8.1.2 [sparse], page 121. This chapter describes the internal format GNU tar uses to store such files.

The support for sparse files in GNU tar has a long history. The earliest version featuring this support that I was able to find was 1.09, released in November, 1990. The format introduced back then is called old GNU sparse format and in spite of the fact that its design contained many flaws, it was the only format GNU tar supported until version 1.14 (May, 2004), which introduced initial support for sparse archives in PAX archives (see Section 8.3.6 [posix], page 126). This format was not free from design flows, either and it was subsequently improved in versions 1.15.2 (November, 2005) and 1.15.92 (June, 2006).

In addition to GNU sparse format, GNU tar is able to read and extract sparse files archived by star.

The following subsections describe each format in detail.

Old GNU Format

The format introduced some time around 1990 (v. 1.09). It was designed on top of standard ustar headers in such an unfortunate way that some of its fields overwrote fields required by POSIX.

An old GNU sparse header is designated by type ‘S’ (GNUTYPE_SPARSE) and has the following layout:

Offset Size Name  Data type  Contents

0  345  N/A  Not used.
345  12  atime  Number  atime of the file.
357  12  ctime  Number  ctime of the file .

369 12 offset Number For multivolume archives: the offset of the start of this volume.

381  4  N/A  Not used.
385  1  N/A  Not used.
386  96  sp  sparse_ header  (4 entries) File map.

482 1 isextended Bool 1 if an extension sparse header follows, 0 otherwise.

483 12 realsize Number Real size of the file.

Each of sparse_header object at offset 386 describes a single data chunk. It has the following structure:

Offset Size Data type  Contents
0  12  Number  Offset of the beginning of the chunk.
12  12  Number  Size of the chunk.

If the member contains more than four chunks, the isextended field of the header has the value 1 and the main header is followed by one or more extension headers. Each such header has the following structure:

Offset Size Name  Data type  Contents
0  21  sp  sparse_ header  (21 entires) File map.

504 1 isextended Bool 1 if an extension sparse header follows, or 0 otherwise.

A header with isextended=0 ends the map.

PAX Format, Versions 0.0 and 0.1

There are two formats available in this branch. The version 0.0 is the initial version of sparse format used by tar versions 1.14–1.15.1. The sparse file map is kept in extended (x) PAX header variables:

GNU.sparse.size

Real size of the stored file

GNU.sparse.numblocks

Number of blocks in the sparse map

GNU.sparse.offset

Offset of the data block

GNU.sparse.numbytes

Size of the data block

The latter two variables repeat for each data block, so the overall structure is like this:

     GNU.sparse.size=size
     GNU.sparse.numblocks=numblocks
     repeat numblocks times
       GNU.sparse.offset=offset
       GNU.sparse.numbytes=numbytes
     end repeat

This format presented the following two problems:

1. Whereas the POSIX specification allows a variable to appear multiple times in a header, it requires that only the last occurrence be meaningful. Thus, multiple occurrences of GNU.sparse.offset and GNU.sparse.numbytes are conflicting with the POSIX specs.

2. Attempting to extract such archives using a third-party tars results in extraction of sparse files in compressed form. If the tar implementation in question does not support POSIX format, it will also extract a file containing extension header attributes. This file can be used to expand the file to its original state. However, posix-aware tars will usually ignore the unknown variables, which makes restoring the file more difficult. See [extracting sparse v.0.x], page 133, for the detailed description of how to restore such members using non-GNU tars.

GNU tar 1.15.2 introduced sparse format version 0.1, which attempted to solve these problems. As its predecessor, this format stores sparse map in the extended POSIX header. It retains

GNU.sparse.size and GNU.sparse.numblocks variables, but instead of
GNU.sparse.offset/GNU.sparse.numbytes pairs it uses a single variable:
GNU.sparse.map

Map of non-null data chunks. It is a string consisting of commaseparated values "offset,size[,offset-1,size-1...]"

To address the 2nd problem, the name field in ustar is replaced with a special name, constructed using the following pattern:

     %d/GNUSparseFile.%p/%f

The real name of the sparse file is stored in the variable GNU.sparse.name. Thus, those tar implementations that are not aware of GNU extensions will at least extract the files into separate directories, giving the user a possibility to expand it afterwards. See [extracting sparse v.0.x], page 133, for the detailed description of how to restore such members using non-GNU tars.

The resulting GNU.sparse.map string can be very long. Although POSIX does not impose any limit on the length of a x header variable, this possibly can confuse some tars.

PAX Format, Version 1.0

The version 1.0 of sparse format was introduced with GNU tar 1.15.92. Its main objective was to make the resulting file extractable with little effort even by non-posix aware tar implementations. Starting from this version, the extended header preceding a sparse member always contains the following variables that identify the format being used:

GNU.sparse.major

Major version

GNU.sparse.minor

Minor version The name field in ustar header contains a special name, constructed using

the following pattern:

     %d/GNUSparseFile.%p/%f

The real name of the sparse file is stored in the variable GNU.sparse.name. The real size of the file is stored in the variable

GNU.sparse.realsize.

The sparse map itself is stored in the file data block, preceding the actual file data. It consists of a series of octal numbers of arbitrary length, delimited by newlines. The map is padded with nulls to the nearest block boundary.

The first number gives the number of entries in the map. Following are map entries, each one consisting of two numbers giving the offset and size of the data block it describes.

The format is designed in such a way that non-posix aware tars and tars not supporting GNU.sparse.* keywords will extract each sparse file in its condensed form with the file map prepended and will place it into a separate directory. Then, using a simple program it would be possible to expand the file to its original form even without GNU tar. See Section 8.3.9.2 [Sparse Recovery], page 131, for the detailed information on how to extract sparse members without GNU tar.

Format of the Incremental Snapshot Files

A snapshot file (or directory file) is created during incremental backups (see Section 5.2 [Incremental Dumps], page 74). It contains the status of the file system at the time of the dump and is used to determine which files were modified since the last backup.

GNU tar version 1.16.1 supports two snapshot file formats. The first format, called format 0, is the one used by GNU tar versions up to 1.15.1. The second format, called format 1 is an extended version of this format, that contains more metadata and allows for further extensions.

‘Format 0’ snapshot file begins with a line containing a decimal number that represents the UNIX timestamp of the beginning of the last archivation. This line is followed by directory metadata descriptions, one per line. Each description has the following format:

     [nfs]dev inode name

where optional nfs is a single plus character (‘+’) if this directory is located on an NFS-mounted partition, dev and inode are the device and inode numbers of the directory, and name is the directory name.

‘Format 1’ snapshot file begins with a line specifying the format of the file. This line has the following structure:

     `GNU tar-'tar-version`-'incr-format-version

where tar-version is the version of GNU tar implementation that created this snapshot, and incr-format-version is the version number of the snapshot format (in this case ‘1’).

The following line contains two decimal numbers, representing the time of the last backup. First number is the number of seconds, the second one is the number of nanoseconds, since the beginning of the epoch.

Following lines contain directory metadata, one line per directory. The line format is:

     [nfs]mtime-sec mtime-nsec dev inode name

where mtime-sec and mtime-nsec represent the last modification time of this directory with nanosecond precision; nfs, dev, inode and name have the same meaning as with ‘format 0’.

Dumpdir

Incremental archives keep information about contents of each dumped directory in special data blocks called dumpdirs.

Dumpdir is a sequence of entries of the following form:

     C filename \0

where C is one of the control codes described below, filename is the name of the file C operates upon, and ‘\0’ represents a nul character (ASCII 0). The white space characters were added for readability, real dumpdirs do not contain them.

Each dumpdir ends with a single nul character.

The following table describes control codes and their meanings:

‘Y’ filename is contained in the archive.

‘N’ filename was present in the directory at the time the archive was made, yet it was not dumped to the archive, because it had not changed since the last backup.

‘D’ filename is a directory.

‘R’ This code requests renaming of the filename to the name specified with the following ‘T’ command.

‘T’ Specify target file name for ‘R’ command (see below).

‘X’ Specify temporary directory name for a rename operation (see below).

Codes ‘Y’, ‘N’ and ‘D’ require filename argument to be a relative file name to the directory this dumpdir describes, whereas codes ‘R’, ‘T’ and ‘X’ require their argument to be an absolute file name.

The three codes ‘R’, ‘T’ and ‘X’ specify a renaming operation. In the simplest case it is:

     R`source'\0T`dest'\0

which means “rename file ‘source’ to file ‘dest’”.

However, there are cases that require using a temporary directory. For example, consider the following scenario:

1. Previous run dumped a directory ‘foo’ which contained the following

three directories:

         a
         b
         c

2. They were renamed cyclically, so that:

        `a' became `b'
        `b' became `c'
        `c' became `a'

3. New incremental dump was made.

This case cannot be handled by three successive renames, since renaming ‘a’ to ‘b’ will destroy existing directory. To handle such case a temporary directory is required. GNU tar will create the following dumpdir (newlines have been added for readability):

     Xfoo\0
     Rfoo/a\0T\0
     Rfoo/b\0Tfoo/c\0
     Rfoo/c\0Tfoo/a\0
     R\0Tfoo/a\0

The first command, ‘Xfoo\0’, instructs the extractor to create a temporary directory in the directory ‘foo’. Second command, ‘Rfoo/aT\0’, says “rename file ‘foo/a’ to the temporary directory that has just been created” (empty file name after a command means use temporary directory). Third and fourth commands work as usual, and, finally, the last command, ‘R\0Tfoo/a\0’ tells tar to rename the temporary directory to ‘foo/a’.

The exact placement of a dumpdir in the archive depends on the archive format (see Chapter 8 [Formats], page 117):

• PAX archives In PAX archives, dumpdir is stored in the extended header of the corresponding directory, in variable GNU.dumpdir.

• GNU and old GNU archives These formats implement special header type ‘D’, which is similar to ustar header ‘5’ (directory), except that it precedes a data block containing the dumpdir.

180 GNU tar 1.16.1

Appendix D Genfile

This appendix describes genfile, an auxiliary program used in the GNU tar testsuite. If you are not interested in developing GNU tar, skip this appendix.

Initially, genfile was used to generate data files for the testsuite, hence its name. However, new operation modes were being implemented as the testsuite grew more sophisticated, and now genfile is a multi-purpose instrument.

There are three basic operation modes:

File Generation This is the default mode. In this mode, genfile generates data files.

File Status In this mode genfile displays status of specified files.

Synchronous Execution.

In this mode genfile executes the given program with ‘--checkpoint’ option and executes a set of actions when specified checkpoints are reached.

D.1 Generate Mode

In this mode genfile creates a data file for the test suite. The size of the file is given with the ‘--length’ (‘-l’) option. By default the file contents is written to the standard output, this can be changed using ‘--file’ (‘-f’) command line option. Thus, the following two commands are equivalent:

     genfile --length 100 > outfile
     genfile --length 100 --file outfile

If ‘--length’ is not given, genfile will generate an empty (zero-length) file.

The command line option ‘--seek=N’ istructs genfile to skip the given number of bytes (N) in the output file before writing to it. It is similar to the ‘seek=N’ of the dd utility.

You can instruct genfile to create several files at one go, by giving it ‘--files-from’ (‘-T’) option followed by a name of file containing a list of file names. Using dash (‘-’) instead of the file name causes genfile to read file list from the standard input. For example:

     # Read file names from file `file.list'
     genfile --files-from file.list
     # Read file names from standard input
     genfile --files-from -

The list file is supposed to contain one file name per line. To use file lists separated by ASCII NUL character, use ‘--null’ (‘-0’) command line option:

     genfile --null --files-from file.list

The default data pattern for filling the generated file consists of first 256 letters of ASCII code, repeated enough times to fill the entire file. This behavior can be changed with ‘--pattern’ option. This option takes a mandatory argument, specifying pattern name to use. Currently two patterns are implemented:

‘--pattern=default’

The default pattern as described above.

‘--pattern=zero’

Fills the file with zeroes. If no file name was given, the program exits with the code 0. Otherwise,

it exits with 0 only if it was able to create a file of the specified length.

Special option ‘--sparse’ (‘-s’) instructs genfile to create a sparse file. Sparse files consist of data fragments, separated by holes or blocks of zeros. On many operating systems, actual disk storage is not allocated for holes, but they are counted in the length of the file. To create a sparse file, genfile should know where to put data fragments, and what data to use to fill them. So, when ‘--sparse’ is given the rest of the command line specifies a socalled file map.

The file map consists of any number of fragment descriptors. Each descriptor is composed of two values: a number, specifying fragment offset from the end of the previous fragment or, for the very first fragment, from the beginning of the file, and contents string, i.e., a string of characters, specifying the pattern to fill the fragment with. File offset can be suffixed with the following quantifiers:

‘k’ ‘K’ The number is expressed in kilobytes.

‘m’ ‘M’ The number is expressed in megabytes.

‘g’ ‘G’ The number is expressed in gigabytes.

For each letter in contents string genfile will generate a block of data, filled with this letter and will write it to the fragment. The size of block is given by ‘--block-size’ option. It defaults to 512. Thus, if the string consists of n characters, the resulting file fragment will contain n*blocksize of data.

Last fragment descriptor can have only file offset part. In this case genfile will create a hole at the end of the file up to the given offset.

For example, consider the following invocation:

     genfile --sparse --file sparsefile 0 ABCD 1M EFGHI 2000K

It will create 3101184-bytes long file of the following structure:

Offset  Length  Contents

0 4*512=2048 Four 512-byte blocks, filled with letters ‘A’, ‘B’, ‘C’ and ‘D’.

2048 1046528 Zero bytes

1050624 5*512=2560 Five blocks, filled with letters ‘E’, ‘F’, ‘G’, ‘H’, ‘I’.

1053184 2048000 Zero bytes

The exit code of genfile --status command is 0 only if created file is actually sparse.

D.2 Status Mode

In status mode, genfile prints file system status for each file specified in the command line. This mode is toggled by ‘--stat’ (‘-S’) command line option. An optional argument to this option specifies output format: a comma-separated list of struct stat fields to be displayed. This list can contain following identifiers :

name The file name.

dev st dev Device number in decimal.

ino st ino Inode number.

mode[.number] st mode[.number]

File mode in octal. Optional number specifies octal mask to be applied to the mode before outputting. For example, --stat mode.777 will preserve lower nine bits of it. Notice, that you can use any punctuation character in place of ‘.’.

nlink st nlink Number of hard links.

uid st uid User ID of owner.

gid st gid Group ID of owner.

size st size File size in decimal.

blksize st blksize The size in bytes of each file block.

blocks st blocks Number of blocks allocated.

atime st atime Time of last access. mtime st mtime Time of last modification

ctime st ctime Time of last status change

sparse A boolean value indicating whether the file is ‘sparse’.

Modification times are displayed in UTC as UNIX timestamps, unless suffixed with ‘H’ (for “human-readable”), as in ‘ctimeH’, in which case usual tar tv output format is used.

The default output format is: ‘name,dev,ino,mode,

nlink,uid,gid,size,blksize,blocks,atime,mtime,ctime’.

For example, the following command will display file names and corresponding times of last access for each file in the current working directory:

     genfile --stat=name,atime *

D.3 Exec Mode

This mode is designed for testing the behavior of paxutils commands when some of the files change during archiving. It is an experimental mode.

The ‘Exec Mode’ is toggled by ‘--run’ command line option (or its alias ‘-r’). The argument to this option gives the command line to be executed. The actual command line is constructed by inserting ‘--checkpoint’ option between the command name and its first argument (if any). Due to this, the argument to ‘--run’ may not use traditional tar option syntax, i.e., the following is wrong:

     # Wrong!
     genfile --run 'tar cf foo bar'

Use the following syntax instead:

     genfile --run 'tar -cf foo bar'

The rest of command line after ‘--run’ or its equivalent specifies checkpoint values and actions to be executed upon reaching them. Checkpoint values are introduced with ‘--checkpoint’ command line option. Argument to this option is the number of checkpoint in decimal.

Any number of actions may be specified after a checkpoint. Available actions are

‘--cut file’
‘--truncate file’

Truncate file to the size specified by previous ‘--length’ option (or 0, if it is not given).

‘--append file’

Append data to file. The size of data and its pattern are given by previous ‘--length’ and ‘pattern’ options.

‘--touch file’

Update the access and modification times of file. These timestamps are changed to the current time, unless ‘--date’ option was given, in which case they are changed to the specified time. Argument to ‘--date’ option is a date specification in an almost arbitrary format (see Chapter 7 [Date input formats], page 109).

‘--exec command’

Execute given shell command. Option ‘--verbose’ instructs genfile to print on standard output no-

tifications about checkpoints being executed and to verbosely describe exit status of the command.

While the command is being executed its standard output remains connected to descriptor 1. All messages it prints to file descriptor 2, except checkpoint notifications, are forwarded to standard error.

Genfile exits with the exit status of the executed command.

186 GNU tar 1.16.1

Appendix E Free Software Needs Free Documentation

The biggest deficiency in the free software community today is not in the software—it is the lack of good free documentation that we can include with the free software. Many of our most important programs do not come with free reference manuals and free introductory texts. Documentation is an essential part of any software package; when an important free software package does not come with a free manual and a free tutorial, that is a major gap. We have many such gaps today.

Consider Perl, for instance. The tutorial manuals that people normally use are non-free. How did this come about? Because the authors of those manuals published them with restrictive terms—no copying, no modification, source files not available—which exclude them from the free software world.

That wasn’t the first time this sort of thing happened, and it was far from the last. Many times we have heard a GNU user eagerly describe a manual that he is writing, his intended contribution to the community, only to learn that he had ruined everything by signing a publication contract to make it non-free.

Free documentation, like free software, is a matter of freedom, not price. The problem with the non-free manual is not that publishers charge a price for printed copies—that in itself is fine. (The Free Software Foundation sells printed copies of manuals, too.) The problem is the restrictions on the use of the manual. Free manuals are available in source code form, and give you permission to copy and modify. Non-free manuals do not allow this.

The criteria of freedom for a free manual are roughly the same as for free software. Redistribution (including the normal kinds of commercial redistribution) must be permitted, so that the manual can accompany every copy of the program, both on-line and on paper.

Permission for modification of the technical content is crucial too. When people modify the software, adding or changing features, if they are conscientious they will change the manual too—so they can provide accurate and clear documentation for the modified program. A manual that leaves you no choice but to write a new manual to document a changed version of the program is not really available to our community.

Some kinds of limits on the way modification is handled are acceptable. For example, requirements to preserve the original author’s copyright notice, the distribution terms, or the list of authors, are ok. It is also no problem to require modified versions to include notice that they were modified. Even entire sections that may not be deleted or changed are acceptable, as long as they deal with nontechnical topics (like this one). These kinds of restrictions are acceptable because they don’t obstruct the community’s normal use of the manual.

However, it must be possible to modify all the technical content of the manual, and then distribute the result in all the usual media, through all the usual channels. Otherwise, the restrictions obstruct the use of the manual, it is not free, and we need another manual to replace it.

Please spread the word about this issue. Our community continues to lose manuals to proprietary publishing. If we spread the word that free software needs free reference manuals and free tutorials, perhaps the next person who wants to contribute by writing documentation will realize, before it is too late, that only free manuals contribute to the free software community.

If you are writing documentation, please insist on publishing it under the GNU Free Documentation License or another free documentation license. Remember that this decision requires your approval—you don’t have to let the publisher decide. Some commercial publishers will use a free license if you insist, but they will not propose the option; it is up to you to raise the issue and say firmly that this is what you want. If the publisher you are dealing with refuses, please try other publishers. If you’re not sure whether a proposed license is free, write to licensing@gnu.org.

You can encourage commercial publishers to sell more free, copylefted manuals and tutorials by buying them, and particularly by buying copies from the publishers that paid for their writing or for major improvements. Meanwhile, try to avoid buying non-free documentation at all. Check the distribution terms of a manual before you buy it, and insist that whoever seeks your business must respect your freedom. Check the history of the book, and try reward the publishers that have paid or pay the authors to work on it.

The Free Software Foundation maintains a list of free documentation published by other publishers, at

http://www.fsf.org/doc/other-free-books.html.

Appendix F Copying This Manual

F.1 GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 51 Franklin St, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct

from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified

Version, as the publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adja-

cent to the other copyright notices. F. Include, immediately after the copyright notices, a license notice

giving the public permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections

and required Cover Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add

to it an item stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included in the Modified Version.

N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in title with any Invariant Section.

O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties— for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one.

The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. See

http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

F.1.1 ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

       Copyright (C)  year  your name.
       Permission is granted to copy, distribute and/or modify this document
       under the terms of the GNU Free Documentation License, Version 1.2
       or any later version published by the Free Software Foundation;
       with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
       Texts.  A copy of the license is included in the section entitled ``GNU
       Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

         with the Invariant Sections being list their titles, with
         the Front-Cover Texts being list, and with the Back-Cover Texts
         being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Appendix G: Index of Command Line Options 197

Appendix G Index of Command Line Options

This appendix contains an index of all GNU tar long command line options. The options are listed without the preceding double-dash. For a cross-reference of short command line options, Section 3.4.3 [Short Option Summary], page 43.

A

absolute-names . . . . . . . . . . . . . . . . . . . . 106 absolute-names, summary . . . . . . . . . . . 29 add-file. . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 after-date . . . . . . . . . . . . . . . . . . . . . . . . 103 after-date, summary. . . . . . . . . . . . . . . . 29 anchored. . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 anchored, summary . . . . . . . . . . . . . . . . . . 29 append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 append, summary . . . . . . . . . . . . . . . . . . . . 28 atime-preserve . . . . . . . . . . . . . . . . . . . . 122 atime-preserve, summary . . . . . . . . . . . 29

B

backup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 backup, summary . . . . . . . . . . . . . . . . . . . . 30 block-number . . . . . . . . . . . . . . . . . . . . . . . 48 block-number, summary . . . . . . . . . . . . . 30 blocking-factor . . . . . . . . . . . . . . . . . . . 143 blocking-factor, summary . . . . . . . . . . 30 bzip2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 bzip2, summary . . . . . . . . . . . . . . . . . . . . . 30

C

catenate. . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 catenate, summary . . . . . . . . . . . . . . . . . . 28 check-links, summary. . . . . . . . . . . . . . . 31 checkpoint . . . . . . . . . . . . . . . . . . . . . . . . . 47 checkpoint, summary. . . . . . . . . . . . . . . . 31 compare . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 compare, summary . . . . . . . . . . . . . . . . . . . 28 compress. . . . . . . . . . . . . . . . . . . . . . . . . . . 120 compress, summary . . . . . . . . . . . . . . . . . . 31 concatenate . . . . . . . . . . . . . . . . . . . . . . . . 57 concatenate, summary. . . . . . . . . . . . . . . 28 confirmation, summary . . . . . . . . . . . . . 31 create, additional options . . . . . . . . . . . . 59 create, complementary notes . . . . . . . . . 51 create, introduced. . . . . . . . . . . . . . . . . . . 11 create, summary . . . . . . . . . . . . . . . . . . . . 28 create, using with ‘--verbose’ . . . . . . . 12 create, using with ‘--verify’ . . . . . . . 157

D

                                        delay-directory-restore . . . . . . . . . . . 66
                                        delay-directory-restore, summary. . 31

delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 delete, summary . . . . . . . . . . . . . . . . . . . . 28 dereference . . . . . . . . . . . . . . . . . . . . . . . 125 dereference, summary. . . . . . . . . . . . . . . 31 diff, summary . . . . . . . . . . . . . . . . . . . . . . 28 directory . . . . . . . . . . . . . . . . . . . . . . . . . 105 directory, summary . . . . . . . . . . . . . . . . . 31

                                        directory, using in ‘--files-from’

argument . . . . . . . . . . . . . . . . . . . . . . . 90

E

exclude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 exclude, potential problems with . . . . . 93 exclude, summary . . . . . . . . . . . . . . . . . . . 31 exclude-caches . . . . . . . . . . . . . . . . . . . . . 92 exclude-caches, summary . . . . . . . . . . . 32 exclude-from . . . . . . . . . . . . . . . . . . . . . . . 92 exclude-from, summary . . . . . . . . . . . . . 32 exclude-tag . . . . . . . . . . . . . . . . . . . . . . . . 92 exclude-tag, summary. . . . . . . . . . . . . . . 32 extract . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 extract, additional options . . . . . . . . . . . 61 extract, complementary notes . . . . . . . . 52 extract, summary . . . . . . . . . . . . . . . . . . . 29 extract, using with

                                            ‘--listed-incremental’ . . . . . . . . . 75

F

file, short description . . . . . . . . . . . . . . . 87 file, summary . . . . . . . . . . . . . . . . . . . . . . 32 file, tutorial . . . . . . . . . . . . . . . . . . . . . . . . . 7 files-from . . . . . . . . . . . . . . . . . . . . . . . . . 90 files-from, summary. . . . . . . . . . . . . . . . 32 force-local, short description . . . . . . 138 force-local, summary. . . . . . . . . . . . . . . 32 format, summary . . . . . . . . . . . . . . . . . . . . 32

G

get, summary . . . . . . . . . . . . . . . . . . . . . . . 29 group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 group, summary . . . . . . . . . . . . . . . . . . . . . 33 gunzip, summary . . . . . . . . . . . . . . . . . . . . 33 gzip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 gzip, summary . . . . . . . . . . . . . . . . . . . . . . 33

H

help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 help, introduction . . . . . . . . . . . . . . . . . . . 45 help, summary . . . . . . . . . . . . . . . . . . . . . . 33

I

ignore-case . . . . . . . . . . . . . . . . . . . . . . . . 96 ignore-case, summary. . . . . . . . . . . . . . . 33

ignore-command-error. . . . . . . . . . . . . . . 68
ignore-command-error, summary . . . . . 33

ignore-failed-read . . . . . . . . . . . . . . . . . 61

ignore-failed-read, summary . . . . . . . 33

ignore-zeros . . . . . . . . . . . . . . . . . . . . . . . 62 ignore-zeros, short description . . . . . 146 ignore-zeros, summary . . . . . . . . . . . . . 33 incremental, summary. . . . . . . . . . . . . . . 33 incremental, using with ‘--list’ . . . . . 76 index-file, summary. . . . . . . . . . . . . . . . 33 info-script . . . . . . . . . . . . . . . . . . . . . . . 152 info-script, short description . . . . . . 139 info-script, summary. . . . . . . . . . . . . . . 34 interactive . . . . . . . . . . . . . . . . . . . . . . . . 48 interactive, summary. . . . . . . . . . . . . . . 34

K

keep-newer-files . . . . . . . . . . . . . . . . . . . 64 keep-newer-files, summary . . . . . . . . . 34 keep-old-files . . . . . . . . . . . . . . . . . . . . . 64 keep-old-files, introduced . . . . . . . . . . 62 keep-old-files, summary . . . . . . . . . . . 34

L

label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 label, summary . . . . . . . . . . . . . . . . . . . . . 34 list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 list, summary . . . . . . . . . . . . . . . . . . . . . . 29 list, using with ‘--incremental’ . . . . . 76

list, using with ‘--listed-incremental’

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 list, using with ‘--verbose’ . . . . . . . . . 15 list, using with file name arguments. . 16 listed-incremental . . . . . . . . . . . . . . . . . 74

listed-incremental, summary . . . . . . . 34
listed-incremental, using with

‘--extract’ . . . . . . . . . . . . . . . . . . . . . 75

listed-incremental, using with ‘--list’

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

M

mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 mode, summary . . . . . . . . . . . . . . . . . . . . . . 34 mtime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 mtime, summary . . . . . . . . . . . . . . . . . . . . . 34 multi-volume . . . . . . . . . . . . . . . . . . . . . . 151 multi-volume, short description . . . . . 139 multi-volume, summary . . . . . . . . . . . . . 35

N

new-volume-script . . . . . . . . . . . . . . . . . 152

new-volume-script, short description

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 new-volume-script, summary. . . . . 34, 35 newer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 newer, summary . . . . . . . . . . . . . . . . . . . . . 35 newer-mtime . . . . . . . . . . . . . . . . . . . . . . . 103 newer-mtime, summary. . . . . . . . . . . . . . . 35 no-anchored . . . . . . . . . . . . . . . . . . . . . . . . 95 no-anchored, summary. . . . . . . . . . . . . . . 35

no-delay-directory-restore . . . . . . . . 66
no-delay-directory-restore, summary

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 no-ignore-case . . . . . . . . . . . . . . . . . . . . . 96 no-ignore-case, summary . . . . . . . . . . . 35

no-ignore-command-error . . . . . . . . . . . 68
no-ignore-command-error, summary. . 36

no-overwrite-dir, summary . . . . . . . . . 36 no-quote-chars, summary . . . . . . . . . . . 36 no-recursion . . . . . . . . . . . . . . . . . . . . . . 104 no-recursion, summary . . . . . . . . . . . . . 36 no-same-owner . . . . . . . . . . . . . . . . . . . . . 123 no-same-owner, summary . . . . . . . . . . . . 36

no-same-permissions, summary . . . . . . 36

no-unquote . . . . . . . . . . . . . . . . . . . . . . . . . 89

Appendix G: Index of Command Line Options 199

no-unquote, summary. . . . . . . . . . . . . . . . 36 no-wildcards . . . . . . . . . . . . . . . . . . . . . . . 95 no-wildcards, summary . . . . . . . . . . . . . 36

no-wildcards-match-slash . . . . . . . . . . 96
no-wildcards-match-slash, summary

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 null . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 null, summary . . . . . . . . . . . . . . . . . . . . . . 36 numeric-owner . . . . . . . . . . . . . . . . . . . . . 123 numeric-owner, summary . . . . . . . . . . . . 36

O

occurrence, summary. . . . . . . . . . . . . . . . 37 old-archive, summary. . . . . . . . . . . . . . . 37 one-file-system . . . . . . . . . . . . . . . . . . . 105 one-file-system, summary . . . . . . . . . . 37 overwrite. . . . . . . . . . . . . . . . . . . . . . . . . . . 63 overwrite, introduced . . . . . . . . . . . . . . . 63 overwrite, summary . . . . . . . . . . . . . . . . . 37 overwrite-dir . . . . . . . . . . . . . . . . . . . . . . 63 overwrite-dir, introduced . . . . . . . . . . . 62 overwrite-dir, summary . . . . . . . . . . . . 37 owner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 owner, summary . . . . . . . . . . . . . . . . . . . . . 37

P

pax-option . . . . . . . . . . . . . . . . . . . . . . . . 126 pax-option, summary. . . . . . . . . . . . . . . . 38 portability, summary. . . . . . . . . . . . . . . 38 posix, summary . . . . . . . . . . . . . . . . . . . . . 38 preserve. . . . . . . . . . . . . . . . . . . . . . . . . . . 124 preserve, summary . . . . . . . . . . . . . . . . . . 38 preserve-order . . . . . . . . . . . . . . . . . . . . . 69 preserve-order, summary . . . . . . . . . . . 38

preserve-permissions. . . . . . . . . . . . . . . 65
preserve-permissions, short description

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

preserve-permissions, summary . . . . . 38

Q

quote-chars, summary. . . . . . . . . . . . . . . 38 quoting-style . . . . . . . . . . . . . . . . . . . . . . 97 quoting-style, summary . . . . . . . . . . . . 38

R

read-full-records . . . . . . . . . . . . . . 61, 62

read-full-records, short description

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 read-full-records, summary . . . . . . . . 39 record-size, summary. . . . . . . . . . . . . . . 39 recursion . . . . . . . . . . . . . . . . . . . . . . . . . 104 recursion, summary . . . . . . . . . . . . . . . . . 39 recursive-unlink . . . . . . . . . . . . . . . . . . . 64 recursive-unlink, summary . . . . . . . . . 39 remove-files . . . . . . . . . . . . . . . . . . . . . . . 68 remove-files, summary . . . . . . . . . . . . . 39 restrict, summary . . . . . . . . . . . . . . . . . . 39 rmt-command, summary. . . . . . . . . . . . . . . 39 rsh-command . . . . . . . . . . . . . . . . . . . . . . . 138 rsh-command, summary. . . . . . . . . . . . . . . 39

S

same-order . . . . . . . . . . . . . . . . . . . . . . . . . 69 same-order, summary. . . . . . . . . . . . . . . . 39 same-owner . . . . . . . . . . . . . . . . . . . . . . . . 123 same-owner, summary. . . . . . . . . . . . . . . . 39 same-permissions . . . . . . . . . . . . . . . . . . . 65 same-permissions, short description

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 same-permissions, summary . . . . . . 38, 40 seek, summary . . . . . . . . . . . . . . . . . . . . . . 35 show-defaults . . . . . . . . . . . . . . . . . . . . . . 46 show-defaults, summary . . . . . . . . . . . . 40 show-omitted-dirs . . . . . . . . . . . . . . . . . . 48 show-omitted-dirs, summary . . . . . . . . 40 show-stored-names . . . . . . . . . . . . . . . . . . 16 show-stored-names, summary . . . . . . . . 40

show-transformed-names . . . . . . . . . . . 101
show-transformed-names, summary. . . 40

sparse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 sparse, summary . . . . . . . . . . . . . . . . . . . . 40 sparse-version . . . . . . . . . . . . . . . . . . . . 122 sparse-version, summary . . . . . . . . . . . 40 starting-file . . . . . . . . . . . . . . . . . . . . . . 68 starting-file, summary . . . . . . . . . . . . 40 strip-components . . . . . . . . . . . . . . . . . . 100 strip-components, summary . . . . . . . . . 40 suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 suffix, summary . . . . . . . . . . . . . . . . . . . . 41

T

tape-length . . . . . . . . . . . . . . . . . . . . . . . 151 tape-length, short description . . . . . . 139 tape-length, summary. . . . . . . . . . . . . . . 41 test-label . . . . . . . . . . . . . . . . . . . . . . . . 155 test-label, summary. . . . . . . . . . . . . . . . 41 to-command . . . . . . . . . . . . . . . . . . . . . . . . . 67 to-command, summary. . . . . . . . . . . . . . . . 41 to-stdout. . . . . . . . . . . . . . . . . . . . . . . . . . . 66 to-stdout, summary . . . . . . . . . . . . . . . . . 41 totals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 totals, summary . . . . . . . . . . . . . . . . . . . . 41 touch . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 122 touch, summary . . . . . . . . . . . . . . . . . . . . . 41 transform . . . . . . . . . . . . . . . . . . . . . . . . . 101 transform, summary . . . . . . . . . . . . . . . . . 38

U

uncompress . . . . . . . . . . . . . . . . . . . . . . . . 120 uncompress, summary . . . . . . . . . . . . 31, 41 ungzip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ungzip, summary. . . . . . . . . . . . . . . . . 33, 41 unlink-first . . . . . . . . . . . . . . . . . . . . . . . 64 unlink-first, introduced . . . . . . . . . . . . 63 unlink-first, summary . . . . . . . . . . . . . 42 unquote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 unquote, summary . . . . . . . . . . . . . . . . . . . 42 update . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 update, summary . . . . . . . . . . . . . . . . . . . . 29 usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

                                        use-compress-program. . . . . . . . . . . . . . 120
                                        use-compress-program, summary . . . . . 42

utc, summary . . . . . . . . . . . . . . . . . . . . . . . 42

V

verbose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 verbose, introduced. . . . . . . . . . . . . . . . . . . 8 verbose, summary . . . . . . . . . . . . . . . . . . . 42 verbose, using with ‘--create’ . . . . . . . 12 verbose, using with ‘--list’ . . . . . . . . . 15

                                        verify, short description . . . . . . . . . 157

verify, summary . . . . . . . . . . . . . . . . . . . . 42 verify, using with ‘--create’ . . . . . . . 157 version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 version, summary . . . . . . . . . . . . . . . . . . . 42 volno-file . . . . . . . . . . . . . . . . . . . . . . . . 152 volno-file, summary. . . . . . . . . . . . . . . . 42

W

wildcards. . . . . . . . . . . . . . . . . . . . . . . . . . . 95 wildcards, summary . . . . . . . . . . . . . . . . . 42

                                        wildcards-match-slash. . . . . . . . . . . . . . 96
                                        wildcards-match-slash, summary . . . . 42

Appendix H: Index 201

Appendix H Index

A

abbreviations for months . . . . . . . . . . . . 111 absolute file names. . . . . . . . . . . . . . . . . . 139 Adding archives to an archive . . . . . . . . . 57 Adding files to an Archive . . . . . . . . . . . . 54 ADMINISTRATOR . . . . . . . . . . . . . . . . . . . . . . 78 Age, excluding files by . . . . . . . . . . . . . . 102 ago in date strings . . . . . . . . . . . . . . . . . . 113 am in date strings . . . . . . . . . . . . . . . . . . . 112 Appending files to an Archive. . . . . . . . . 54 archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Archive creation . . . . . . . . . . . . . . . . . . . . . 87 archive member . . . . . . . . . . . . . . . . . . . . . . . 2 Archive Name . . . . . . . . . . . . . . . . . . . . . . . 87 Archive, creation of . . . . . . . . . . . . . . . . . . 10 Archives, Appending files to . . . . . . . . . . 54 Archiving Directories. . . . . . . . . . . . . . . . . 14

ARGP_HELP_FMT, environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 authors of get_date. . . . . . . . . . . . . . . . . 115 Avoiding recursion in directories . . . . . 104

B

backup options . . . . . . . . . . . . . . . . . . . . . . 69 backup suffix . . . . . . . . . . . . . . . . . . . . . . . . 70 BACKUP_DIRS . . . . . . . . . . . . . . . . . . . . . . . . 78 BACKUP_FILES . . . . . . . . . . . . . . . . . . . . . . . 79 BACKUP_HOUR . . . . . . . . . . . . . . . . . . . . . . . . 78 backups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 beginning of time, for POSIX . . . . . . . . 114 Bellovin, Steven M. . . . . . . . . . . . . . . . . . 115 Berets, Jim. . . . . . . . . . . . . . . . . . . . . . . . . 115 Berry, K.. . . . . . . . . . . . . . . . . . . . . . . . . . . 116 Block number where error occurred … 48 BLOCKING. . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 blocking factor . . . . . . . . . . . . . . . . . . . . . 146 Blocking Factor. . . . . . . . . . . . . . . . . . . . . 143 Blocks per record . . . . . . . . . . . . . . . . . . . 143 bug reports. . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Bytes per record . . . . . . . . . . . . . . . . . . . . 143

C

calendar date item . . . . . . . . . . . . . . . . . . 110 case, ignored in dates . . . . . . . . . . . . . . . 110 cat vs concatenate . . . . . . . . . . . . . . . . . . 58 Changing directory mid-stream . . . . . . 105 Character class, excluding characters from

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 Choosing an archive file . . . . . . . . . . . . . . 87 comments, in dates . . . . . . . . . . . . . . . . . 110 Compressed archives . . . . . . . . . . . . . . . . 118 concatenate vs cat . . . . . . . . . . . . . . . . . . 58 Concatenating Archives . . . . . . . . . . . . . . 57 corrupted archives . . . . . . . . . . . . . . . 73, 120 Creation of the archive . . . . . . . . . . . . . . . 10

D

DAT blocking . . . . . . . . . . . . . . . . . . . . . . 147 Data Modification time, excluding files by

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 Data modification times of extracted files

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 date format, iso 8601 . . . . . . . . . . . . . . . 111 date input formats . . . . . . . . . . . . . . . . . . 109 day in date strings . . . . . . . . . . . . . . . . . . 113 day of week item. . . . . . . . . . . . . . . . . . . . 112 Deleting files from an archive . . . . . . . . . 58 Deleting from tape archives . . . . . . . . . . . 58 Descending directories, avoiding. . . . . . 104 Directories, Archiving . . . . . . . . . . . . . . . . 14 Directories, avoiding recursion . . . . . . . 104 Directory, changing mid-stream . . . . . . 105 DIRLIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 displacement of dates . . . . . . . . . . . . . . . 113 doc-opt-col . . . . . . . . . . . . . . . . . . . . . . . 162 Double-checking a write operation. … 157 DUMP_BEGIN . . . . . . . . . . . . . . . . . . . . . . . . . 81 DUMP_END. . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 DUMP_REMIND_SCRIPT . . . . . . . . . . . . . . . . . 79 dumps, full . . . . . . . . . . . . . . . . . . . . . . . . . . 73 dup-args. . . . . . . . . . . . . . . . . . . . . . . . . . . 161 dup-args-note . . . . . . . . . . . . . . . . . . . . . 162

E

Eggert, Paul. . . . . . . . . . . . . . . . . . . . . . . . 115 End-of-archive blocks, ignoring. . . . . . . . 62 End-of-archive info script . . . . . . . . . . . . 152 entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 epoch, for POSIX . . . . . . . . . . . . . . . . . . . 114 Error message, block number of . . . . . . . 48 Exabyte blocking . . . . . . . . . . . . . . . . . . . 147 exclude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 exclude-caches . . . . . . . . . . . . . . . . . . . . . 92 exclude-from . . . . . . . . . . . . . . . . . . . . . . . 92 exclude-tag . . . . . . . . . . . . . . . . . . . . . . . . 92 Excluding characters from a character

class . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 Excluding file by age . . . . . . . . . . . . . . . . 102 Excluding files by file system . . . . . . . . . 91 Excluding files by name and pattern… 91 Exec Mode, genfile . . . . . . . . . . . . . . . . 184 existing backup method . . . . . . . . . . . . . . 70 exit status. . . . . . . . . . . . . . . . . . . . . . . . . . . 22 extraction . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Extraction . . . . . . . . . . . . . . . . . . . . . . . . . . 17

F

FDL, GNU Free Documentation License

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 File lists separated by NUL characters

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 File Name arguments, alternatives …. 89 File name arguments, using ‘--list’ with

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 File names, excluding files by . . . . . . . . . 91 File names, terminated by NUL . . . . . . . . 91 File names, using symbolic links . . . . . 125 File system boundaries, not crossing. . 105 FILELIST. . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 first in date strings . . . . . . . . . . . . . . . . 109 Format Options . . . . . . . . . . . . . . . . . . . . 143 Format Parameters . . . . . . . . . . . . . . . . . 143 Format, old style . . . . . . . . . . . . . . . . . . . 125 fortnight in date strings . . . . . . . . . . . 113 free documentation . . . . . . . . . . . . . . . . . 187 full dumps . . . . . . . . . . . . . . . . . . . . . . . . . . 73 future time stamps. . . . . . . . . . . . . . . . . . 129

G

general date syntax . . . . . . . . . . . . . . . . . 109 Generate Mode, genfile . . . . . . . . . . . . 181 genfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 genfile, create file . . . . . . . . . . . . . . . . . 181 genfile, creating sparse files . . . . . . . . 182 genfile, generate mode . . . . . . . . . . . . . 181 genfile, reading a list of file names . . 181 genfile, seeking to a given offset …. 181 get_date. . . . . . . . . . . . . . . . . . . . . . . . . . . 109 Getting program version number . . . . . . 44 GNU archive format . . . . . . . . . . . . . . . . 126

                                        GNU.sparse.major, extended header

variable . . . . . . . . . . . . . . . . . . . . . . 177

                                        GNU.sparse.map, extended header

variable . . . . . . . . . . . . . . . . . . . . . . 176

                                        GNU.sparse.minor, extended header

variable . . . . . . . . . . . . . . . . . . . . . . 177

                                        GNU.sparse.name, extended header

variable . . . . . . . . . . . . . . . . . . . . . . 176

                                        GNU.sparse.name, extended header
                                            variable, in v.1.0 . . . . . . . . . . . . 177
                                        GNU.sparse.numblocks, extended header

variable . . . . . . . . . . . . . . . . . . . . . . 175

                                        GNU.sparse.numbytes, extended header

variable . . . . . . . . . . . . . . . . . . . . . . 175

                                        GNU.sparse.offset, extended header

variable . . . . . . . . . . . . . . . . . . . . . . 175

                                        GNU.sparse.realsize, extended header

variable . . . . . . . . . . . . . . . . . . . . . . 177

                                        GNU.sparse.size, extended header

variable . . . . . . . . . . . . . . . . . . . . . . 175 gnupg, using with tar . . . . . . . . . . . . . . . 120 gpg, using with tar. . . . . . . . . . . . . . . . . . 120

H

header-col . . . . . . . . . . . . . . . . . . . . . . . . 163 hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 hour in date strings . . . . . . . . . . . . . . . . . 113

I

Ignoring end-of-archive blocks. . . . . . . . . 62 Info script. . . . . . . . . . . . . . . . . . . . . . . . . . 152 Interactive operation . . . . . . . . . . . . . . . . . 48 iso 8601 date format . . . . . . . . . . . . . . . . 111 items in date strings . . . . . . . . . . . . . . . . 109

L

Labeling an archive . . . . . . . . . . . . . . . . . 155 Labeling multi-volume archives . . . . . . 155 Labels on the archive media . . . . . . . . . 155 language, in dates. . . . . . . . . . . . . . . . . . . 110 Large lists of file names on small machines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 large values . . . . . . . . . . . . . . . . . . . . . . . . 129 last day . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 last in date strings . . . . . . . . . . . . . . . . . 109 Listing all tar options. . . . . . . . . . . . . . . . 45 listing member and file names. . . . . . . . . 15 Listing volume label . . . . . . . . . . . . . . . . 155 Lists of file names. . . . . . . . . . . . . . . . . . . . 89 Local and remote archives . . . . . . . . . . . . 88 long-opt-col . . . . . . . . . . . . . . . . . . . . . . 162

M

MacKenzie, David . . . . . . . . . . . . . . . . . . 115 member . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 member name . . . . . . . . . . . . . . . . . . . . . . . . 2 Members, replacing with other members

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Meyering, Jim . . . . . . . . . . . . . . . . . . . . . . 115 Middle of the archive, starting in the . . 69 midnight in date strings. . . . . . . . . . . . . 112 minute in date strings . . . . . . . . . . . . . . . 113 minutes, time zone correction by . . . . . 112 Modes of extracted files . . . . . . . . . . . . . . 65 Modification time, excluding files by. . 102 Modification times of extracted files … 64 month in date strings . . . . . . . . . . . . . . . . 113 month names in date strings . . . . . . . . . 111 months, written-out . . . . . . . . . . . . . . . . . 110 MT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 MT_BEGIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 MT_OFFLINE . . . . . . . . . . . . . . . . . . . . . . . . . 80 MT_REWIND. . . . . . . . . . . . . . . . . . . . . . . . . . . 80 MT_STATUS. . . . . . . . . . . . . . . . . . . . . . . . . . . 80 Multi-volume archives . . . . . . . . . . . . . . . 151 Mutli-volume archives in PAX format,

extracting using non-GNU tars . . 130 Mutli-volume archives, extracting using

non-GNU tars . . . . . . . . . . . . . . . . . . 130

N

Naming an archive . . . . . . . . . . . . . . . . . . . 87 negative time stamps. . . . . . . . . . . . . . . . 129 next day . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 next in date strings . . . . . . . . . . . . . . . . . 109 noon in date strings . . . . . . . . . . . . . . . . . 112 now in date strings . . . . . . . . . . . . . . . . . . 113 ntape device . . . . . . . . . . . . . . . . . . . . . . . 147 NUL terminated file names . . . . . . . . . . . . 91 Number of blocks per record . . . . . . . . . 143 Number of bytes per record. . . . . . . . . . 143 numbered backup method . . . . . . . . . . . . 70 numbers, written-out. . . . . . . . . . . . . . . . 109

O

Obtaining help . . . . . . . . . . . . . . . . . . . . . . 45 Obtaining total status information …. 47 Old GNU archive format . . . . . . . . . . . . 126 Old GNU sparse format . . . . . . . . . . . . . 174 Old style archives . . . . . . . . . . . . . . . . . . . 125 Old style format . . . . . . . . . . . . . . . . . . . . 125 opt-doc-col . . . . . . . . . . . . . . . . . . . . . . . 163 option syntax, traditional. . . . . . . . . . . . . 26 Options when reading archives . . . . . . . . 61 Options, archive format specifying … 143 Options, format specifying. . . . . . . . . . . 143 ordinal numbers . . . . . . . . . . . . . . . . . . . . 109 Overwriting old files, prevention . . . . . . 62

P

pattern, genfile. . . . . . . . . . . . . . . . . . . . 182 PAX archive format. . . . . . . . . . . . . . . . . 126 Permissions of extracted files . . . . . . . . . 65 Pinard, F. . . . . . . . . . . . . . . . . . . . . . . . . . . 116 pm in date strings . . . . . . . . . . . . . . . . . . . 112 POSIX archive format . . . . . . . . . . . . . . 126 Progress information . . . . . . . . . . . . . . . . . 47 Protecting old files . . . . . . . . . . . . . . . . . . . 63 pure numbers in date strings. . . . . . . . . 114

R

Reading file names from a file. . . . . . . . . 89 Reading incomplete records . . . . . . . . . . . 61 Record Size . . . . . . . . . . . . . . . . . . . . . . . . 143 Records, incomplete. . . . . . . . . . . . . . . . . . 61 Recursion in directories, avoiding …. 104 relative items in date strings . . . . . . . . . 113 Remote devices . . . . . . . . . . . . . . . . . . . . . . 88 remote tape drive . . . . . . . . . . . . . . . . . . . 139 Removing files from an archive. . . . . . . . 58 Replacing members with other members

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . 4 RESTORE_BEGIN . . . . . . . . . . . . . . . . . . . . . . 81 RESTORE_END . . . . . . . . . . . . . . . . . . . . . . . . 81 Resurrecting files from an archive . . . . . 17 Retrieving files from an archive . . . . . . . 17 return status . . . . . . . . . . . . . . . . . . . . . . . . 22 rmargin. . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 rmt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 RSH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 RSH_COMMAND . . . . . . . . . . . . . . . . . . . . . . . . 79 Running out of space. . . . . . . . . . . . . . . . . 68

S

Salz, Rich . . . . . . . . . . . . . . . . . . . . . . . . . . 115 short-opt-col . . . . . . . . . . . . . . . . . . . . . 162 simple backup method . . . . . . . . . . . . . . . 70

SIMPLE_BACKUP_SUFFIX. . . . . . . . . . . . . . . 70

SLEEP_MESSAGE . . . . . . . . . . . . . . . . . . . . . . 79 SLEEP_TIME . . . . . . . . . . . . . . . . . . . . . . . . . 79 Small memory . . . . . . . . . . . . . . . . . . . . . . . 68 Sparse Files . . . . . . . . . . . . . . . . . . . . . . . . 121 sparse files v.0.0, extracting with

non-GNU tars . . . . . . . . . . . . . . . . . . 133 sparse files v.0.1, extracting with

non-GNU tars . . . . . . . . . . . . . . . . . . 133 sparse files v.1.0, extracting with

non-GNU tars . . . . . . . . . . . . . . . . . . 131 Sparse files, creating using genfile . . 182 sparse files, extracting with non-GNU tars

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 sparse formats . . . . . . . . . . . . . . . . . . . . . . 174 sparse formats, defined . . . . . . . . . . . . . . 122 sparse formats, Old GNU. . . . . . . . . . . . 174 sparse formats, v.0.0 . . . . . . . . . . . . . . . . 175 sparse formats, v.0.1 . . . . . . . . . . . . . . . . 176 sparse formats, v.1.0 . . . . . . . . . . . . . . . . 177 sparse versions. . . . . . . . . . . . . . . . . . . . . . 174 Specifying archive members. . . . . . . . . . . 88 Specifying files to act on. . . . . . . . . . . . . . 88 Standard input and output . . . . . . . . . . . 87 Standard output, writing extracted files to

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Storing archives in compressed format

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 Symbolic link as file name . . . . . . . . . . . 125

T

TAPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 tape blocking . . . . . . . . . . . . . . . . . . . . . . . 146 tape marks . . . . . . . . . . . . . . . . . . . . . . . . . 148 tape positioning . . . . . . . . . . . . . . . . . . . . 148 TAPE_FILE. . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Tapes, using ‘--delete’ and . . . . . . . . . . 58 tar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 TAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 tar archive . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Tar archive formats . . . . . . . . . . . . . . . . . 117 tar entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 tar file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 tar to a remote device . . . . . . . . . . . . . . . . 88 tar to standard input and output . . . . . 87

TAR_ARCHIVE, info script environment

variable . . . . . . . . . . . . . . . . . . . . . . 152

TAR_ATIME, to-command environment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

                                        TAR_CTIME, to-command environment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

                                        TAR_FILENAME, to-command environment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

                                        TAR_FILETYPE, to-command environment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

                                        TAR_FORMAT, info script environment

variable . . . . . . . . . . . . . . . . . . . . . . 152

                                        TAR_GID, to-command environment. . . . 68
                                        TAR_GNAME, to-command environment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

                                        TAR_MODE, to-command environment . . 67
                                        TAR_MTIME, to-command environment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

                                        TAR_OPTIONS, environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

                                        TAR_REALNAME, to-command environment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

                                        TAR_SIZE, to-command environment . . 68
                                        TAR_SUBCOMMAND, info script
                                            environment variable . . . . . . . . . . 152
                                        TAR_UID, to-command environment. . . . 68
                                        TAR_UNAME, to-command environment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

                                        TAR_VERSION, info script environment

variable . . . . . . . . . . . . . . . . . . . . . . 152

                                        TAR_VOLUME, info script environment

variable . . . . . . . . . . . . . . . . . . . . . . 152 tarcat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 this in date strings . . . . . . . . . . . . . . . . . 113 time of day item . . . . . . . . . . . . . . . . . . . . 111 time zone correction . . . . . . . . . . . . . . . . 112 time zone item . . . . . . . . . . . . . . . . . 110, 112 today in date strings . . . . . . . . . . . . . . . . 113 tomorrow in date strings. . . . . . . . . . . . . 113 TZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

U

Ultrix 3.1 and write failure . . . . . . . . . . 140 unpacking . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Updating an archive . . . . . . . . . . . . . . . . . 56 usage-indent . . . . . . . . . . . . . . . . . . . . . . 163 Using encrypted archives . . . . . . . . . . . . 120 ustar archive format . . . . . . . . . . . . . . . . 126 uuencode. . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

V

v7 archive format . . . . . . . . . . . . . . . . . . . 125 Verbose operation. . . . . . . . . . . . . . . . . . . . 46 Verifying a write operation . . . . . . . . . . 157 Verifying the currency of an archive … 59 Version of the tar program . . . . . . . . . . . 44 version-control Emacs variable . . . . . 70 VERSION_CONTROL . . . . . . . . . . . . . . . . . . . . 70 volno file . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 VOLNO_FILE . . . . . . . . . . . . . . . . . . . . . . . . . 79 Volume label, listing . . . . . . . . . . . . . . . . 155 Volume number file . . . . . . . . . . . . . . . . . 152

W

week in date strings . . . . . . . . . . . . . . . . . 113 Where is the archive? . . . . . . . . . . . . . . . . 87 Working directory, specifying . . . . . . . . 105 Writing extracted files to standard output

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Writing new archives . . . . . . . . . . . . . . . . . 87

X

XLIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 xsparse. . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

Y

year in date strings . . . . . . . . . . . . . . . . . 113 yesterday in date strings . . . . . . . . . . . 113

206 GNU tar 1.16.1

Short Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Tutorial Introduction to tar. . . . . . . . . . . . . . . . . . . . . . . . 5 3 Invoking GNU tar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4 GNU tar Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 5 Performing Backups and Restoring Files . . . . . . . . . . . . . . . 73 6 Choosing Files and Names for tar . . . . . . . . . . . . . . . . . . . 87 7 Date input formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 8 Controlling the Archive Format . . . . . . . . . . . . . . . . . . . . 117 9 Tapes and Other Archive Media. . . . . . . . . . . . . . . . . . . . 137 A Changes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 B Configuring Help Summary . . . . . . . . . . . . . . . . . . . . . . . 161 C Tar Internals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 D Genfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 E Free Software Needs Free Documentation . . . . . . . . . . . . . 187 F Copying This Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 G Index of Command Line Options . . . . . . . . . . . . . . . . . . . 197 H Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 ii GNU tar 1.16.1

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 What this Book Contains . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Some Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 What tar Does . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 How tar Archives are Named . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.5 GNU tar Authors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.6 Reporting bugs or suggestions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Tutorial Introduction to tar . . . . . . . . . . . . . . . . 5

2.1 Assumptions this Tutorial Makes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 Stylistic Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.3 Basic tar Operations and Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.4 The Three Most Frequently Used Operations . . . . . . . . . . . . . . . . . . 7 2.5 Two Frequently Used Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

The ‘--file’ Option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 The ‘--verbose’ Option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Getting Help: Using the ‘--help’ Option . . . . . . . . . . . . . . . . . . . . . . 10

2.6 How to Create Archives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.6.1 Preparing a Practice Directory for Examples . . . . . . . . . . . . . 11 2.6.2 Creating the Archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.6.3 Running ‘--create’ with ‘--verbose’ . . . . . . . . . . . . . . . . . . . 12 2.6.4 Short Forms with ‘create’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.6.5 Archiving Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2.7 How to List Archives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

Listing the Contents of a Stored Directory . . . . . . . . . . . . . . . . . . . . . 16 2.8 How to Extract Members from an Archive . . . . . . . . . . . . . . . . . . . 17

2.8.1 Extracting an Entire Archive . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.8.2 Extracting Specific Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.8.3 Extracting Files that are Directories. . . . . . . . . . . . . . . . . . . . . 18 2.8.4 Extracting Archives from Untrusted Sources . . . . . . . . . . . . . 19 2.8.5 Commands That Will Fail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

2.9 Going Further Ahead in this Manual. . . . . . . . . . . . . . . . . . . . . . . . . 20

3 Invoking GNU tar . . . . . . . . . . . . . . . . . . . . . . . . 21

3.1 General Synopsis of tar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.2 Using tar Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.3 The Three Option Styles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

3.3.1 Long Option Style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.3.2 Short Option Style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.3.3 Old Option Style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.3.4 Mixing Option Styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.4 All tar Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

3.4.1 Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.4.2 tar Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.4.3 Short Options Cross Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 43

3.5 GNU tar documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.6 Obtaining GNU tar default values . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.7 Checking tar progress . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.8 Asking for Confirmation During Operations . . . . . . . . . . . . . . . . . . 48

4 GNU tar Operations . . . . . . . . . . . . . . . . . . . . . . 51

4.1 Basic GNU tar Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 4.2 Advanced GNU tar Operations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

4.2.1 The Five Advanced tar Operations . . . . . . . . . . . . . . . . . . . . . 52 4.2.2 How to Add Files to Existing Archives: ‘--append’ . . . . . . . 53

4.2.2.1 Appending Files to an Archive . . . . . . . . . . . . . . . . . . . . . 54 4.2.2.2 Multiple Members with the Same Name. . . . . . . . . . . . . 55

4.2.3 Updating an Archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

4.2.3.1 How to Update an Archive Using ‘--update’ . . . . . . . . 56 4.2.4 Combining Archives with ‘--concatenate’ . . . . . . . . . . . . . . 57 4.2.5 Removing Archive Members Using ‘--delete’ . . . . . . . . . . . 58 4.2.6 Comparing Archive Members with the File System . . . . . . . 59

4.3 Options Used by ‘--create’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 4.3.1 Overriding File Metadata. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 4.3.2 Ignore Fail Read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

4.4 Options Used by ‘--extract’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

4.4.1 Options to Help Read Archives . . . . . . . . . . . . . . . . . . . . . . . . . 61 Reading Full Records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Ignoring Blocks of Zeros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

4.4.2 Changing How tar Writes Files . . . . . . . . . . . . . . . . . . . . . . . . . 62 Options Controlling the Overwriting of Existing Files . . . . . . . . 62 Overwrite Old Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Keep Old Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Keep Newer Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Unlink First . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Recursive Unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Setting Data Modification Times . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Setting Access Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 Directory Modification Times and Permissions. . . . . . . . . . . . . . . 65 Writing to Standard Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Writing to an External Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Removing Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

4.4.3 Coping with Scarce Resources . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Starting File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Same Order. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

4.5 Backup options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 4.6 Notable tar Usages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

4.7 Looking Ahead: The Rest of this Manual . . . . . . . . . . . . . . . . . . . . 71

5 Performing Backups and Restoring Files … 73

5.1 Using tar to Perform Full Dumps . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 5.2 Using tar to Perform Incremental Dumps . . . . . . . . . . . . . . . . . . . . 74 5.3 Levels of Backups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 5.4 Setting Parameters for Backups and Restoration . . . . . . . . . . . . . . 77

5.4.1 General-Purpose Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 5.4.2 Magnetic Tape Control. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 5.4.3 User Hooks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 5.4.4 An Example Text of ‘Backup-specs’ . . . . . . . . . . . . . . . . . . . . 81

5.5 Using the Backup Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 5.6 Using the Restore Script . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

6 Choosing Files and Names for tar . . . . . . . . . 87

6.1 Choosing and Naming Archive Files . . . . . . . . . . . . . . . . . . . . . . . . . 87 6.2 Selecting Archive Members. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 6.3 Reading Names from a File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

6.3.1 NUL Terminated File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 6.4 Excluding Some Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

Problems with Using the exclude Options . . . . . . . . . . . . . . . . . . . . . 93 6.5 Wildcards Patterns and Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

Controlling Pattern-Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 6.6 Quoting Member Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 6.7 Modifying File and Member Names . . . . . . . . . . . . . . . . . . . . . . . . . 100 6.8 Operating Only on New Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 6.9 Descending into Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 6.10 Crossing File System Boundaries . . . . . . . . . . . . . . . . . . . . . . . . . . 105

6.10.1 Changing the Working Directory. . . . . . . . . . . . . . . . . . . . . . 105 6.10.2 Absolute File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

7 Date input formats . . . . . . . . . . . . . . . . . . . . . . 109

7.1 General date syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 7.2 Calendar date items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 7.3 Time of day items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 7.4 Time zone items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 7.5 Day of week items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 7.6 Relative items in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 7.7 Pure numbers in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 7.8 Seconds since the Epoch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 7.9 Specifying time zone rules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 7.10 Authors of get_date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

8 Controlling the Archive Format . . . . . . . . . . 117

8.1 Using Less Space through Compression . . . . . . . . . . . . . . . . . . . . . 118 8.1.1 Creating and Reading Compressed Archives . . . . . . . . . . . . 118 8.1.2 Archiving Sparse Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

8.2 Handling File Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 8.3 Making tar Archives More Portable . . . . . . . . . . . . . . . . . . . . . . . . 124

8.3.1 Portable Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 8.3.2 Symbolic Links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 8.3.3 Old V7 Archives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 8.3.4 Ustar Archive Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 8.3.5 GNU and old GNU tar format . . . . . . . . . . . . . . . . . . . . . . . . . 126 8.3.6 GNU tar and POSIX tar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

8.3.6.1 Controlling Extended Header Keywords . . . . . . . . . . . . 126 8.3.7 Checksumming Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 8.3.8 Large or Negative Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 8.3.9 How to Extract GNU-Specific Data Using Other tar

Implementations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 8.3.9.1 Extracting Members Split Between Volumes. . . . . . . . 130 8.3.9.2 Extracting Sparse Members . . . . . . . . . . . . . . . . . . . . . . . 131

8.4 Comparison of tar and cpio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134

9 Tapes and Other Archive Media . . . . . . . . . . 137

9.1 Device Selection and Switching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 9.2 The Remote Tape Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 9.3 Some Common Problems and their Solutions . . . . . . . . . . . . . . . . 140 9.4 Blocking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

9.4.1 Format Variations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 9.4.2 The Blocking Factor of an Archive . . . . . . . . . . . . . . . . . . . . . 143

9.5 Many Archives on One Tape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 9.5.1 Tape Positions and Tape Marks. . . . . . . . . . . . . . . . . . . . . . . . 149 9.5.2 The mt Utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

9.6 Using Multiple Tapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 9.6.1 Archives Longer than One Tape or Disk . . . . . . . . . . . . . . . . 151 9.6.2 Tape Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 9.6.3 Concatenate Volumes into a Single Archive . . . . . . . . . . . . . 154

9.7 Including a Label in the Archive. . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 9.8 Verifying Data as It is Stored . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 9.9 Write Protection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

Appendix A Changes. . . . . . . . . . . . . . . . . . . . . . 159

Appendix B Configuring Help Summary … 161 Appendix C Tar Internals . . . . . . . . . . . . . . . . . 165

Basic Tar Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 GNU Extensions to the Archive Format . . . . . . . . . . . . . . . . . . . . . . . . . . 173 Storing Sparse Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174

Old GNU Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 PAX Format, Versions 0.0 and 0.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 PAX Format, Version 1.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176

Format of the Incremental Snapshot Files . . . . . . . . . . . . . . . . . . . . . . . . 177 Dumpdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178

Appendix D Genfile . . . . . . . . . . . . . . . . . . . . . . . 181

D.1 Generate Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 D.2 Status Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 D.3 Exec Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184

Appendix E Free Software Needs Free Documentation . . . . . . . . . . . . . . . . . . . . . . . . . 187

Appendix F Copying This Manual . . . . . . . . . 189

F.1 GNU Free Documentation License. . . . . . . . . . . . . . . . . . . . . . . . . . 189

F.1.1 ADDENDUM: How to use this License for your documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196

Appendix G Index of Command Line Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197

Appendix H Index . . . . . . . . . . . . . . . . . . . . . . . . 201

viii GNU tar 1.16.1